### v1

# TropNNC: Structured Neural Network Compression Using Tropical Geometry

Abstract
We present TropNNC, a structured pruning framework for compressing neural networks with linear and convolutional layers and ReLU activations. Our approximation is based on a geometrical approach to machine/deep learning, using tropical geometry and extending the work of Misiakos et al. (2022). We use the Hausdorff distance of zonotopes in its standard continuous form to achieve a tighter approximation bound for tropical polynomials compared to Misiakos et al. (2022). This enhancement allows for superior functional approximations of neural networks, leading to a more effective compression algorithm. Our method is significantly easier to implement compared to other frameworks, and does not depend on the availability of training data samples. We validate our framework through extensive empirical evaluations on the MNIST, CIFAR, and ImageNet datasets. Our results demonstrate that TropNNC achieves performance on par with the state-of-the-art method ThiNet, even surpassing it in compressing linear layers, and to the best of our knowledge, it is the first method that achieves this using tropical geometry.
Keywords: tropical geometry, zonotopes, Hausdorff distance, neural network compression
1 Introduction
In recent years, deep neural networks (DNNs) have significantly advanced the field of computer vision, excelling in tasks such as image classification (Krizhevsky et al., 2012; Simonyan and Zisserman, 2015; He et al., 2016; Rawat and Wang, 2017), object detection (Zou et al., 2023), semantic segmentation (Noh et al., 2015; Guo et al., 2018), and image captioning (Jia et al., 2015; Hossain et al., 2019). These networks have outperformed traditional methods reliant on manually crafted visual features. However, despite their success, deploying DNNs on resource-constrained devices, such as mobile phones or embedded systems, remains challenging due to their substantial computational and storage demands. For example, the VGG-16 model (Simonyan and Zisserman, 2015), with its approximately $138$ million parameters, requires over $500$MB of storage and performs about $31$ billion FLOPs to classify a single image. Such resource-intensive models exceed the capabilities of small devices and pose significant challenges, including energy limitations recently highlighted by the deployment of large language models (Samsi et al., 2023). This has led to a growing focus on network compression techniques (Blalock et al., 2020) as a critical area of research.
Initial attempts to reduce network complexity date back to the 1990s, when LeCun et al. (1989) demonstrated that removing negligible weights from a network had minimal impact on performance. This approach, known as neural network pruning, has since evolved. Han et al. (2015) proposed a simple yet effective pruning strategy: iteratively removing connections with weights below a certain threshold and fine-tuning the network to recover accuracy, resulting in highly sparse models.
These methods belong to the category of unstructured pruning, which alters the network’s structure by eliminating individual weights. While effective in reducing network size, unstructured pruning presents challenges in practical applications. Such pruned networks often require specialized hardware and software for efficient inference due to issues like poor cache locality and jumping memory accesses, which can negate the benefits of sparsity and sometimes even degrade performance (Wen et al., 2016).
To overcome the limitations of unstructured pruning, structured pruning methods, such as channel-level pruning, have been proposed (He and Xiao, 2024). Notably, the ThiNet framework by Luo et al. (2017) prunes entire filters or channels, maintaining the network’s original structure. In another work, Wen et al. (2016) proposed the structured sparsity learning approach, which also includes filter- or channel-wise pruning. This structured approach ensures compatibility with existing deep learning libraries and offers several advantages: it significantly reduces memory footprint, facilitates further compression and acceleration through methods like parameter quantization (Gong et al., 2014; Chen et al., 2015; Wu et al., 2016), and accelerates various vision tasks.
Parallel to these advancements, tropical geometry (Maclagan and Sturmfels, 2021) has emerged as a promising mathematical framework with applications in neural networks and machine learning (Maragos et al., 2021). Tropical geometry, rooted in algebraic geometry, operates over the tropical semiring (Cuninghame-Green, 1994), where conventional addition and multiplication are replaced by min and plus operations, respectively. This transformation turns polynomials into piecewise linear functions, making tropical geometry highly relevant to the study of neural networks with piecewise linear activations like ReLU.
Tropical mathematics encompasses a wide range of applications, which include operations research and scheduling (Cuninghame-Green, 1979), control and optimization (Baccelli et al., 2001; Cohen et al., 2004; Butkovič, 2010; Akian et al., 2012), mathematical physics (Litvinov et al., 2001), speech recognition (Mohri et al., 2002), polyhedral geometry (Gaubert and Katz, 2011), tropical geometry (Maclagan and Sturmfels, 2021), dynamical systems on weighted lattices (Maragos, 2017), finite state transducers (Theodosis and Maragos, 2018), convex regression (Maragos and Theodosis, 2020), and machine learning (Maragos et al., 2021; Gärtner and Jaggi, 2008). Recently, tropical geometry has been applied to the theoretical study of neural networks. For example, Zhang et al. (2018) demonstrated the equivalence of ReLU-activated neural networks with tropical rational mappings, and used zonotopes to compute a bound on the number of linear regions of networks, equal to the one in (Montufar et al., 2014). Other works, like those of Charisopoulos and Maragos (2018); Alfarra et al. (2022); Smyrnis and Maragos (2020), have used tropical geometry to compute bounds on the number of linear regions in neural networks and to develop methods for pruning and compressing these networks.
In this paper, we explore the application of tropical geometry to neural network compression. We build on the work of Misiakos et al. (2022), who utilized tropical geometry for structured neural network compression by geometrically approximating the zonotopes corresponding to the network layers. Our contributions include:
  • Obtaining a tighter bound for the functional approximation of tropical rational mappings than that presented in (Misiakos et al., 2022) using the standard continuous Hausdorff distance.
  • Designing an algorithm that leverages tropical geometry and the Hausdorff distance for the structured compression of neural networks. Our algorithm refines the choice of weights in the compressed network from the clustering done in (Misiakos et al., 2022) and we also provide an iterative technique to further optimize them. Our method is simple to implement and does not depend on the availability of training data samples like other frameworks such as ThiNet (Luo et al., 2017). It is also the first method based on the theory of tropical geometry that compresses convolutional layers.
  • Evaluating our algorithm through experiments on MNIST, Fashion-MNIST, CIFAR-10, CIFAR-100, and ImageNet datasets. Our method outperforms that of Misiakos et al. (2022) and achieves competitive or superior performance compared to ThiNet, especially in compressing linear layers, and superior performance compared to CUP (Duggal et al., 2021), particularly in the VGG architecture.
This work demonstrates the potential of tropical geometry in enhancing neural network compression techniques, providing a theoretical foundation and a practical algorithm that improves efficiency and maintains accuracy.
2 Related Work
Seminal works have showcased that neural networks tend to be hyperparameterized. Denil et al. (2013) demonstrated that neural networks can be reconstructed using a small subset of parameters, which sparked significant interest in network compression techniques. Blalock et al. (2020) reviewed various methods for reducing model size without compromising performance, highlighting the ongoing advancements in this area.
Early methods for network compression focused on unstructured pruning. LeCun et al. (1989) showed that removing unimportant weights leads to minimal performance loss. This approach was further refined by Han et al. (2015), who proposed pruning weights below a certain threshold followed by fine-tuning the network. Despite their effectiveness, these methods require specialized hardware for efficient inference due to their unstructured nature.
Structured pruning techniques maintain the network’s architecture, facilitating better compatibility and efficiency. Wen et al. (2016) introduced structured sparsity learning to compress networks by finding compact representations of filters and channels, preserving the architecture. Luo et al. (2017) proposed ThiNet, a method that greedily prunes filters based on their impact on the next layer, thereby maintaining network structure and efficiency. Yu et al. (2018) presented the Neuron Importance Score Propagation (NISP) algorithm, which extended the ideas of ThiNet by pruning neurons and channels based on their impact on the final layer output. Lin et al. (2020) introduced HRank, which uses feature map information to rank and prune filters. Duggal et al. (2021) proposed Cluster Pruning (CUP), which prunes similar filters by clustering them based on features derived from both incoming and outgoing weight connections. CUP addresses the limitations of prior work by efficiently determining the ideal number of filters to prune in each layer and integrating pruning into the initial training phase, leading to significant savings in training time and maintaining performance. Wang et al. (2021) introduced a novel approach to pruning convolutional neural networks that focuses on reducing structural redundancy rather than merely pruning individual neurons or channels. Their method identifies and eliminates redundant structures within the network architecture, thereby achieving more efficient pruning without compromising the network’s performance. Smyrnis and Maragos (2020) proposed a tropical division algorithm for structured network compression, further enhancing the capabilities of structured pruning techniques. Misiakos et al. (2022) developed Neural Path K-means, which uses clustering techniques based on tropical geometry and functional approximation for structured compression.
Beyond pruning, other state-of-the-art strategies for network compression include parameter quantization and low-rank approximation. Parameter quantization, as explored by Zhou et al. (2017), Zhou et al. (2018) and Han et al. (2015), reduces the number of bits per parameter, significantly lowering memory and computational requirements. Behdin et al. (2023) presented an optimization-based framework for efficient post-training quantization of large language models. Polino et al. (2018) presented quantized distillation, a method that combines quantization with teacher-student distillation. Low-rank approximation (LRA) techniques decompose weight matrices into smaller matrices, reducing parameters and computations. Studies by Denton et al. (2014), Jaderberg et al. (2014), Sindhwani et al. (2015), Ioannou et al. (2016), Tai et al. (2016), and Chen et al. (2018) have shown that LRA can achieve significant speedups and compression with minimal accuracy loss, although it typically requires iterative optimization. These methods can be combined with structured pruning approaches like ThiNet to achieve further model compression and efficiency.
Our framework is a structured pruning method that extends the approach of Misiakos et al. (2022). It prunes networks layer-by-layer, aiming to minimize the error introduced to the input of the next layer. In that sense, it is similar to the ThiNet framework. However, our algorithm is simpler to implement and does not require a training dataset. Our framework leverages clustering and can be naturally extended to non-uniform pruning, drawing parallels to the CUP framework. In fact, it improves upon steps 1 and 3 of CUP. For step 1, we use similar clustering vectors (referred to as filter features by CUP authors Duggal et al. (2021)) for linear layers, and introduce a new method for constructing clustering vectors in convolutional layers. Our main contribution lies in step 3, where we employ tropical geometry to select a better representative for each cluster, rather than simply choosing the neuron or channel with the most energy. While it is tempting to compare our framework with those of NISP (Yu et al., 2018) and HRank (Lin et al., 2020) given that they are both structured pruning methods, it is important to recognize the distinctions between them and ours. NISP and HRank are advanced frameworks specifically designed to address the limitations of layer-by-layer pruning approaches, such as those employed by ThiNet and our own method. In contrast, our framework leverages a rich theoretical foundation and does not rely on training data samples, offering a unique advantage in certain contexts. We believe our results are promising and present significant advantages. Our work highlights the potential of tropical geometrical techniques and paves the way for further research that could enhance their competitiveness with current state-of-the-art methods.
3 Tropical Geometry of Neural Networks
In this section, we introduce the fundamental background of tropical geometry and its applications to neural network compression.
3.1 Background on Tropical Geometry
Tropical algebra is a field of mathematics that studies the tropical semiring. Tropical geometry is the counterpart of algebraic geometry in the tropical setting. The tropical semiring can refer to either the min-plus semiring or the max-plus semiring (Butkovič, 2010). In this work, we adhere to the convention of using the max-plus semiring $(\mathbb{R}_{\mathrm{max}},\vee,+)$, defined as the set $\mathbb{R}_{\mathrm{max}}=\mathbb{R}\cup\{-\infty\}$ equipped with two binary operations: $\vee$ (ordinary max) and $+$ (ordinary sum).
3.1.1 Tropical Polynomials and Rational Mappings
Within the max-plus semiring, we can define polynomials. A tropical polynomial $f$ in $d$ variables $\mathbf{x}=(x_{1},\ldots,x_{d})$ is defined as the function:
$f(\mathbf{x})=\bigvee_{i=1}^{n}\left\{\mathbf{a}_{i}^{T}\mathbf{x}+b_{i}\right\}=\max_{i\in[n]}\left\{\mathbf{a}_{i}^{T}\mathbf{x}+b_{i}\right\},$ (1)
where $[n]=\{1,...,n\}$. Here, $n$ represents the rank of the tropical polynomial. Each monomial term $\{\mathbf{a}_{i}^{T}\mathbf{x}+b_{i}\}$ of the polynomial has an exponent or slope $\mathbf{a}_{i}\in\mathbb{R}^{d}$ and a coefficient or bias $b_{i}\in\mathbb{R}$. Each monomial term corresponds to a plane in $\mathbb{R}^{d+1}$. Consequently, tropical polynomials are piecewise linear convex functions. Specifically, every tropical polynomial is a continuous piecewise linear convex function, and every continuous piecewise linear convex function can be expressed (though not uniquely) as a tropical polynomial (Maclagan and Sturmfels, 2021). The set of tropical polynomials in $\mathbf{x}$ defines the semiring $\mathbb{R}_{\mathrm{max}}[\mathbf{x}]$. Figures 1(a), 1(b) illustrate examples of tropical polynomials in one and multiple variables, respectively.
Tropical rational mappings extend the concept of rational polynomial functions to tropical algebra. A tropical rational mapping or function is defined as the tropical multiplication of a tropical polynomial $p$ by the tropical multiplicative inverse $q^{-1}$ of another tropical polynomial $q$. In conventional arithmetic, this operation corresponds to the difference of two tropical polynomials $p$ and $q$:
$r(\mathbf{x})=p(\mathbf{x})-q(\mathbf{x})$
Tropical rational mappings correspond to general piecewise linear functions. Specifically, every tropical rational mapping is a continuous piecewise linear function, and every continuous piecewise linear function can be expressed (though not uniquely) as a tropical rational mapping. Figures 1(c), 1(d) provide examples of tropical rational mappings.
(a) $f(x)=\max\{x,1\}$
(b) $f(x,y)=\max\{x,y,1\}$
(c) $f(x)=\max\{2x-2,-x\}-\max\{-2x-2,x\}$
(d) $f(x,y)=\max\{x,-x\}-\max\{y,-y\}$
Figure 1: (a) depicts a single-variate tropical polynomial, (b) depicts a multi-variate tropical polynomial, (c) depicts a single-variate tropical rational mapping, (d) depicts a multi-variate tropical rational mapping
3.1.2 Newton Polytopes
As with algebraic geometry (Hartshorne, 2013), in tropical geometry we can define the Newton polytope (Monical et al., 2019) of a tropical polynomial. Newton polytopes can be used to analyse the behavior of a polynomial. They connect tropical geometry with polytope theory, an extensively studied field of mathematics (Ziegler, 2012). For a tropical polynomial $f$ as defined in (1), we define its Newton polytope as the convex hull of the slopes $\mathbf{a}_{i}$ of $f$
$\mathrm{Newt}(f):=\mathrm{conv}\left\{\mathbf{a}_{i},i\in[n]\right\}.$
Additionally, we define the extended Newton polytope of a tropical polynomial $f$ as the convex hull of the slopes $\mathbf{a}_{i}$ of $f$ extended in their last dimension by the coefficient $b_{i}$
$\mathrm{ENewt}(f):=\mathrm{conv}\left\{(\mathbf{a}_{i}^{T},b_{i}),i\in[n]\right\}.$
The following proposition allows us to calculate the (extended) Newton polytope of expressions of tropical polynomials.
Proposition 1 (Zhang et al., 2018)
Let $f,g\in\mathbb{R}_{\mathrm{max}}[\mathbf{x}]$ be two tropical polynomials in $\mathbf{x}$. For the extended Newton polytope, the following holds:
$\mathrm{ENewt}\left(f\vee g\right)$ $=\mathrm{conv}\left\{\mathrm{ENewt}(f)\cup\mathrm{ENewt}(g)\right\}$
$\mathrm{ENewt}(f+g)$ $=\mathrm{ENewt}(f)\oplus\mathrm{ENewt}(g).$
Here, $\vee$ denotes the tropical addition (taking the maximum), and $+$ denotes the tropical multiplication (ordinary addition). The Minkowski sum $\oplus$ of two polytopes (or more generally subsets of $\mathbb{R}^{d}$) $P$ and $Q$ is defined as:
$P\oplus Q:=\{p+q\mid p\in P,q\in Q\}.$
Using the principle of induction, Proposition 1 can be generalized to any finite tropical expression of tropical polynomials.
The extended Newton polytope provides a geometrical interpretation for studying tropical polynomials. The following propositions establish the relationship between tropical polynomials and their extended Newton polytopes. The upper envelope or upper hull $UF(P)$ of an extended Newton polytope is defined as the set of all points $(\mathbf{a}^{T},b)$ of the polytope $P$ that are not ”shadowed” by any other part of the polytope when viewed from above (last dimension). This means that there is no $b^{\prime}>b$ such that $(\mathbf{a}^{T},b^{\prime})$ belongs to $P$. We have the following useful lemma (proof in Appendix A.1).
Lemma 2
Let $p\in\mathbb{R}_{\mathrm{max}}[\mathbf{x}]$ a tropical polynomial in $d$ variables with extended Newton polytope $P=\mathrm{ENewt}(p)$. If $(\mathbf{a}^{T},b)$ lies below the upper envelope of $P$, then $\forall\mathbf{x}\in\mathbb{R}^{d},\mathbf{a}^{T}\mathbf{x}+b\leq p(\mathbf{x})$. The inequality is strict if $(\mathbf{a}_{i}^{T},b_{i})$ lies strictly below the upper envelope.
An immediate consequence of the above lemma is the following theorem:
Theorem 3
Let $f\in\mathbb{R}_{\mathrm{max}}[\mathbf{x}]$ be a tropical polynomial. The values of $f$ are fully determined by the upper envelope $UF(\mathrm{ENewt}(f))$ of its extended Newton polytope.
Theorem 4 (Charisopoulos and Maragos, 2018)
The linear regions of a tropical polynomial $f\in\mathbb{R}_{\mathrm{max}}[\mathbf{x}]$ are in one-to-one correspondence with the vertices of $UF(\mathrm{ENewt}(f))$. In fact, each vertex $(\mathbf{a}_{i}^{T},b_{i})\in UF(\mathrm{ENewt}(f))$ gives exactly one linear region $D$ where $f(\mathbf{x})=\mathbf{a}_{i}^{T}\mathbf{x}+b_{i}$.
The above theorem highlights the direct relationship between the geometric structure of the extended Newton polytope and the piecewise linear nature of the tropical polynomial. An immideate consequence of Theorems 3 and 4 is the following theorem:
Theorem 5
For any two tropical polynomials $f,g\in\mathbb{R}_{\mathrm{max}}[\mathbf{x}]$, the following holds:
$f=g\Leftrightarrow UF(\mathrm{ENewt}(f))=UF(\mathrm{ENewt}(g))$
This implies that two tropical polynomials are functionally identical if and only if their extended Newton polytopes have the same upper envelope.
The above theorems indicate that a tropical polynomial is fully functionally determined by the upper envelope of its extended Newton polytope, as shown by the following example.
Example 1
Consider the polynomials corresponding to Figure 2.
$f(x,y)=\max\{0,-y+1,y+1\}$
$g(x,y)=\max\{x+1,-x+1\}.$
We have that
$(f\vee g)(x,y)=\max\{0,x+1,y+1,-y+1,-x+1\}$
$(f+g)(x,y)=\max\{x+1,-x+1,x-y+2,-x-y+2,x+y+2,-x+y+2\}.$
The extended Newton polytopes of $f,g,f\vee g,f+g$ are shown in Figure 2. The polynomial $f\vee g$ can be reduced as follows:
$(f\vee g)(x,y)=\max\{x+1,y+1,-y+1,-x+1\},$
which correspond to the vertices of the upper envelope of $\mathrm{ENewt}(f\vee g)$.
(a) $\mathrm{ENewt}(f)$
(b) $\mathrm{ENewt}(g)$
(c) $\mathrm{ENewt}(f\vee g)$
(d) $\mathrm{ENewt}(f+g)$
Figure 2: Operations on tropical polynomials. $\mathrm{ENewt}(f\vee g)$ corresponds to the convex hull of the union of the vertices of the polytopes $\mathrm{ENewt}(f),\mathrm{ENewt}(g)$. $\mathrm{ENewt}(f+g)$ corresponds to the Minkowski sum of $\mathrm{ENewt}(f),\mathrm{ENewt}(g)$. For the polytope $\mathrm{ENewt}(f\vee g)$ we illustrate with blue the upper envelope, which consists of a single face. The vertices of the upper envelope are the only non reduntant terms of the polynomial $f\vee g$.
3.1.3 Zonotopes
In polytope theory, zonotopes are a special class of convex polytopes that can be defined as the Minkowski sum of a finite set of line segments (or edges). Formally, given a set of line segments $g_{1},\ldots,g_{n}$, the zonotope is defined as
$Z:=\bigoplus_{i\in[n]}g_{i}$
The line segments $g_{i}$ are referred to as the zonotope’s generators. Each generator contributes to the shape and size of the zonotope, which is formed by the space swept out by these line segments.
Alternatively, a zonotope can be expressed equivalently by a set of vectors $\mathbf{v}_{1},\ldots,\mathbf{v}_{n}\in\mathbb{R}^{d}$ and a starting point $\mathbf{s}\in\mathbb{R}^{d}$. By taking the generators to be the segments $[\mathbf{0},\mathbf{v}_{1}],\ldots,[\mathbf{0},\mathbf{v}_{n}]$ and translating the first segment by $\mathbf{s}$, we obtain the equivalent form:
$Z=\left\{\mathbf{s}+\sum_{i=1}^{n}\lambda_{i}\mathbf{v}_{i}\mid 0\leq\lambda_{i}\leq 1\right\}$
In this context, the vectors $\mathbf{v}_{i}$ are sometimes referred to as the zonotope’s generators, meaning the segments $[\mathbf{0},\mathbf{v}_{i}]$. When the starting point $\mathbf{s}$ is not mentioned, it is assumed to be the origin $\mathbf{0}$.
For a zonotope with a starting point $\mathbf{s}\in\mathbb{R}^{d}$ and generators $\mathbf{v}_{1},\ldots,\mathbf{v}_{n}\in\mathbb{R}^{d}$, a vertex $\mathbf{u}\in V_{Z}$ corresponds to points where $\lambda_{i}=0\ \text{or}\ 1$. The vertex $\mathbf{u}$ can be expressed as:
$\mathbf{u}=\mathbf{s}+\sum_{i\in I}\mathbf{v}_{i},$
where $I\subset[n]$. It can be shown that zonotopes have up to exponential many verticesIn fact, for generators in general positions it has exactly $2\sum_{j=0}^{d-1}nCr(n-1,j)$ vertices (Gritzmann and Sturmfels, 1993).
Zonotopes exhibit several interesting properties. Notably, they are centrally symmetric. This symmetry, along with their convex structure, make zonotopes particularly useful in various fields, including optimization (Bern and Eppstein, 2001) and computational geometry (Ziegler, 2012).
3.2 Neural Networks with Piecewise Linear Activations
Tropical geometry provides a mathematical framework for analyzing neural networks with piecewise linear activation functions. Leveraging the properties of tropical polynomials and Newton polytopes, it is used to model the behavior of various activation functions commonly used in neural networks. In this work, we focus on ReLU-activated networks. In this section, we outline the basic tropical geometrical properties of ReLU-activated networks and establish our notations.
ReLU Activations. Consider a network which consists of an input layer $\mathbf{x}=(x_{1},\ldots,x_{d})$, a hidden layer $\mathbf{f}=(f_{1},\ldots,f_{n})$ of ReLU units, and an output layer $\mathbf{v}=(v_{1},\ldots,v_{m})$. The input, hidden, and output layers are connected through linear transformations represented by matrices $\mathbf{A}$ and $\mathbf{C}$. Each neuron $i$ has input weights and bias given by $\mathbf{A}_{i,:}=(\mathbf{a}_{i}^{T},b_{i})$ and output weights $\mathbf{C}_{:,i}^{T}=(c_{1i},\ldots,c_{mi})$. We assume the output layer has no bias. Such a network is depicted in Figure 3.
Figure 3: Neural network with one hidden ReLU layer. The first linear layer has weights $\{\mathbf{a}_{i}^{T}\}$ with bias $\{b_{i}\}$ corresponding to $i$-th node $\forall i\in[n]$ and the second has weights $\{c_{ji}\}$, $\forall j\in[m],i\in[n]$
The output of the ReLU unit $i$ is given by:
$f_{i}(\mathbf{x})=\mathrm{ReLU}(\mathbf{a}_{i}^{T}\mathbf{x}+b_{i})=\max\{\mathbf{a}_{i}^{T}\mathbf{x}+b_{i},0\}.$
This expression represents a tropical polynomial of rank 2, with one term being the constant 00. The extended Newton polytope $\mathrm{ENewt}(f_{i})$ of $f_{i}$ is an edge with one endpoint at the origin $\mathbf{0}$ and the other endpoint at $(\mathbf{a}_{i}^{T},b_{i})$. The $j$-th component of the output layer $v_{j}$ can be computed as follows:
$v_{j}=\sum_{i\in[n]}c_{ji}f_{i}=\sum_{i:c_{ji}>0}|c_{ji}|f_{i}-\sum_{i:c_{ji}<0}|c_{ji}|f_{i}=p_{j}-q_{j}.$
In the above expression, $|c_{ji}|f_{i}$ are tropical polynomials. Thus, $p_{j}$ and $q_{j}$ are tropical polynomials formed by the addition of tropical polynomials. Consequently, $v_{j}$ is a tropical rational mapping. We call $p_{j}$ the positive polynomial and $q_{j}$ the negative polynomial of $v_{j}$. This result can be extended to deeper networks, as suggested by the following proposition:
Theorem 6 (Zhang et al., 2018)
A ReLU-activated deep neural network $\mathbf{F}:\mathbb{R}^{d}\rightarrow\mathbb{R}^{m}$ is a tropical rational mapping.
The extended Newton polytope of $|c_{ji}|f_{i}$ is an edge with one endpoint at the origin $\mathbf{0}$ and the other at $|c_{ji}|(\mathbf{a}_{i}^{T},b_{i})$. The extended Newton polytope $P_{j}$ of $p_{j}$ is the Minkowski sum of the positive generators $\{|c_{ji}|(\mathbf{a}_{i}^{T},b_{i}):c_{ji}>0\}$, and the polytope $Q_{j}$ of $q_{j}$ is the Minkowski sum of the negative generators $\{|c_{ji}|(\mathbf{a}_{i}^{T},b_{i}):c_{ji}<0\}$. Thus, $P_{j},Q_{j}$ are zonotopes. We refer to $P_{j}$ as the positive zonotope and $Q_{j}$ as the negative zonotope of $v_{j}$.
4 Approximation based on Hausdorff distance
In this section, we present our improved theorem, which uses the Hausdorff distance in its standard continuous form to bound the error between the original and approximate neural networks. Misiakos et al. (2022) derived a bound for the error in approximating the tropical polynomials that represent a neural network. The motivation for geometrically bounding the error of these polynomials stems from Theorem 5. This theorem indicates that two polynomials with the same extended Newton polytopes are functionally equivalent. Consequently, it is expected that two tropical polynomials with approximately equal extended Newton polytopes will attain similar values.
[…]
Definition 7 (Hausdorff distance)
Let $S,\tilde{S}$ be two subsets of $\mathbb{R}^{d}$. The Hausdorff distance $H(S,\tilde{S})$ of the two sets is defined as
$H(S,\tilde{S}):=\max\left\{\sup_{\mathbf{u}\in S}\mathrm{dist}(\mathbf{u},\tilde{S}),\sup_{\mathbf{v}\in\tilde{S}}\mathrm{dist}(S,\mathbf{v})\right\}$
In the case of polytopes, the following lemma (proof in Appendix A.2) reassures us that the suprema in the above expression are attained, and in fact by points in the vertex sets $V_{P},V_{\tilde{P}}$ of the polytopes.
Lemma 8
Due to the convexity and compactness of polytopes, we have that
$H(P,\tilde{P})=\max\left\{\max_{\mathbf{u}\in V_{P}}\mathrm{dist}(\mathbf{u},\tilde{P}),\max_{\mathbf{v}\in V_{\tilde{P}}}\mathrm{dist}(P,\mathbf{v})\right\}$
In their work, Misiakos et al. (2022) used the discrete form of the Hausdorff distance, defined as the Hausdorff distance of the vertex sets of the two polytopes ($DH(P,\tilde{P}):=H(V_{P},V_{\tilde{P}})$), to obtain the following result.
Proposition 9 (Misiakos et al., 2022)
Let $p,\tilde{p}\in\mathbb{R}_{\mathrm{max}}[\mathbf{x}]$ be two tropical polynomials with extended Newton polytopes $P=\mathrm{ENewt}(p)$ and $\tilde{P}=\mathrm{ENewt}(\tilde{p})$. Then,
$\frac{1}{\rho}\max_{\mathbf{x}\in B}|p(\mathbf{x})-\tilde{p}(\mathbf{x})|\leq DH(P,\tilde{P})$
where $B=\{\mathbf{x}\in\mathbb{R}^{d}:\|\mathbf{x}\|\leq r\}$ and $\rho=\sqrt{r^{2}+1}$.
We extend the result of Misiakos et al. (2022). The proof is provided in the appendix.
Proposition 10
Let $p,\tilde{p}\in\mathbb{R}_{\mathrm{max}}[\mathbf{x}]$ be two tropical polynomials with extended Newton polytopes $P=\mathrm{ENewt}(p)$ and $\tilde{P}=\mathrm{ENewt}(\tilde{p})$. Then,
$\frac{1}{\rho}\max_{\mathbf{x}\in B}|p(\mathbf{x})-\tilde{p}(\mathbf{x})|\leq H(P,\tilde{P})$
where $B=\{\mathbf{x}\in\mathbb{R}^{d}:\|\mathbf{x}\|\leq r\}$ and $\rho=\sqrt{r^{2}+1}$.
Since $V_{P}\subset P,V_{\tilde{P}}\subset\tilde{P}$, we have that $\mathrm{dist}(\mathbf{u},\tilde{P})\leq\mathrm{dist}(\mathbf{u},V_{\tilde{P}}),\mathrm{dist}(P,\mathbf{v})\leq\mathrm{dist}(V_{P},\mathbf{v})$. The equality in the aforementioned inequalities is achieved only in special cases and thus the bound $H(P,\tilde{P})$ we provide is in general tighter than $DH(P,\tilde{P})$ from Misiakos et al. (2022). We demonstrate one example where the two formulations give equal distances and one where the standard Hausdorff distance is strictly smaller than the discrete one.
Example 2
Let $P$ be the zonotope formed by taking the generators $(1,0)$ and $(\cos 60^{\circ}=\frac{1}{2},\sin 60^{\circ}=\frac{\sqrt{3}}{2})$. Form zonotope $\tilde{P}$ by taking as a generator the mean of the two generators of $P$, which is the vector $(\frac{3}{4},\frac{2+\sqrt{3}}{4})$. Refer to Figure 4(a) for a visual representation.
The discrete Hausdorff distance between the two polytopes is given by:
$DH(P,\tilde{P})=\max\left\{\frac{\sqrt{3}}{2},\frac{1}{2}\right\}=\frac{\sqrt{3}}{2}.$
In this case, the Hausdorff distance is equal to the discrete one:
$H(P,\tilde{P})=\max\left\{\frac{\sqrt{3}}{2},\sin 30^{\circ}\cdot|\text{generator of }\tilde{P}|\right\}$
$=\max\left\{\frac{\sqrt{3}}{2},\frac{1}{2}\cdot\cos 30^{\circ}=\frac{\sqrt{3}}{4}\right\}=\frac{\sqrt{3}}{2}.$
Example 3
Let zonotope $P$ be formed by taking the generators $(1,0)$ and $(\cos 60^{\circ}=\frac{1}{2},\sin 60^{\circ}=\frac{\sqrt{3}}{2})$. Form zonotope $\tilde{P}$ by taking as a generator the sum of the two generators, which results in the vector $(\frac{3}{2},\frac{2+\sqrt{3}}{2})$. Refer to Figure 4(b) for a visual representation.
The discrete Hausdorff distance between the two polytopes is calculated as:
$DH(P,\tilde{P})=\max\{0,1\}=1.$
Here, the Hausdorff distance is strictly smaller and equal to:
$H(P,\tilde{P})=\max\left\{0,\frac{1}{2}\right\}=\frac{1}{2}.$
(a) Example 2
(b) Example 3
Figure 4: Polytopes $P$, $\tilde{P}$ of examples 2, 3
Notice that if our goal were to approximate the polytope $P$ by the polytope $\tilde{P}$, then according to the discrete Hausdorff distance, taking the mean is the optimal strategy and gives distance $\sqrt{3}/2$ (example 2), better than $1$ (example 3). On the other hand, according to the Hausdorff distance in its standard form, taking the sum is the optimal strategy and gives distance $1/2$ (example 2), better than $\sqrt{3}/2$ (example 3). This implies that an algorithm based on the Hausdorff distance would imply a better approximation of $\tilde{P}$ in the above example.
We derive the following result, which differs from (Misiakos et al., 2022) by using the standard form of the Hausdorff distance and thus providing a tighter bound. For the proof we apply triangle inequality and then repeatedly Proposition 10 to each output of the network. The full proofs of Proposition 10 and Theorem 11 are provided in Appendix A.3.
Theorem 11
Let $v$ and $\tilde{v}$ be the outputs of two neural networks as in Figure 3. Then, the following inequality holds:
$\frac{1}{\rho}\max_{\mathbf{x}\in B}\|v(\mathbf{x})-\tilde{v}(\mathbf{x})\|_{1}\leq\sum_{j=1}^{m}\left(H(P_{j},\tilde{P}_{j})+H(Q_{j},\tilde{Q}_{j})\right),$
where $\|\cdot\|_{1}$ denotes the $L^{1}$ norm.
5 Compression Algorithm
In this section we present our algorithm TropNNC for neural network compression. Our method, based on Theorem 11, is a structured neural network compression method that reduces the number of neurons in the hidden layers and produces a compressed network that is a functional approximation of the original. We differentiate between single-output and multi-output networks and introduce novel approaches for both scenarios. The main achievement of our method is to preserve the neural network as a function, achieving performance comparable to state-of-the-art frameworks, all without relying on training data samples.
5.1 Single Output
First, we focus on single output networks, as the one depicted in Figure 3, with $m=1$. The network has weight matrices $\mathbf{A}\in\mathbb{R}^{n\times d}$, $\mathbf{C}\in\mathbb{R}^{1\times n}$, and a single output $v$ with corresponding positive and negative zonotopes $P$ and $Q$. We aim to approximate this network with a smaller one that has fewer neurons in the hidden layer, say $K<n$, resulting in a new hidden layer $\tilde{\mathbf{f}}=(\tilde{f}_{1},\ldots,\tilde{f}_{K})$ with new weight matrices $\widetilde{\mathbf{A}}$ and $\widetilde{\mathbf{C}}$. Our goal is to ensure that $\tilde{v}(\mathbf{x})\approx v(\mathbf{x})$ for all $\mathbf{x}\in B$. According to Theorem 11, it is sufficient to choose the new weight matrices $\widetilde{\mathbf{A}}$ and $\widetilde{\mathbf{C}}$ such that $\widetilde{P}\approx P$ and $\widetilde{Q}\approx Q$, where the approximation relation is in terms of the Hausdorff distance between the zonotopes.
In the original network with $n$ hidden neurons, each neuron contributes a generator $|c_{i}|(\mathbf{a}_{i}^{T},b_{i})$, which can be either positive or negative. In the approximating network with $K$ hidden neurons, each neuron contributes a generator $|\tilde{c}_{i}|(\tilde{\mathbf{a}}_{i}^{T},\tilde{b}_{i})$. We aim to pick generators $|\tilde{c_{i}}|(\tilde{\mathbf{a}}_{i}^{T},\tilde{b}_{i})$ such that the generated positive and negative zonotopes $\widetilde{P}$ and $\widetilde{Q}$ approximate the positive and negative zonotopes $P$ and $Q$, respectively.
In essence, we have translated the neural network approximation problem into a zonotope approximation problem—specifically, the task of approximating a zonotope with another zonotope that has fewer generators. This problem is known in the literature as zonotope order reduction (Yang and Scott, 2018). In our case, the approximation must happen in terms of the Hausdorff distance.
5.1.1 Algorithm for single output network
Misiakos et al. (2022) proposed the algorithm ”Zonotope K-means” for compressing single-output neural networks. Their approach uses K-means clustering on the set of positive and negative generators, and replaces the generators of each formed cluster by a single representative, the center of the cluster (i.e. the mean of the generators of the cluster).
In contrast, our approach also uses K-means to cluster the generators, but instead of taking the representative to be the center of the cluster, we take the representative to be the sum of the generators of the cluster. We show later that this compression produces a tighter approximation of the neural network compared to (Misiakos et al., 2022). We demonstrate this with Example 4 and prove this formally with Proposition 12 and Corollary 13. Our Algorithm 1 is depicted below.
Algorithm 1 TropNNC, Single output
1:  Split the generators $|c_{i}|(\mathbf{a}_{i}^{T},b_{i})$ into positive and negative generators:
[…]
2:  Execute K-means with $K/2$ centers on the positive generators $\{|c_{i}|(\mathbf{a}_{i}^{T},b_{i}):c_{i}>0\}$, and $K/2$ centers on the negative generators $\{|c_{i}|(\mathbf{a}_{i}^{T},b_{i}):c_{i}<0\}$.
3:  Obtain positive and negative cluster representatives:
[…]
where $C_{+}\sqcup C_{-}=[K]$ and $|\tilde{c}_{i}|(\tilde{\mathbf{a}}_{i}^{T},\tilde{b}_{i})$ is the sum of the generators of cluster $i$.
4:  For each $i\in[K]$ construct a (hidden layer) neuron with input weights and bias $|\tilde{c}_{i}|(\tilde{\mathbf{a}}_{i}^{T},\tilde{b}_{i})$.
5:  For each constructed neuron $i$ set the output weight to $1$ if the neuron corresponds to a representative of a positive cluster ($i\in C_{+}$), otherwise set it to -1.
Example 4
Suppose we have a positive zonotope consisting of 3 positive generators $g_{1},g_{2},g_{3}$ like the pink zonotope (hexagon) in Figure 5(a), and suppose we have $K/2=2$. The Zonotope K-means algorithm of Misiakos et al. (2022) will form 2 clusters of generators, the black cluster which consists of a single generator $g_{1}$, and the pink cluster which consists of the pink generators $g_{2},g_{3}$. Zonotope K-means will take as cluster representatives the centers of the clusters. In our example, this will be the black generator $\tilde{g}_{1}=g_{1}$ for the black cluster, and the blue generator $\tilde{g}_{2}=\frac{g_{2}+g_{3}}{2}$ for the pink cluster, as in Figure 5(b). The blue zonotope is the zonotope generated by the representatives of the clusters (black and blue generators $\tilde{g}_{1},\tilde{g}_{2}$ respectively), and it is the approximation that Zonotope K-means uses to approximate the pink zonotope.
In contrast, TropNNC will also form the same 2 clusters of generators, but instead of taking the center, it will take the sum of the generators of each cluster as the cluster representative. In our example, these will be the black generator $\tilde{g}_{1}=g_{1}$ and the blue generator $\tilde{g}_{2}=g_{2}+g_{3}$ of Figure 5(c), which form a different blue zonotope which is used as the approximation of the pink zonotope.
In this example, one can visually verify that according to the Hausdorff distance, the approximation of TropNNC is better than the approximation of Zonotope K-Means.
(a) Original Zonotope
(b) Zonotope K-means
(c) TropNNC
Figure 5: Example execution of Zonotope K-means and Improved Zonotope K-means
The result of Example 4 can be generalized. Indeed, the following bound and its corollary hold (proofs in Appendix A.4 and A.5).
Proposition 12
Suppose the clusters K are enough so that for every cluster, no 2 generators of the cluster form an obtuse angle. Then, single-output TropNNC produces a neural network with output $\tilde{v}$ satisfying:
$\frac{1}{\rho}\max_{\mathbf{x}\in B}|u(\mathbf{x})-\tilde{v}(\mathbf{x})|\leq\sum_{i\in I}\min\{||c_{i}(\mathbf{a}_{i}^{T},b_{i})||,\delta_{max}\},$
where $\delta_{max}$ is the largest distance from a point to its corresponding cluster center.
Corollary 13
The above bound is tighter than the bound of Zonotope K-means of Misiakos et al. (2022).
5.2 Multi-output
Now we consider the case of multi-output networks as in Figure 3, for general $m\in\mathbb{N}$. We consider a network with output $\mathbf{v}=(v_{1},\ldots,v_{m})$ and corresponding positive and negative zonotopes $P_{j}$ and $Q_{j}$ for each output node $v_{j}$. An interesting property of these polytopes is that they share the directions $(\mathbf{a}_{i}^{T},b_{i})$ of their generators. For instance, output $v_{1}$ might have a positive generator $|c_{1i}|(\mathbf{a}_{i}^{T},b_{i})$ of zonotope $P_{1}$, while output $v_{2}$ might have a negative generator $|c_{2i}|(\mathbf{a}_{i}^{T},b_{i})$ of zonotope $Q_{2}$. These generators are parallel to each other, with common direction $(\mathbf{a}_{i}^{T},b_{i})$.
As before, we aim to approximate this network with a smaller network that has fewer neurons in the hidden layer, say $K<n$, resulting in a new hidden layer $\tilde{\mathbf{f}}=(\tilde{f}_{1},\ldots,\tilde{f}_{K})$ with new weight matrices $\widetilde{\mathbf{A}}$ and $\widetilde{\mathbf{C}}$. Our goal is to ensure that $\tilde{\mathbf{v}}(\mathbf{x})\approx\mathbf{v}(\mathbf{x})$ for all $\mathbf{x}\in B$. As we have seen, this can be translated to having $\widetilde{P}_{j}\approx P_{j}$ and $\widetilde{Q}_{j}\approx Q_{j}$, where the approximation relation is in terms of the Hausdorff distance between the zonotopes.
Notice that for each new neuron $\tilde{f}_{i}$, its input weight-bias $(\tilde{\mathbf{a}}_{i}^{T},\tilde{b}_{i})$ defines a generator direction for every output $\tilde{v}_{j}$, where $j\in[m]$. For this generator direction, the magnitudes of the generators change across different outputs by adjusting the output weight $\tilde{c}_{ji}$ for each $j\in[m]$, resulting in parallel generators for each output: $|\tilde{c}_{1i}|(\tilde{\mathbf{a}}_{i}^{T},\tilde{b}_{i}),\ldots,|\tilde{c}_{mj}|(\tilde{\mathbf{a}}_{i}^{T},\tilde{b}_{i})$. These generators can be either positive or negative according to the sign of $\tilde{c}_{ji}$.
Thus, we start with original positive and negative zonotopes, with parallel generators, and aim to approximate them with new positive and negative zonotopes, also with parallel generators. We refer to this complex approximation problem as simultaneous zonotope approximation.
Example 5
Suppose we have a neural network with a single hidden layer, as in Figure 3, with dimensions $d+1=n=m=2$. Consider input weights $(\mathbf{a}_{1}^{T},b_{1})=(1,0),(\mathbf{a}_{2}^{T},b_{2})=(0,1)$ and output weights $c_{11}=3,c_{12}=5,c_{21}=4,c_{22}=2$. In this example, for simplicity we took all output weights to be positive so that we only deal with positive zonotopes. The zonotopes of the two outputs will be two parallelograms with parallel edges, as illustrated in the Figure 6. The zonotope of the first output is generated by $c_{11}(\mathbf{a}_{1}^{T},b_{1})=3(1,0)=(3,0)$ and $c_{12}(\mathbf{a}_{2}^{T},b_{2})=5(0,1)=(0,5)$, and of the second output by $c_{21}(\mathbf{a}_{1}^{T},b_{1})=4(1,0)=(4,0)$ and $c_{22}(\mathbf{a}_{2}^{T},b_{2})=2(0,1)=(0,2)$. Say we want to reduce the hidden neurons to $K=1,\tilde{\mathbf{f}}=(\tilde{f_{1}})$. If we could approximate each output’s zonotope separately, we could simply apply the single output algorithm and approximate each parallelogram by its diagonal. However, these diagonals are not parallel to each other, and thus can not occur by a single hidden neuron $\tilde{f}_{1}$ with input weights $(\tilde{\mathbf{a}}_{1}^{T},\tilde{b}_{1})$. Instead, we have to choose a single common direction $(\tilde{\mathbf{a}}_{1}^{T},\tilde{b}_{1})$ for both output zonotopes. We can however choose a different magnitude for each output along this common direction. As will be presented in the algorithm below, for the common direction we choose the vector $(\tilde{\mathbf{a}}_{i}^{T},\tilde{b}_{i})=\frac{(\mathbf{a}_{1}^{T},b_{1})+(\mathbf{a}_{2}^{T},b_{2})}{2}=(0.5,0.5)$. For each output $j$, $\tilde{c}_{j1}$ is chosen so that the edge is as close to the diagonal as possible. Specifically, we choose $\tilde{c}_{11}=3+5=8$ and $\tilde{c}_{21}=2+4=6$. The approximation procedure can be seen in Figure 6(c).
(a) Zonotope of $1^{st}$ output
(b) Zonotope of $2^{nd}$ output
(c) Simultaneous zonotope approximation
Figure 6: Example of simultaneous zonotope approximation for a network with 2 outputs and 2 hidden neurons
5.2.1 Non-iterative Algorithm for multi-output network
To tackle this problem of simultaneous zonotope approximation, Misiakos et al. (2022) proposed the ”Neural Path K-means” algorithm. Their approach to the problem is to use K-means clustering on the set of vectors $\left\{(\mathbf{a}_{i}^{T},b_{i},\mathbf{C}_{:,i}^{T}),i\in[n]\right\}$, and replace the vectors of each formed cluster by a single representative, the center of the cluster (i.e. the mean of the vectors of the cluster). Our approach also uses K-means to cluster the same vectors, but instead of taking the representative to be the center of the cluster, we form the representative using the following heuristic:
Suppose you have 2 neurons with input weights $(\mathbf{a}_{1}^{T},b_{1}),(\mathbf{a}_{2}^{T},b_{2})$ and output weights $\mathbf{C}_{:,1}^{T},\mathbf{C}_{:,2}^{T}$, that get clustered together. If $(\mathbf{a}_{1}^{T},b_{1})\approx(\mathbf{a}_{2}^{T},b_{2})\approx\frac{(\mathbf{a}_{1}^{T},b_{1})+(\mathbf{a}_{2}^{T},b_{2})}{2}$, then for every output $j$ the following holds:
$C_{j,1}(\mathbf{a}_{1}^{T},b_{1})+C_{j,2}(\mathbf{a}_{2}^{T},b_{2})\approx(C_{j,1}+C_{j,2})\frac{(\mathbf{a}_{1}^{T},b_{1})+(\mathbf{a}_{2}^{T},b_{2})}{2}.$
Thus, in order to approximate the 2 neurons, it suffices to take the mean of their input weights and the sum of their output weights. This heuristic can also be motivated by viewing at the network: it corresponds to taking the mean in terms of the input activations of the two neurons, and taking the sum as the joint output activations of the two neurons.
We generalize the above to clusters with more neurons. For every cluster $k\in[K]$ with clustered neuron indexes $I_{k}$ and vectors $\left\{(\mathbf{a}_{i}^{T},b_{i},\mathbf{C}_{:,i}^{T}),i\in I_{k}\right\}$ take $(\tilde{\mathbf{a}}_{k}^{T},\tilde{b}_{k})$ to be the mean of $\left\{(\mathbf{a}_{i}^{T},b_{i}),i\in I_{k}\right\}$, take $\widetilde{\mathbf{C}}_{:,k}^{T}$ to be the sum of $\left\{\mathbf{C}_{:,i}^{T},i\in I_{k}\right\}$, and form the representative of the cluster $(\tilde{\mathbf{a}}_{k}^{T},\tilde{b}_{k},\widetilde{\mathbf{C}}_{:,k}^{T})$. This leads to Algorithm 2.
Algorithm 2 (Non-iterative) TropNNC for Multi-output networks
1:  Execute KMeans with $K$ centers on the vectors $(\mathbf{a}_{i}^{T},b_{i},\mathbf{C}_{:,i}^{T})$ for $i\in[n]$, forming $K$ clusters $\{(\mathbf{a}_{i}^{T},b_{i},\mathbf{C}_{:,i}^{T})\mid i\in I_{k}\}$ for $k\in[K]$.
2:  For each $k\in[K]$, form the cluster representative $(\tilde{\mathbf{a}}_{k}^{T},\tilde{b}_{k},\widetilde{\mathbf{C}}_{:,k}^{T})$ as follows:
  • (i)
    Compute $(\tilde{\mathbf{a}}_{k}^{T},\tilde{b}_{k})$ as the mean of the input weights and biases of the vectors in the cluster, i.e., the mean of the set $\{(\mathbf{a}_{i}^{T},b_{i})\mid i\in I_{k}\}$.
  • (ii)
    Compute $\widetilde{\mathbf{C}}_{:,k}^{T}$ as the sum of the output weights of the vectors in the cluster, i.e., the sum of the set $\{\mathbf{C}_{:,i}^{T}\mid i\in I_{k}\}$.
3:  Construct the new hidden layer:
  • (i)
    For the input weights, set the $k$-th row of the weight-bias matrix to $(\tilde{\mathbf{a}}_{k}^{T},\tilde{b}_{k})$.
  • (ii)
    For the output weights, set the $k$-th column to $\widetilde{\mathbf{C}}_{:,k}$.
There are several variations of Algorithm 2. For instance, one could use the sum instead of the mean for the input weights. However, in our experiments, this approach yielded worse results. Another variation involves clustering neurons based on cosine similarity rather than the L2 distance of the vectors, recognizing that neurons with parallel input weights can be clustered together regardless of their output weights. Additionally, one could normalize the vectors based on their input weights by shifting the norm to the output weights before applying K-means. Despite testing these methods, none provided significant advantages. Therefore, we simply recommend Algorithm 2.
5.2.2 Iterative Algorithm for multi-output network
To improve the approximation of Algorithm 2, we again make use of tropical geometry. Specifically, we formulate an optimization problem that takes the output of Algorithm 2 and with an iterative process produces weights that achieve a better simultaneous zonotope approximation.
Motivated by Algorithm 1, assuming the number of null neurons are few (see Misiakos et al. (2022) and Appendix A.6), we wish in terms of every output $j$ the cluster representative $\widetilde{C}_{j,k}(\tilde{\mathbf{a}}_{k}^{T},\tilde{b}_{k})$ to be as close to the cluster sum $\sum_{i\in I_{k}}C_{j,i}(\mathbf{a}_{i}^{T},b_{i})$ as possible. Thus, for every cluster $k$, we have unknowns $(\tilde{\mathbf{a}}_{k}^{T},\tilde{b}_{k},\widetilde{\mathbf{C}}_{:,k}^{T})=(\tilde{a}_{k1},\ldots,\tilde{a}_{kd},\tilde{b}_{k},\widetilde{C}_{1k},\ldots)$, and we wish to find a solution which minimizes the following criterion:
$\sum_{j=1}^{m}\left[\tilde{C}_{jk}(\tilde{\mathbf{a}}_{k}^{T},\tilde{b}_{k})-\sum_{i\in I_{k}}C_{ji}(\mathbf{a}_{i}^{T},b_{i})\right]^{2},$
where $m$ is the number of outputs, and $I_{k}$ is the set of neurons of cluster $k$.
The above optimization problem can be solved by means of iterative alternating minimization. Specifically:
1. Holding the input weights $(\tilde{\mathbf{a}}_{k}^{T},\tilde{b}_{k})$ constant and minimizing with respect to $\widetilde{\mathbf{C}}_{:,k}$, we notice that the terms of the sum are independent and thus can be minimized separately. For each term of the sum, the minimization occurs if we project the sum $\sum_{i\in I_{k}}C_{ji}(\mathbf{a}_{i}^{T},b_{i})$ onto $(\tilde{\mathbf{a}}_{k}^{T},\tilde{b}_{k})$. We have:
$\widetilde{C}_{jk}=\frac{\left\langle\sum_{i\in I_{k}}C_{ji}(\mathbf{a}_{i}^{T},b_{i}),(\tilde{\mathbf{a}}_{k}^{T},\tilde{b}_{k})\right\rangle}{\left\|(\tilde{\mathbf{a}}_{k}^{T},\tilde{b}_{k})\right\|^{2}}$
2. Holding the output weights $\widetilde{\mathbf{C}}_{:,k}$ constant, and minimizing with respect to the input weights $(\tilde{\mathbf{a}}_{k}^{T},\tilde{b}_{k})$, we take the derivative of the criterion with respect to $(\tilde{\mathbf{a}}_{k}^{T},\tilde{b}_{k})$ and set it to zero. We have:
$2\sum_{j=1}^{m}\left[\widetilde{C}_{jk}(\tilde{\mathbf{a}}_{k}^{T},\tilde{b}_{k})-\sum_{i\in I_{k}}C_{ji}(\mathbf{a}_{i}^{T},b_{i})\right]\widetilde{C}_{jk}=0\Leftrightarrow$
$\Leftrightarrow(\tilde{\mathbf{a}}_{k}^{T},\tilde{b}_{k})\sum_{j=1}^{m}\widetilde{C}_{jk}^{2}=\sum_{j=1}^{m}\sum_{i\in I_{k}}C_{ji}(\mathbf{a}_{i}^{T},b_{i})\widetilde{C}_{jk}\Leftrightarrow$
$(\tilde{\mathbf{a}}_{k}^{T},\tilde{b}_{k})=\frac{\sum_{j=1}^{m}\sum_{i\in I_{k}}C_{ji}(\mathbf{a}_{i}^{T},b_{i})\widetilde{C}_{jk}}{\sum_{j=1}^{m}\widetilde{C}_{jk}^{2}}$
We iterate over the above alternating steps for num_iter iterations. Additionally, we can initialize the iteration with the representative obtained from Algorithm 2, which should provide a good starting point. This initialization is expected to reduce the number of required epochs and speed up the compression algorithm. The resulting procedure is detailed in Algorithm 3. We should note that, in practise, the number of null generators is not negligible, and thus our optimization criterion also constitutes a heuristic method. For this reason, the number of iterations should not be excessive. We provide the following bound for the approximation error of our algorithm (proof in Appendix A.6).
Algorithm 3 Iterative TropNNC for Multi-output networks
1:  Execute K-means with K centers on the vectors $(\mathbf{a}_{i}^{T},b_{i},\mathbf{C}_{:,i}^{T}),i\in[n]$, and form K clusters $\{(\mathbf{a}_{i}^{T},b_{i},\mathbf{C}_{:,i}^{T}),i\in I_{k}\},k\in[K]$
2:  for $k\in[K]$ do
3:     Form the initial cluster representative $(\tilde{\mathbf{a}}_{k}^{T},\tilde{b}_{k},\widetilde{\mathbf{C}}_{:,k}^{T})$ as follows:
4:      2i. Take $(\tilde{\mathbf{a}}_{k}^{T},\tilde{b}_{k})$ to be the mean of the input weights + bias of the vectors of the cluster, i.e., the mean of the set of vectors $\{(\mathbf{a}_{i}^{T},b_{i}),i\in I_{k}\}$
5:      2ii. Take $\widetilde{\mathbf{C}}_{:,k}^{T}$ to be the sum of the output weights of the vectors of the cluster, i.e., the sum of the set of vectors $\{\mathbf{C}_{:,i}^{T},i\in I_{k}\}$
6:  end for
7:  for iter=1 to $\text{iter}=1\text{ to }num\_iter$ do
8:     for $k\in[K]$ do
9:        for $j\in[m]$ do
10:           $\widetilde{C}_{jk}\leftarrow\frac{\langle\sum_{i\in I_{k}}C_{ji}(\mathbf{a}_{i}^{T},b_{i}),(\tilde{\mathbf{a}}_{k}^{T},\tilde{b}_{k})\rangle}{\|(\tilde{\mathbf{a}}_{k}^{T},\tilde{b}_{k})\|^{2}}$
11:        end for
12:        $(\tilde{\mathbf{a}}_{k}^{T},\tilde{b}_{k})\leftarrow\frac{\sum_{j=1}^{m}\sum_{i\in I_{k}}C_{ji}(\mathbf{a}_{i}^{T},b_{i})\widetilde{C}_{jk}}{\sum_{j=1}^{m}\widetilde{C}_{jk}^{2}}$
13:     end for
14:  end for
15:  Construct the new linear layer:
16:   For the input weights, the k-th row of weight-bias matrix is $(\tilde{\mathbf{a}}_{k}^{T},\tilde{b}_{k})$
17:   For the output weights, the k-th column is $\widetilde{\mathbf{C}}_{:,k}^{T}$
Proposition 14
Suppose the clusters K are enough so that for every cluster, no two $(\mathbf{a}_{i}^{T},b_{i}),(\mathbf{a}_{i^{\prime}}^{T},b_{i^{\prime}})$ of the cluster form an obtuse angle. Then, a variant (see Appendix) of Iterative TropNNC produces a neural network with output $\tilde{v}$ satisfying:
$\frac{1}{\rho}\max_{\mathbf{x}\in B}\|v(\mathbf{x})-\tilde{v}(\mathbf{x})\|_{1}\leq\sqrt{m}\sum_{i=1}^{n}\min\left\{\|\mathbf{C}_{:,i}\|\|(\mathbf{a}_{i}^{T},b_{i})\|,\frac{l_{k(i)}}{N_{min}}+\|\epsilon_{:,i}\|_{F}\right\}\\ +\sum_{j=1}^{m}\sum_{i\in N_{j}}|c_{ji}|\|(\mathbf{a}_{i}^{T},b_{i})\|$
where:
  • $N_{j}$ is the set of null neurons with respect to output $j$.
  • $k(i)$ is the cluster of neuron $i$.
  • $N_{min}$ is the minimum cardinality of the non-null generators of a cluster.
  • $l_{k}$ is the objective value of the optimization criterion for cluster $k$.
  • $\epsilon_{j,i}$ is the difference/error between $c_{ji}(\mathbf{a}_{i}^{T},b_{i})$ and the cluster mean $\frac{\sum_{i\in I_{jk}}c_{ji}(\mathbf{a}_{i}^{T},b_{i})}{|I_{jk}|}$.
  • $\epsilon_{:,i}=[\epsilon_{1,i},\dots,\epsilon_{m,i}]$.
  • $\|\cdot\|_{F}$ is the Frobenius norm.
6 Discussion and Limitations
Before proceeding to our experiments, we would like to discuss the applicability of our algorithm, highlight some limitations, and explore potential strategies to overcome these challenges.
Deep and Convolutional Networks. Our methods were analyzed previously for the case of networks with a single hidden linear layer with ReLU activations. However, they can be generalized both for deep and convolutional neural networks. In the first case, where we have a deep network with multiple layers, we can view each layer as an instance for our algorithm, by having an input from the previous layer and providing an output to the next layer. Therefore, we can recursively apply our multi-output algorithm to compress each hidden layer of the network.
In the case of convolutional networks, our methods can be extended for convolutional layers with ReLU and maxout or average pooling layers in between them. Here, instead of neurons, we have channels, and instead of weights that connect the neurons, we have kernels that connect the channels. For each hidden layer, the input and output weight matrices $\mathbf{A}$ and $\mathbf{C}$ are 4-dimensional, where each element of the row $(\mathbf{a}_{i}^{T},b_{i})$ and the column $\mathbf{C}_{:,i}$ is a kernel. Compressing a convolutional layer in a structured manner involves clustering channels. To apply our multi-output algorithm, we first flatten the matrices $\mathbf{A}$ and $\mathbf{C}$ so that they become 2-dimensional, by unravelling the kernels row-wise for $\mathbf{A}$ and column-wise for $\mathbf{C}$. We then apply our multi-output algorithm on the flattened matrices. Finally, after obtaining the new $\widetilde{\mathbf{A}}$ and $\widetilde{\mathbf{C}}$, we reshape the matrices to restore their original dimensions as convolutional layer weight matrices. We emphasize that we are the first to apply a tropical geometrical method to convolutional layers.
Batch normalization and other limitations. Our algorithm can not be applied in the case of networks with batch normalization layers in between the convolutional and linear layers. The reason for this is that batch normalization layers have a non-negligible effect on the zonotopes of the network we approximate. Here, we explain two methodologies that can be used to adapt networks with batch normalization to be applicable for our algorithm. These methodologies were used in the experiments to convert the VGG and ResNet models to equivalent models that our algorithm can be applied to. We propose two methodologies that allow most networks with batch normalization layers to be expressed equivalently as networks without batch normalization. As a first idea, we can fuse together a pair (convolutional (or linear) layer, batch normalization layer) into a single convolutional layer as follows: We take the pair we wish to fuse and form a network with just these two layers. We form a second network with a single, new convolutional layer. We then train the second network to give the same output as the first network on random noise input. This should produce a new convolutional layer that approximates the pair (convolutional (or linear) layer, batch normalization layer) with very little accuracy loss. Finally, we replace each such pair with its approximate convolutional (or linear) layer, which gives us an approximately equivalent network without batch normalization. An alternative second method is to shift the computations of the normalization layer directly to the convolutional or linear layer. Specifically, we can view the batch normalization layer being applied as a multiplication with corrective terms and addition with an extra bias term. We can incorporate these computations into the weights and bias of the convolutional or linear layer.
As for other limitations, we should note that our algorithm is designed to work strictly on layers as described above. Namely, convolutional or linear layers with ReLU and maxout or average pooling layers in between. Layers that apply computations, like batch normalization layers or the loops of residual blocks of ResNet networks fundamentally change the theory for the analysis of the network. For example, similary to ThiNet, our methods can not be applied for the compression of the terminal convolutional layer of residual blocks.
Non-uniform pruning. The techniques we have presented apply a uniform pruning ratio to all layers of the network. However, they can easily be extended to non-uniform pruning if instead of K-means we use hierarchical clustering with a global threshold parameter, like in the CUP framework (Duggal et al., 2021). In fact, when used with hierarchical clustering, TropNNC differs from CUP in step 1, where we choose a different approach to build the cluster vectors (or filter features as Duggal et al. (2021) call them) of a convolutional layer, and more importantly in step 3, where we choose a different cluster representative based on tropical geometry.
7 Experiments
We conduct experiments targeting the compression of the linear and convolutional layers of both convolutional and deep neural networks. Our method TropNNC performs layer-by-layer compression of both linear and convolutional layers and is suitable for regression and classification tasks. The empirical results from these experiments demonstrate that our algorithm achieves the expected performance as claimed. Throughout our experiments, we reserve the use of the iterative variant of TropNNC for cases where its slower performance is justified by the expectation of significantly better results compared to the non-iterative variant.
Baselines. We compare our uniform pruning framework with state-of-the-art methods that conduct structured pruning without requiring re-training (or fine-tuning). Specifically, we compare it with the algorithms proposed by Misiakos et al. (2022), ThiNet (Luo et al., 2017), CUP (Duggal et al., 2021) and the simple baselines random and L1 structured pruning. The algorithms presented in (Misiakos et al., 2022) are originally designed for linear layers only. To enable a fair comparison in compressing convolutional layers, we extended their approach using our proposed technique, making them applicable to convolutional layers as well. ThiNet employs a greedy criterion to remove neurons and channels that contribute least to the subsequent layer’s input. For the non-uniform pruning variant of our framework, we compare it with CUP. CUP assigns to each neuron/channel a feature, performs hierarchical clustering based on these features, and replaces each neuron in each cluster by a cluster representative according to a maximum norm criterion. To ensure a fair comparison, we use identical cluster vectors for both frameworks, as defined in CUP, to emphasize the advantages of our tropical geometry-based approach for selecting cluster representatives. Furthermore, the trivial baseline of random structured pruning discards neurons or channels based on a uniform probability distribution, while L1 structured pruning targets those with the smallest L1 norm of their weights or kernels.
Datasets and networks. We evaluate our framework on the MNIST and CIFAR datasets, testing it across various models including simple multi-layer perceptrons (MLPs), convolutional neural networks (CNNs), LeNet (LeCun et al., 1998), AlexNet (Krizhevsky et al., 2012), and VGG (Simonyan and Zisserman, 2015). The non-uniform variant of our algorithm is tested on the CIFAR and ImageNet datasets, with its performance evaluated across models such as VGG, ResNet56-bn, and ResNet18 (He et al., 2016).
7.1 MNIST Dataset, Pairs 3-5 and 4-9
The first experiment is performed on the binary classification tasks of pairs 3/5 and 4/9 of the MNIST dataset and so we can utilize both single-output and multi-output versions of our algorithm for the compression of the final hidden linear layer. In Tables 1 and 2, we compare single-output TropNNC and TropNNC with the corresponding single-output and multi-output algorithms from Misiakos et al. (2022). We use the same network, a simple CNN network with two fully connected layers and a final hidden linear layer of size $1000$, and the same pruning ratios as in the corresponding experiment of Misiakos et al. (2022). According to Tables 1 and 2, our algorithm performs better than Misiakos et al. (2022) with particular difference in performance when using $0.3\%$ of the nodes of the hidden layer.
Percentage of Remaining Neurons Zonotope K-means TropNNC, single output Neural Path K-means TropNNC
100.0 99.42 $\pm$ 0.05 99.42 $\pm$ 0.05 99.44 $\pm$ 0.06 99.44 $\pm$ 0.06
10.0 99.38 $\pm$ 0.04 99.43 $\pm$ 0.04 99.41 $\pm$ 0.06 99.34 $\pm$ 0.16
5.0 99.37 $\pm$ 0.06 99.42 $\pm$ 0.06 99.37 $\pm$ 0.04 99.31 $\pm$ 0.16
1.0 99.33 $\pm$ 0.05 99.41 $\pm$ 0.04 99.32 $\pm$ 0.06 99.33 $\pm$ 0.05
0.5 99.12 $\pm$ 0.28 99.43 $\pm$ 0.04 99.40 $\pm$ 0.08 99.37 $\pm$ 0.04
0.3 95.98 $\pm$ 3.16 99.43 $\pm$ 0.04 98.39 $\pm$ 0.99 99.39 $\pm$ 0.10
Table 1: Comparison of methods on MNIST-3_5
Percentage of Remaining Neurons Zonotope K-means TropNNC, single output Neural Path K-means TropNNC
100.0 99.55 $\pm$ 0.04 99.55 $\pm$ 0.04 99.56 $\pm$ 0.04 99.56 $\pm$ 0.04
10.0 99.60 $\pm$ 0.07 99.54 $\pm$ 0.04 99.52 $\pm$ 0.04 99.50 $\pm$ 0.05
5.0 99.56 $\pm$ 0.04 99.54 $\pm$ 0.04 99.52 $\pm$ 0.04 99.50 $\pm$ 0.03
1.0 99.35 $\pm$ 0.29 99.54 $\pm$ 0.04 99.49 $\pm$ 0.14 99.47 $\pm$ 0.11
0.5 98.03 $\pm$ 0.71 99.54 $\pm$ 0.04 99.01 $\pm$ 0.50 99.48 $\pm$ 0.12
0.3 86.38 $\pm$ 7.59 99.54 $\pm$ 0.04 94.25 $\pm$ 4.57 99.47 $\pm$ 0.09
Table 2: Comparison of methods on MNIST-4_9
7.2 MNIST and Fashion-MNIST Datasets
The second experiment is performed on the MNIST and FASHION-MNIST datasets. This classification task is multiclass, and thus only our multi-output algorithm may be used. Table 3 compares TropNNC with Neural Path K-means of Misiakos et al. (2022) for the same CNN network as above, and for the same pruning ratios as in (Misiakos et al., 2022). We again compress the final linear layer. As shown in the results, for both datasets, our algorithm outperforms Neural Path K-means.
To evaluate the performance of our algorithm in compressing linear layers of deeper networks, we applied TropNNC to ”deepNN,” a fully connected neural network with layer sizes $28\times 28,512,256,128$, and $10$. The performance plots are provided in Figures 7(a) and 7(b). As illustrated, TropNNC outperforms both Neural Path K-means from (Misiakos et al., 2022) and ThiNet from (Luo et al., 2017).
To assess the performance of our algorithm in compressing convolutional layers, we applied TropNNC to ”deepCNN2D,” a LeNet-type convolutional neural network with ReLU activations. The performance plots are provided in Figures 7(c) and 7(d). The results demonstrate that TropNNC outperforms Neural Path K-means from (Misiakos et al., 2022) and matches, or even surpasses, ThiNet from (Luo et al., 2017).
Percentage of Remaining Neurons MNIST FASHION-MNIST
Neural Path K-means TropNNC Neural Path K-means TropNNC
100.0 98.54 $\pm$ 0.16 98.54 $\pm$ 0.16 89.16 $\pm$ 0.21 89.16 $\pm$ 0.21
50.0 97.85 $\pm$ 0.39 98.49 $\pm$ 0.14 88.17 $\pm$ 0.46 89.00 $\pm$ 0.25
25.0 96.69 $\pm$ 1.06 98.36 $\pm$ 0.14 86.33 $\pm$ 0.87 88.70 $\pm$ 0.22
10.0 96.25 $\pm$ 1.39 97.96 $\pm$ 0.35 84.91 $\pm$ 1.28 88.24 $\pm$ 0.40
5.0 95.17 $\pm$ 2.36 97.06 $\pm$ 0.73 81.48 $\pm$ 3.90 87.42 $\pm$ 0.46
Table 3: Comparison of Neural Path K-means and TropNNC on MNIST and FASHION-MNIST
(a) deepNN, MNIST
(b) deepNN, F-MNIST
(c) deepCNN2D, MNIST
(d) deepCNN2D, F-MNIST
Figure 7: Compression of linear and convolutional layers of ReLU neural networks on MNIST datasets.
7.3 CIFAR Datasets
In this experiment, we compress AlexNet and VGG trained on the CIFAR-10 and CIFAR-100 datasets to assess the performance of each compression method. Figures 8(a) and 8(b) illustrate the compression of the linear layers of AlexNet on CIFAR-10 and CIFAR-100, respectively. Additionally, Figures 8(c) and 8(d) show the compression of VGG’s convolutional layers for these datasets.
The results indicate that for larger datasets, TropNNC consistently outperforms Neural Path K-means from (Misiakos et al., 2022). Furthermore, it at least matches or even surpasses the performance of ThiNet from (Luo et al., 2017) for the compression of linear layers. These findings highlight the effectiveness of TropNNC in handling more complex and larger-scale data scenarios. For the compression of convolutional layers of VGG, TropNNC matches ThiNet. For the compression of convolutional layers of VGG, we do not present results for iterative TropNNC due to the excessively long execution time required.
(a) AlexNet, linear, CIFAR10
(b) AlexNet, linear, CIFAR100
(c) VGG, conv., CIFAR10
(d) VGG, conv., CIFAR100
Figure 8: Compression of linear and convolutional layers of AlexNet on CIFAR datasets.
7.4 Non-ReLU Activation
In this experiment, we evaluate the performance of our algorithm in compressing networks that do not use ReLU activations, such as LeNet (LeCun et al., 1998). We conduct experiments on the MNIST and FASHION-MNIST datasets, focusing on compressing the linear and convolutional layers of LeNet. The results are presented in Figures 9(a), 9(b), 9(c) and 9(d). Interestingly, our algorithm demonstrate robust performance even on networks with non-ReLU activations.
(a) LeNet, linear, MNIST
(b) LetNet, linear, F-MNIST
(c) LeNet, conv., MNIST
(d) LetNet, conv., F-MNIST
Figure 9: Compression of LeNet on MNIST Datasets.
7.5 Non-uniform Pruning
In this section, we evaluate the effectiveness of the non-uniform pruning variant of TropNNC. Our experiments here compress various models, such as VGG on CIFAR-10, ResNet56-bn on CIFAR-10, and ResNet18 on ImageNet. The results are summarized in Tables 4, 5, and 6. We should emphasize that we did not apply any form of fine-tuning or re-training.
Our findings indicate that the method we use for selecting cluster representatives demonstrates a clear advantage. Notably, our approach shows a significant performance improvement for the VGG model, while the benefits are comparatively modest for ResNet models. The substantial advantage observed with VGG remains somewhat unexplained, and we have yet to determine why this effect is less pronounced for ResNet architectures. Further investigation may be required to fully understand these discrepancies and to optimize our approach across different model types.
Method Threshold #params $\downarrow$ FLOPS $\downarrow$ Acc. $\uparrow$
Original - 14.7M 0.63G 93.64
CUP 0.15 6.24M 0.46G 92.34
TropNNC 0.17 5.90M 0.45G 93.27
CUP 0.20 4.62M 0.42G 67.82
TropNNC 0.22 4.58M 0.41G 92.91
CUP 0.25 3.61M 0.39G 18.10
TropNNC 0.30 3.42M 0.38G 79.84
Table 4: Comparison of CUP and TropNNC accuracy across different pruning thresholds on CIFAR10, VGG.
Method Threshold #params $\downarrow$ FLOPS $\downarrow$ Acc. $\uparrow$
Original - 0.85M 0.25G 93.67
CUP 0.50 0.83M 0.22G 87.10
TropNNC-3it. 0.50 0.83M 0.22G 88.34
CUP 0.55 0.82M 0.21G 82.18
TropNNC-3it. 0.55 0.82M 0.21G 83.32
CUP 0.60 0.79M 0.20G 59.97
TropNNC-3it. 0.60 0.79M 0.20G 64.16
Table 5: Comparison of CUP and TropNNC (3 iterations) accuracy across different pruning thresholds on CIFAR10, ResNet56-bn.
Method Threshold #params $\downarrow$ FLOPS $\downarrow$ Acc. $\uparrow$
Original - 11.68M 3.63G 69.10
CUP 0.60 11.65M 3.55G 62.30
TropNNC 0.60 11.65M 3.55G 64.80
CUP 0.65 11.60M 3.48G 54.10
TropNNC 0.65 11.60M 3.48G 56.60
CUP 0.70 11.43M 3.32G 29.20
TropNNC 0.70 11.43M 3.32G 33.20
Table 6: Comparison of CUP and TropNNC accuracy across different pruning thresholds on ImageNet, ResNet18.
8 Conclusions and Future work
We improved upon the theoretical results of Misiakos et al. (2022) by using the (usual) Hausdorff distance instead of its discrete counterpart, to obtain a tighter bound for the approximation of tropical polynomials. Leveraging this enhancement we refined the choice of compressed node’s weights that correspond to the K-means clusters from (Misiakos et al., 2022). Our novel algorithm, TropNNC, applied to linear and convolutional layers of deep neural networks outperforms Neural Path K-means Misiakos et al. (2022), and manages to match or even surpass the performance of ThiNet, particularly for the compression of linear layers. In non-uniform pruning it performs better than CUP, with significant improvement in the case of the VGG network architecture. These achievements underscore the potential of tropical geometry in the realm of neural network compression. One promising direction for further investigation could involve optimizing the approximation techniques by considering the upper hulls of the network’s tropical polynomials.
Appendix A Proofs of Propositions and Lemmas
A.1 Proof of Lemma 2
Proof  Since $(\mathbf{a}^{T},b)$ lies below the upper envelope of $P$, there exists a point
$(\mathbf{a}^{T},b^{\prime})=\sum_{i=1}^{k}\lambda_{i}\mathbf{v}_{i},\quad\sum_{i=1}^{k}\lambda_{i}=1$
on a face of the upper envelope of $P$ defined by the points $\mathbf{v}_{1},\ldots,\mathbf{v}_{k}\in V_{UF(P)}$ such that $b^{\prime}\geq b$. Therefore, we have:
$\mathbf{a}^{T}\mathbf{x}+b\leq\mathbf{a}^{T}\mathbf{x}+b^{\prime}=\langle(\mathbf{a}^{T},b^{\prime}),(\mathbf{x},1)\rangle$
$=\sum_{i=1}^{k}\langle\mathbf{v}_{i},(\mathbf{x},1)\rangle\leq\max_{i=1,\ldots,k}\langle\mathbf{v}_{i},(\mathbf{x},1)\rangle\leq p(\mathbf{x})$
If $(\mathbf{a}_{i}^{T},b_{i})$ lies strictly below the upper envelope of $P$, then $b^{\prime}>b$, and the inequality is strict.
A.2 Proof of Lemma 8
Proof  We will prove that
$\sup_{\mathbf{u}\in P}\operatorname{dist}(\mathbf{u},\tilde{P})=\max_{\mathbf{u}\in V_{P}}\operatorname{dist}(\mathbf{u},\tilde{P}),$
i.e., the supremum is attained at some vertex of $P$.
The polytope $P$, the domain of the supremum, is convex and compact. Thus, it suffices to prove that the function
$f(\mathbf{u})=\operatorname{dist}(\mathbf{u},\tilde{P})=\min_{\tilde{\mathbf{u}}\in\tilde{P}}\|\mathbf{u}-\tilde{\mathbf{u}}\|$
is convex in terms of $\mathbf{u}$.
Let $\mathbf{u}_{1},\mathbf{u}_{2}\in P$. By the compactness of $\tilde{P}$, there exist points $\tilde{\mathbf{u}}_{1},\tilde{\mathbf{u}}_{2}\in\tilde{P}$ such that $f(\mathbf{u}_{1})=\operatorname{dist}(\mathbf{u}_{1},\tilde{P})=\|\mathbf{u}_{1}-\tilde{\mathbf{u}}_{1}\|$ and $f(\mathbf{u}_{2})=\operatorname{dist}(\mathbf{u}_{2},\tilde{P})=\|\mathbf{u}_{2}-\tilde{\mathbf{u}}_{2}\|$.
For every $\lambda\in[0,1]$, we have that
$\lambda f(\mathbf{u}_{1})+(1-\lambda)f(\mathbf{u}_{2})$ $=\lambda\|\mathbf{u}_{1}-\tilde{\mathbf{u}}_{1}\|+(1-\lambda)\|\mathbf{u}_{2}-\tilde{\mathbf{u}}_{2}\|$
$\geq\|\lambda(\mathbf{u}_{1}-\tilde{\mathbf{u}}_{1})+(1-\lambda)(\mathbf{u}_{2}-\tilde{\mathbf{u}}_{2})\|$
$=\|\lambda\mathbf{u}_{1}+(1-\lambda)\mathbf{u}_{2}-\lambda\tilde{\mathbf{u}}_{1}-(1-\lambda)\tilde{\mathbf{u}}_{2}\|.$
By the convexity of $\tilde{P}$, $\tilde{\mathbf{u}}=\lambda\tilde{\mathbf{u}}_{1}+(1-\lambda)\tilde{\mathbf{u}}_{2}\in\tilde{P}$. Hence,
$\lambda f(\mathbf{u}_{1})+(1-\lambda)f(\mathbf{u}_{2})\geq\|\lambda\mathbf{u}_{1}+(1-\lambda)\mathbf{u}_{2}-\tilde{\mathbf{u}}\|\geq f(\lambda\mathbf{u}_{1}+(1-\lambda)\mathbf{u}_{2}),$
which concludes the proof.
A.3 Proof of Proposition 10 and Theorem 11
Proof  Consider a point $\mathbf{x}\in B$ and assume that $p(\mathbf{x})=\mathbf{a}^{T}\mathbf{x}+b$ and $\tilde{p}(\mathbf{x})=\mathbf{c}^{T}\mathbf{x}+d$. Take an arbitrary $(\mathbf{u}^{T},v)\in\tilde{P}$. This point lies below the upper envelope of $\tilde{P}$. Thus, by Lemma 2, we have that $\tilde{p}(\mathbf{x})\geq\mathbf{u}^{T}\mathbf{x}+v$. Choose $(\mathbf{u}^{T},v)$ to be the closest point to $(\mathbf{a}^{T},b)$. Then,
$p(\mathbf{x})-\tilde{p}(\mathbf{x})$ $\leq p(\mathbf{x})-(\mathbf{u}^{T},v)\begin{pmatrix}\mathbf{x}\\ 1\end{pmatrix}$
$=((\mathbf{a}^{T},b)-(\mathbf{u}^{T},v))\begin{pmatrix}\mathbf{x}\\ 1\end{pmatrix}$
$\leq\left\|(\mathbf{a}^{T},b)-(\mathbf{u}^{T},v)\right\|\left\|\begin{pmatrix}\mathbf{x}\\ 1\end{pmatrix}\right\|$
$\leq\operatorname{dist}((\mathbf{a}^{T},b),\tilde{P})\cdot\rho$
$\leq\max_{(\mathbf{a}^{T},b)\in V_{P}}\operatorname{dist}((\mathbf{a}^{T},b),\tilde{P})\cdot\rho,$
where the second inequality is due to the Cauchy-Schwarz inequality.
In a similar manner, take an arbitrary $(\mathbf{r}^{T},s)\in P$. This point lies below the upper envelope of $P$. Thus, by Lemma 2, we have that $p(\mathbf{x})\geq\mathbf{r}^{T}\mathbf{x}+s$. Choose $(\mathbf{r}^{T},s)$ to be the closest point to $(\mathbf{c}^{T},d)$. Then,
$p(\mathbf{x})-\tilde{p}(\mathbf{x})$ $\geq(\mathbf{r}^{T},s)\begin{pmatrix}\mathbf{x}\\ 1\end{pmatrix}-\tilde{p}(\mathbf{x})$
$=((\mathbf{r}^{T},s)-(\mathbf{c}^{T},d))\begin{pmatrix}\mathbf{x}\\ 1\end{pmatrix}$
$\geq-\left\|(\mathbf{r}^{T},s)-(\mathbf{c}^{T},d)\right\|\left\|\begin{pmatrix}\mathbf{x}\\ 1\end{pmatrix}\right\|$
$\geq-\operatorname{dist}((\mathbf{r}^{T},s),\tilde{P})\cdot\rho$
$\geq-\max_{(\mathbf{c}^{T},d)\in V_{\tilde{P}}}\operatorname{dist}((\mathbf{c}^{T},d),P)\cdot\rho.$
Finally, we obtain that
$-\max_{(\mathbf{c}^{T},d)\in V_{\tilde{P}}}\operatorname{dist}((\mathbf{c}^{T},d),P)\cdot\rho\leq p(\mathbf{x})-\tilde{p}(\mathbf{x})\leq\max_{(\mathbf{a}^{T},b)\in V_{P}}\operatorname{dist}((\mathbf{a}^{T},b),\tilde{P})\cdot\rho,$
which implies
$\frac{1}{\rho}\lvert p(\mathbf{x})-\tilde{p}(\mathbf{x})\rvert\leq\max\left\{\max_{(\mathbf{a}^{T},b)\in V_{P}}\operatorname{dist}((\mathbf{a}^{T},b),\tilde{P}),\max_{(\mathbf{c}^{T},d)\in V_{\tilde{P}}}\operatorname{dist}((\mathbf{c}^{T},d),P)\right\},\quad\forall\mathbf{x}\in B.$
Therefore, by Lemma 8, we have,
$\frac{1}{\rho}\max_{\mathbf{x}\in B}\lvert p(\mathbf{x})-\tilde{p}(\mathbf{x})\rvert\leq H(P,\tilde{P}),$
which concludes the proof of Proposition 10.
Finally, notice that we may write
$\|\mathbf{v}(\mathbf{x})-\tilde{\mathbf{v}}(\mathbf{x})\|_{1}$ $=\sum_{j=1}^{m}\lvert\mathbf{v}_{j}(\mathbf{x})-\tilde{\mathbf{v}}_{j}(\mathbf{x})\rvert$
$=\sum_{j=1}^{m}\lvert(p_{j}(\mathbf{x})-q_{j}(\mathbf{x}))-(\tilde{p}_{j}(\mathbf{x})-\tilde{q}_{j}(\mathbf{x}))\rvert$
$=\sum_{j=1}^{m}\lvert(p_{j}(\mathbf{x})-\tilde{p}_{j}(\mathbf{x}))-(q_{j}(\mathbf{x})-\tilde{q}_{j}(\mathbf{x}))\rvert$
$\leq\sum_{j=1}^{m}\left(\lvert p_{j}(\mathbf{x})-\tilde{p}_{j}(\mathbf{x})\rvert+\lvert q_{j}(\mathbf{x})-\tilde{q}_{j}(\mathbf{x})\rvert\right)$
By Proposition 10 we derive
$\frac{1}{\rho}\max_{\mathbf{x}\in B}\|\mathbf{v}(\mathbf{x})-\tilde{\mathbf{v}}(\mathbf{x})\|_{1}\leq\sum_{j=1}^{m}\left(H(P_{j},\tilde{P}_{j})+H(Q_{j},\tilde{Q}_{j})\right),$
A.4 Proof of Proposition 12
Proof  For the output function, it holds that
$v(\mathbf{x})=p(\mathbf{x})-q(\mathbf{x}),\quad\tilde{v}(\mathbf{x})=\tilde{p}(\mathbf{x})-\tilde{q}(\mathbf{x})$
From the triangle inequality, we deduce
$\frac{1}{\rho}\left|v(\mathbf{x})-\tilde{v}(\mathbf{x})\right|\leq H(P,\tilde{P})+H(Q,\tilde{Q})$
Thus, it suffices to get a bound on $H(P,\tilde{P})$ and $H(Q,\tilde{Q})$.
Let $I_{+}\subset[n]$ and $I_{-}\subset[n]$ be the sets of positive and negative generators, respectively. First, we deal with $H(P,\tilde{P})$. Notice that $\forall\mathbf{v}\in V_{\tilde{P}},\mathbf{v}$ is the sum of generators of some clusters of $P$. Thus $\forall\mathbf{v}\in V_{\tilde{P}},\mathbf{v}$ is a vertex of $P$. Hence,
$\operatorname{dist}(P,\mathbf{v})=0,\quad\forall\mathbf{v}\in V_{\tilde{P}}.$
Let $I_{k}\subset[n]$ be the set of generators that belong to cluster $k$. Let $C_{+}\subset[K]$ be the set of clusters of positive generators, and $C_{-}\subset[K]$ the set of clusters of negative generators.
Consider any vertex $\mathbf{u}$ of $P$. This vertex can be written as the sum of generators $c_{i}(\mathbf{a}_{i}^{T},b_{i})$, for some subset $I_{+}^{\prime}\subset I_{+}$. Thus,
$\mathbf{u}=\sum_{i\in I_{+}^{\prime}}c_{i}(\mathbf{a}_{i}^{T},b_{i})$
For every generator $c_{i}(\mathbf{a}_{i}^{T},b_{i})$ that belongs to cluster $k$, define
$x_{i}=\arg\min_{x}\left\{c_{i}(\mathbf{a}_{i}^{T},b_{i})-x\tilde{c}_{k}(\tilde{\mathbf{a}}_{k}^{T},\tilde{b}_{k})\right\}$
i.e., project the generator onto its cluster representative. Since every inner-cluster pair of generators forms an acute angle, any generator and its cluster representative (sum of generators of the cluster) will also form an acute angle, and thus $x_{i}\geq 0$.
Define $I_{k}^{\prime}=I_{k}\cap I_{+}^{\prime}$. For every cluster $k$, define
$\tilde{x}_{k}=\sum_{i\in I_{k}^{\prime}}x_{i}$
Since the cluster representative is the sum of the generators of the cluster, we have
$\sum_{i\in I_{k}}x_{i}=1\Rightarrow\tilde{x}_{k}=\sum_{i\in I_{k}^{\prime}}x_{i}\leq 1$
Thus, for every cluster $k$, the point $\tilde{x}_{k}\tilde{c}_{k}(\tilde{\mathbf{a}}_{k}^{T},\tilde{b}_{k}^{T})$ lies inside the segment $[\mathbf{0},\tilde{c}_{k}(\tilde{\mathbf{a}}_{k}^{T},\tilde{b}_{k}^{T})]$ and thus belongs to $\tilde{P}$.
For the vertex $\mathbf{u}$, we choose to compare it with the vertex
$\sum_{k\in C_{+}}\tilde{x}_{k}\tilde{c}_{k}(\tilde{\mathbf{a}}_{k}^{T},\tilde{b}_{k}^{T})\in\tilde{P}.$
Thus, we have that
$\operatorname{dist}(\mathbf{u},\tilde{P})$ $\leq\left\|\sum_{i\in I_{+}^{\prime}}c_{i}(\mathbf{a}_{i}^{T},b_{i})-\sum_{k\in C_{+}}\tilde{x}_{k}\tilde{c}_{k}(\tilde{\mathbf{a}}_{k}^{T},\tilde{b}_{k})\right\|$
$\leq\sum_{k\in C_{+}}\left\|\sum_{i\in I_{k}^{\prime}}c_{i}(\mathbf{a}_{i}^{T},b_{i})-\tilde{x}_{k}\tilde{c}_{k}(\tilde{\mathbf{a}}_{k}^{T},\tilde{b}_{k})\right\|$
$\leq\sum_{k\in C_{+}}\left\|\sum_{i\in I_{k}^{\prime}}\left[c_{i}(\mathbf{a}_{i}^{T},b_{i})-x_{i}\tilde{c}_{k}(\tilde{\mathbf{a}}_{k}^{T},\tilde{b}_{k})\right]\right\|$
$\leq\sum_{k\in C_{+}}\sum_{i\in I_{k}^{\prime}}\left\|c_{i}(\mathbf{a}_{i}^{T},b_{i})-x_{i}\tilde{c}_{k}(\tilde{\mathbf{a}}_{k}^{T},\tilde{b}_{k})\right\|$
$=\sum_{k\in C_{+}}\sum_{i\in I_{k}^{\prime}}\left\|c_{i}(\mathbf{a}_{i}^{T},b_{i})-x_{i}|I_{k}|(c_{i}(\mathbf{a}_{i}^{T},b_{i})+\epsilon_{i})\right\|$
$=\sum_{k\in C_{+}}\sum_{i\in I_{k}^{\prime}}\min_{x}\Bigl{\|}c_{i}(\mathbf{a}_{i}^{T},b_{i})-x|I_{k}|(c_{i}(\mathbf{a}_{i}^{T},b_{i})+\epsilon_{i})\Bigr{\|}$
$=\sum_{k\in C_{+}}\sum_{i\in I_{k}^{\prime}}\min_{x}\left\{|1-x|I_{k}||\cdot\left\|c_{i}(\mathbf{a}_{i}^{T},b_{i})\right\|+|x|I_{k}||\cdot\left\|\epsilon_{i}\right\|\right\}$
$=\sum_{k\in C_{+}}\sum_{i\in I_{k}^{\prime}}\min\left\{\left\|c_{i}(\mathbf{a}_{i}^{T},b_{i})\right\|,\left\|\epsilon_{i}\right\|\right\}$
$\leq\sum_{k\in C_{+}}\sum_{i\in I_{k}^{\prime}}\min\left\{\left\|c_{i}(\mathbf{a}_{i}^{T},b_{i})\right\|,\delta_{\max}\right\}$
$=\sum_{i\in I_{+}^{\prime}}\min\left\{\left\|c_{i}(\mathbf{a}_{i}^{T},b_{i})\right\|,\delta_{\max}\right\}$
where $\epsilon_{i}$ is the error between generator $i$ and the cluster center/mean of K-means.
The maximum value of the upper bound occurs when $I_{+}^{\prime}=I_{+}$. Thus,
$\max_{\mathbf{u}\in V_{P}}\operatorname{dist}(\mathbf{u},\tilde{P})\leq\sum_{i\in I_{+}}\min\left\{\left\|c_{i}(\mathbf{a}_{i}^{T},b_{i})\right\|,\delta_{\max}\right\}$
Finally, we have
$H(P,\tilde{P})$ $=\max\left\{\max_{\mathbf{u}\in V_{P}}\operatorname{dist}(\mathbf{u},\tilde{P}),\max_{\mathbf{v}\in V_{\tilde{P}}}\operatorname{dist}(P,\mathbf{v})\right\}$
$\leq\max\left\{\sum_{i\in I_{+}}\min\left\{\left\|c_{i}(\mathbf{a}_{i}^{T},b_{i})\right\|,\delta_{\max}\right\},0\right\}$
$=\sum_{i\in I_{+}}\min\left\{\left\|c_{i}(\mathbf{a}_{i}^{T},b_{i})\right\|,\delta_{\max}\right\}$
Similarly, for $H(Q,\tilde{Q})$, we have
$H(Q,\tilde{Q})\leq\sum_{i\in I_{-}}\min\left\{\left\|c_{i}(\mathbf{a}_{i}^{T},b_{i})\right\|,\delta_{\max}\right\}$
Combining, we get
$\frac{1}{\rho}\left|v(\mathbf{x})-\tilde{v}(\mathbf{x})\right|\leq\sum_{i\in I}\min\left\{\left\|c_{i}(\mathbf{a}_{i}^{T},b_{i})\right\|,\delta_{\max}\right\}$
which concludes the proof of Proposition 12.
A.5 Proof of Corollary 13
Proof  The bound of Misiakos et al. (2022) is the following:
$\frac{1}{\rho}\max_{\mathbf{x}\in B}|v(\mathbf{x})-\tilde{v}(\mathbf{x})|\leq K\delta_{\max}+\left(1-\frac{1}{N_{\max}}\right)\sum_{i=1}^{n}|c_{i}|\|(\mathbf{a}_{i}^{T},b_{i})\|.$
We will show that
$K\delta_{\max}+\left(1-\frac{1}{N_{\max}}\right)\sum_{i=1}^{n}|c_{i}|\|(\mathbf{a}_{i}^{T},b_{i})\|\geq\sum_{i\in I}\min\left\{\left\|c_{i}(\mathbf{a}_{i}^{T},b_{i})\right\|,\delta_{\max}\right\}.$
This can be rewritten as
$K\delta_{\max}+\left(1-\frac{1}{N_{\max}}\right)\sum_{i=1}^{n}|c_{i}|\|(\mathbf{a}_{i}^{T},b_{i})\|\geq\sum_{i\in I}\left\|c_{i}(\mathbf{a}_{i}^{T},b_{i})\right\|+\sum_{i\in I}\min\left\{0,\delta_{\max}-\left\|c_{i}(\mathbf{a}_{i}^{T},b_{i})\right\|\right\}.$
Further simplifying, we get
$K\delta_{\max}\geq\frac{1}{N_{\max}}\sum_{i=1}^{n}|c_{i}|\|(\mathbf{a}_{i}^{T},b_{i})\|+\sum_{i=1}^{n}\min\left\{0,\delta_{\max}-\left\|c_{i}(\mathbf{a}_{i}^{T},b_{i})\right\|\right\}.$
It suffices to show that for every cluster $k$, we have:
$\delta_{\max}\geq\frac{1}{|I_{k}|}\sum_{i\in I_{k}}|c_{i}|\|(\mathbf{a}_{i}^{T},b_{i})\|+\sum_{i\in I_{k}}\min\left\{0,\delta_{\max}-|c_{i}|\|(\mathbf{a}_{i}^{T},b_{i})\|\right\}.$
However, it holds that
$\sum_{i\in I_{k}}\min\left\{0,\delta_{\max}-|c_{i}|\|(\mathbf{a}_{i}^{T},b_{i})\|\right\}\leq\delta_{\max}-\max_{i\in I_{k}}|c_{i}|\|(\mathbf{a}_{i}^{T},b_{i})\|,$
and
$\max_{i\in I_{k}}|c_{i}|\|(\mathbf{a}_{i}^{T},b_{i})\|\geq\frac{1}{|I_{k}|}\sum_{i\in I_{k}}|c_{i}|\|(\mathbf{a}_{i}^{T},b_{i})\|.$
Hence, we have
$\sum_{i\in I_{k}}\min\left\{0,\delta_{\max}-|c_{i}|\|(\mathbf{a}_{i}^{T},b_{i})\|\right\}+\frac{1}{|I_{k}|}\sum_{i\in I_{k}}|c_{i}|\|(\mathbf{a}_{i}^{T},b_{i})\|\\ \leq\delta_{\max}-\max_{i\in I_{k}}|c_{i}|\|(\mathbf{a}_{i}^{T},b_{i})\|+\frac{1}{|I_{k}|}\sum_{i\in I_{k}}|c_{i}|\|(\mathbf{a}_{i}^{T},b_{i})\|\leq\delta_{\max},$
which concludes the proof.
A.6 Proof of Proposition 14
Before we proceed with the proof of Proposition 14, we first give the definition of null neurons and generators.
Definition 15 (Null neuron/generator)
A neuron/generator $i\in[n]$ that belongs to cluster $k\in[K]$ is a null neuron/generator with respect to output $j\in[m]$ if $\tilde{c}_{jk}c_{ji}\leq 0$. $N_{j}$ is the set of all null neurons with respect to output $j$.
Proof  Assume the algorithm iterative scheme has reached a stationary point (otherwise, assume the last step is an output weight update and the proof works fine). First, we focus on a single output, say $j$-th output. We will bound $H(P_{j},\tilde{P}_{j}),H(Q_{j},\tilde{Q}_{j})$ for all $j\in[m]$.
Let $I_{j+},I_{j-}$ be the sets of positive and negative generators of output $j$. Let $I_{k}$ be the set of neurons that belong to cluster k. Let $C_{j+}$ be the set of positive clusters for output $j$ (i.e. clusters for which $\tilde{c}_{jk}>0$), and $C_{j-}$ be the set of negative clusters for output $j$. Let $I_{jk}=I_{k}\cap I_{j+}$ if $k$ is a positive cluster, else $I_{jk}=I_{k}\cap I_{j-}$
Consider any vertex $\mathbf{u}$ of $P_{j}$. This vertex can be written as the sum of generators $c_{ji}(\mathbf{a}_{i}^{T},b_{i})$, for some subset $I_{j+}^{\prime}\subset I_{j+}$. Thus,
$u=\sum_{i\in I_{j+}^{\prime}}c_{ji}(\mathbf{a}_{i}^{T},b_{i})$
For every generator $c_{ji}(\mathbf{a}_{i}^{T},b_{i})$ that belongs to cluster k, define
$x_{ji}=arg\min_{x}\left\{c_{ji}(\mathbf{a}_{i}^{T},b_{i})-x\tilde{c}_{jk}(\tilde{\mathbf{a}}_{k}^{T},\tilde{b}_{k})\right\}$
i.e., project the generator onto $\tilde{c}_{jk}(\tilde{\mathbf{a}}_{k}^{T},\tilde{b}_{k})$ of its cluster $k$.
Assume a variant of the algorithm, where the optimization criterion is the following:
$\sum_{j=1}^{m}\left[|\tilde{C}_{jk}|(\tilde{\mathbf{a}}_{k}^{T},\tilde{b}_{k})-\sum_{i\in I_{jk}}|C_{ji}|(\mathbf{a}_{i}^{T},b_{i})\right]^{2}.$
The sign of $\tilde{C}_{jk}$ never changes and gets fixed based on the initial solution. The set $I_{jk}$ of the non-null generators of cluster $k$ in terms of output $j$ depends on the sign of $\tilde{C}_{jk}$, and it is determined by the initial solution. The output weight update rule changes: We update the absolute value of the weight $|\tilde{C}_{jk}|$. The update is performed as normal if the result is positive, otherwise we set $|\tilde{C}_{jk}|=0$. We deduce that throughout the execution of the algorithm, after every output weight update step the following holds:
  • If $|\tilde{C}_{jk}|=0$ then, every generator of cluster $k$ is a null generator by definition.
  • If $|\tilde{C}_{jk}|>0$ then the following argument holds.
By hypothesis, for every cluster $k$, the vectors of the set $\{(\mathbf{a}_{i}^{T},b_{i})|i\in I_{k}\}$ form pair-wise acute angles. Every vector of the set of vectors $\{\sum_{i\in I_{jk}}|c_{ji}|(\mathbf{a}_{i}^{T},b_{i})|j\in[m]\}$ lies inside the cone of the set of vectors $\{(\mathbf{a}_{i}^{T},b_{i})|i\in I_{k}\}$. It is easy to verify that the representative $|\tilde{c}_{jk}|(\tilde{\mathbf{a}}_{k},\tilde{b}_{k})$ that the iterative algorithm produces lies inside the cone of the set of vectors $\{\sum_{i\in I_{jk}}|c_{ji}|(\mathbf{a}_{i}^{T},b_{i})|j\in[m]\}$, which is a subset of the cone of the set of vectors $\{(\mathbf{a}_{i}^{T},b_{i})|i\in I_{k}\}$. Thus, the representative forms an acute angle with every vector of the set $\{(\mathbf{a}_{i}^{T},b_{i})|i\in I_{k}\}$, and thus $x_{ji}\geq 0$
Define $I_{jk}^{\prime}=I_{k}\cap I_{j+}^{\prime}$. For every cluster k, define
$\tilde{x}_{jk}=\sum_{i\in I_{jk}^{\prime}}x_{ji}$
Since $x_{ji}\geq 0$, and by the definition of the update step for the output weights, we have
$\sum_{i\in I_{jk}}x_{ji}=1\Rightarrow\tilde{x}_{jk}=\sum_{i\in I_{jk}^{\prime}}x_{ji}\leq 1$
Thus, for every cluster $k$, the point $\tilde{x}_{jk}\tilde{c}_{jk}(\tilde{\mathbf{a}}_{k}^{T},\tilde{b}_{k})$ lies inside the segment $[\text{O},\tilde{c}_{jk}(\tilde{\mathbf{a}}_{k}^{T},\tilde{b}_{k})$], and thus belongs to $\tilde{P_{j}}$.
For every vertex $u$, we choose to compare it with the vertex
$\sum_{k\in C_{j+}}\tilde{x}_{jk}\tilde{c}_{jk}(\tilde{\mathbf{a}}_{k}^{T},\tilde{b}_{k})\in\tilde{P_{j}}$
We have that
$\tilde{c}_{jk}(\tilde{\mathbf{a}}_{k}^{T},\tilde{b}_{k})=\sum_{i\in I_{jk}}c_{ji}(\mathbf{a}_{i}^{T},b_{i})+l_{jk}=|I_{jk}|\frac{\sum_{i\in I_{jk}}c_{ji}(\mathbf{a}_{t}^{T},b_{i})}{|I_{jk}|}+l_{jk}=|I_{jk}|(c_{ji}(\mathbf{a}_{i}^{T},b_{i})+\epsilon_{ji})+l_{jk}$
where $\sum_{j=1}^{m}l_{jk}^{2}=l_{k}^{2}$ the optimization criterion loss, and $\epsilon_{ji}$ is the error between $c_{ji}(\mathbf{a}_{i}^{T},b_{i})$ and the mean $\frac{\sum_{i\in I_{jk}}c_{ji}(\mathbf{a}_{t}^{T},b_{i})}{|I_{jk}|}$. It is easy to verify that $\epsilon_{ji}$ tends to zero as $\delta_{max}$ tends to 0.
Thus, we have that
$\operatorname{dist}(\mathbf{u},\tilde{P}_{j})$ $\leq\left\|\sum_{i\in I_{j+}^{\prime}}c_{ji}(\mathbf{a}_{i}^{T},b_{i})-\sum_{k\in C_{j+}}\tilde{x}_{jk}\tilde{c}_{jk}(\tilde{\mathbf{a}}_{k}^{T},\tilde{b}_{k})\right\|$
$\leq\sum_{k\in C_{j+}}\left\|\sum_{i\in I_{jk}^{\prime}}c_{ji}(\mathbf{a}_{i}^{T},b_{i})-\tilde{x}_{jk}\tilde{c}_{jk}(\tilde{\mathbf{a}}_{k}^{T},\tilde{b}_{k})\right\|+\sum_{i\in N_{j+}}|c_{ji}|\|(\mathbf{a}_{i}^{T},b_{i})\|$
$\leq\sum_{k\in C_{j+}}\left\|\sum_{i\in I_{jk}^{\prime}}\left[c_{ji}(\mathbf{a}_{i}^{T},b_{i})-x_{ji}\tilde{c}_{jk}(\tilde{\mathbf{a}}_{k}^{T},\tilde{b}_{k})\right]\right\|+\sum_{i\in N_{j+}}|c_{ji}|\|(\mathbf{a}_{i}^{T},b_{i})\|$
$\leq\sum_{k\in C_{j+}}\sum_{i\in I_{jk}^{\prime}}\left\|c_{ji}(\mathbf{a}_{i}^{T},b_{i})-x_{ji}\tilde{c}_{jk}(\tilde{\mathbf{a}}_{k}^{T},\tilde{b}_{k})\right\|+\sum_{i\in N_{j+}}|c_{ji}|\|(\mathbf{a}_{i}^{T},b_{i})\|$
$\leq\sum_{k\in C_{j+}}\sum_{i\in I_{jk}^{\prime}}\min_{x}\Bigl{\|}c_{ji}(\mathbf{a}_{i}^{T},b_{i})-x\left(|I_{jk}|(c_{ji}(\mathbf{a}_{i}^{T},b_{i})+\epsilon_{ji})+l_{jk}\right)\Bigr{\|}+\sum_{i\in N_{j+}}|c_{ji}|\|(\mathbf{a}_{i}^{T},b_{i})\|$
$\leq\sum_{k\in C_{j+}}\sum_{i\in I_{jk}^{\prime}}\min_{x}\Bigl{\|}(1-x|I_{jk}|)c_{ji}(\mathbf{a}_{i}^{T},b_{i})-xl_{jk}-x|I_{jk}|\epsilon_{ji}\Bigr{\|}+\sum_{i\in N_{j+}}|c_{ji}|\|(\mathbf{a}_{i}^{T},b_{i})\|$
$\leq\sum_{k\in C_{j+}}\sum_{i\in I_{jk}^{\prime}}\min_{x}\Biggl{\{}|1-x|I_{jk}||\|c_{ji}(\mathbf{a}_{i}^{T},b_{i})\|+|x|I_{jk}||\left\|\frac{l_{jk}}{|I_{jk}|}+\epsilon_{ji}\right\|\Biggr{\}}+\sum_{i\in N_{j+}}|c_{ji}|\|(\mathbf{a}_{i}^{T},b_{i})\|$
$\leq\sum_{k\in C_{j+}}\sum_{i\in I_{jk}^{\prime}}\min\left\{\|c_{ji}(\mathbf{a}_{i}^{T},b_{i})\|,\left\|\frac{l_{jk}}{|I_{jk}|}+\epsilon_{ji}\right\|\right\}+\sum_{i\in N_{j+}}|c_{ji}|\|(\mathbf{a}_{i}^{T},b_{i})\|$
$\leq\sum_{k\in C_{j+}}\sum_{i\in I_{jk}^{\prime}}\min\left\{\|c_{ji}(\mathbf{a}_{i}^{T},b_{i})\|,\frac{\|l_{jk}\|}{|I_{jk}|}+\|\epsilon_{ji}\|\right\}+\sum_{i\in N_{j+}}|c_{ji}|\|(\mathbf{a}_{i}^{T},b_{i})\|$
The maximum value of the upper bound occurs when $I_{j+}^{\prime}=I_{j+}$. Thus, we have
$\max_{\mathbf{u}\in V_{P_{j}}}\operatorname{dist}(u,\tilde{P}_{j})\leq\sum_{k\in C_{j+}}\sum_{i\in I_{jk}}\min\left\{\|c_{ji}(\mathbf{a}_{i}^{T},b_{i})\|,\frac{\|l_{jk}\|}{|I_{jk}|}+\|\epsilon_{ji}\|\right\}+\sum_{i\in N_{j+}}|c_{ji}|\|(\mathbf{a}_{i}^{T},b_{i})\|$
To obtain a bound for $\max_{\mathbf{v}\in V_{\tilde{P}_{j}}}\operatorname{dist}(P_{j},\mathbf{v})$, we write $\mathbf{v}=\sum_{k\in C_{j+}^{\prime}}\tilde{c}_{jk}(\tilde{\mathbf{a}}_{k}^{T},\tilde{b}_{k})\in\tilde{P}_{j}$ and choose vertex $\mathbf{u}=\sum_{i\in I_{j+}^{\prime}}c_{ji}(\mathbf{a}_{i}^{T},b_{i})$, with $I_{j+}^{\prime}=\{i\in I_{j+}|i\in I_{k},k\in C_{j+}^{\prime}\}$. For this set we have $\tilde{x}_{jk}=1$, and thus this distance has already been taken into account in the calculation of $\max_{\mathbf{u}\in V_{P_{j}}}\operatorname{dist}(\mathbf{u},\tilde{P}_{j})$.
At last, we have
$H(P_{j},\tilde{P}_{j})\leq\sum_{k\in C_{j+}}\sum_{i\in I_{jk}}\min\left\{\|c_{ji}(\mathbf{a}_{i}^{T},b_{i})\|,\frac{\|l_{jk}\|}{|I_{jk}|}+\|\epsilon_{ji}\|\right\}+\sum_{i\in N_{j+}}|c_{ji}|\|(\mathbf{a}_{i}^{T},b_{i})\|$
Similarly, for $H(Q_{j},\tilde{Q}_{j})$ we have
$H(Q_{j},\tilde{Q}_{j})\\ \leq\sum_{k\in C_{j-}}\sum_{i\in I_{jk}}\min\left\{\|c_{ji}(\mathbf{a}_{i}^{T},b_{i})\|,\frac{\|l_{jk}\|}{|I_{jk}|}+\|\epsilon_{ji}\|\right\}\\ +\sum_{i\in N_{j-}}|c_{ji}|\|(\mathbf{a}_{i}^{T},b_{i})\|$
Combining, we obtain
$\frac{1}{\rho}\max_{\mathbf{x}\in B}|v_{j}(\mathbf{x})-\tilde{v}_{j}(\mathbf{x})|\leq\sum_{k=1}^{m}\sum_{i\in I_{jk}}\min\left\{\|c_{ji}(\mathbf{a}_{i}^{T},b_{i})\|,\frac{\|l_{jk}\|}{|I_{jk}|}+\|\epsilon_{ji}\|\right\}\\ +\sum_{i\in N_{j}}|c_{ji}|\|(\mathbf{a}_{i}^{T},b_{i})\|$
Using the fact that $I_{jk}\subset I_{k}$ and $N_{min}\leq|I_{jk}|,\forall j,k$ we have
$\frac{1}{\rho}\max_{\mathbf{x}\in B}|v_{j}(\mathbf{x})-\tilde{v}_{j}(\mathbf{x})|\leq\sum_{k=1}^{m}\sum_{i\in I_{k}}\min\left\{\|c_{ji}(\mathbf{a}_{i}^{T},b_{i})\|,\frac{\|l_{jk}\|}{N_{min}}+\|\epsilon_{ji}\|\right\}+\sum_{i\in N_{j}}|c_{ji}|\|(\mathbf{a}_{i}^{T},b_{i})\|$
We make use of the following inequality, which is a direct consequence of Cauchy-Schwartz Inequality
$\sum_{j=1}^{m}|u_{j}|\leq\sqrt{m}\sqrt{\sum_{j=1}^{m}|u_{j}|^{2}}=\sqrt{m}\|(u_{1},\dots,u_{m})\|$
We have
$\sum_{j=1}^{m}|c_{ji}|\leq\sqrt{m}\|\mathbf{C}_{:,i}\|$
$\sum_{j=1}^{m}\|l_{jk}\|\leq\sqrt{m}\sqrt{\sum_{j=1}^{m}\|l_{jk}\|^{2}}=\sqrt{m}\cdot l_{k}$
$\sum_{j=1}^{m}\|\epsilon_{ji}\|\leq\sqrt{m}\sqrt{\sum_{j=1}^{m}\|\epsilon_{ji}\|^{2}}=\sqrt{m}\|\epsilon_{:,i}\|_{F}$
Using the above inequalities, the fact that $\sum\min\leq\min\sum$, and the fact that $\max\sum\leq\sum\max$ we get
$\frac{1}{\rho}\max_{\mathbf{x}\in B}\|v(\mathbf{x})-\tilde{v}(\mathbf{x})\|_{1}$ $\leq\sum_{k=1}^{m}\sum_{i\in I_{k}}\min\left\{\sum_{j=1}^{m}|c_{ji}|\|(\mathbf{a}_{i}^{T},b_{i})\|,\frac{\sum_{j=1}^{m}\|l_{jk}\|}{N_{min}}+\sum_{j=1}^{m}\|\epsilon_{ji}\|\right\}$
$+\sum_{j=1}^{m}\sum_{i\in N_{j}}|c_{ji}|\|(\mathbf{a}_{i}^{T},b_{i})\|$
$\leq\sqrt{m}\sum_{i=1}^{n}\min\left\{\|\mathbf{C}_{:,i}\|\|(\mathbf{a}_{i}^{T},b_{i})\|,\frac{l_{k(i)}}{N_{min}}+\|\epsilon_{:,i}\|_{F}\right\}$
$+\sum_{j=1}^{m}\sum_{i\in N_{j}}|c_{ji}|\|(\mathbf{a}_{i}^{T},b_{i})\|$
as desired

### v2

@@
 Abstract
-We present TropNNC, a structured pruning framework for compressing neural networks with linear and convolutional layers and ReLU activations.
-Our approximation is based on a geometrical approach to machine/deep learning, using tropical geometry and extending the work of Misiakos et al.
+We present TropNNC, a framework for compressing neural networks with linear and convolutional layers and ReLU activations.
+TropNNC is a structured compression framework based on a geometrical approach to machine/deep learning, using tropical geometry and extending the work of Misiakos et al.
 (2022).
-We use the Hausdorff distance of zonotopes in its standard continuous form to achieve a tighter approximation bound for tropical polynomials compared to Misiakos et al.
-(2022).
-This enhancement allows for superior functional approximations of neural networks, leading to a more effective compression algorithm.
+We use the Hausdorff distance of zonotopes in its standard continuous form to achieve a tighter approximation bound for tropical polynomials compared to previous work.
+This enhancement leads to the development of an effective compression algorithm that achieves superior functional approximations of neural networks.
 Our method is significantly easier to implement compared to other frameworks, and does not depend on the availability of training data samples.
 We validate our framework through extensive empirical evaluations on the MNIST, CIFAR, and ImageNet datasets.
-Our results demonstrate that TropNNC achieves performance on par with the state-of-the-art method ThiNet, even surpassing it in compressing linear layers, and to the best of our knowledge, it is the first method that achieves this using tropical geometry.
+Our results demonstrate that TropNNC achieves performance on par with state-of-the-art methods like ThiNet (even surpassing it in compressing linear layers) and CUP.
+To the best of our knowledge, it is the first method that achieves this using tropical geometry.
 Keywords: tropical geometry, zonotopes, Hausdorff distance, neural network compression
@@
 This transformation turns polynomials into piecewise linear functions, making tropical geometry highly relevant to the study of neural networks with piecewise linear activations like ReLU.
-Tropical mathematics encompasses a wide range of applications, which include operations research and scheduling (Cuninghame-Green, 1979), control and optimization (Baccelli et al., 2001; Cohen et al., 2004; Butkovič, 2010; Akian et al., 2012), mathematical physics (Litvinov et al., 2001), speech recognition (Mohri et al., 2002), polyhedral geometry (Gaubert and Katz, 2011), tropical geometry (Maclagan and Sturmfels, 2021), dynamical systems on weighted lattices (Maragos, 2017), finite state transducers (Theodosis and Maragos, 2018), convex regression (Maragos and Theodosis, 2020), and machine learning (Maragos et al., 2021; Gärtner and Jaggi, 2008).
+Tropical mathematics encompasses a wide range of applications, which include operations research and scheduling (Cuninghame-Green, 1979), control and optimization (Baccelli et al., 2001; Cohen et al., 2004; Butkovič, 2010; Akian et al., 2012), mathematical physics (Litvinov et al., 2001), speech recognition (Mohri et al., 2002), polyhedral geometry (Gaubert and Katz, 2011), tropical geometry (Maclagan and Sturmfels, 2021), dynamical systems on weighted lattices (Maragos, 2017), sparsity theory on weighted lattices (Tsilivis et al., 2022), matrix factorization (Kordonis et al., 2024), finite state transducers (Theodosis and Maragos, 2018), convex regression (Maragos and Theodosis, 2020), and machine learning (Maragos et al., 2021; Gärtner and Jaggi, 2008).
 Recently, tropical geometry has been applied to the theoretical study of neural networks.
@@
 Our contributions include:
-  • Obtaining a tighter bound for the functional approximation of tropical rational mappings than that presented in (Misiakos et al., 2022) using the standard continuous Hausdorff distance.
+  • Obtaining a tighter bound for the functional approximation of tropical polynomials than that presented in (Misiakos et al., 2022) using the standard, continuous Hausdorff distance.
   • Designing an algorithm that leverages tropical geometry and the Hausdorff distance for the structured compression of neural networks.
@@
 Our framework leverages clustering and can be naturally extended to non-uniform pruning, drawing parallels to the CUP framework.
-In fact, it improves upon steps 1 and 3 of CUP.
+In fact, it improves upon steps 1, 2, and 3 of CUP.
 For step 1, we use similar clustering vectors (referred to as filter features by CUP authors Duggal et al.
 (2021)) for linear layers, and introduce a new method for constructing clustering vectors in convolutional layers.
+For step 2, we introduce two approaches for choosing the distance threshold of the hierarchical clustering.
 Our main contribution lies in step 3, where we employ tropical geometry to select a better representative for each cluster, rather than simply choosing the neuron or channel with the most energy.
@@
 In this work, we adhere to the convention of using the max-plus semiring $(\mathbb{R}_{\mathrm{max}},\vee,+)$, defined as the set $\mathbb{R}_{\mathrm{max}}=\mathbb{R}\cup\{-\infty\}$ equipped with two binary operations: $\vee$ (ordinary max) and $+$ (ordinary sum).
-3.1.1 Tropical Polynomials and Rational Mappings
+3.1.1 Tropical Polynomials and Rational Functions
 Within the max-plus semiring, we can define polynomials. A tropical polynomial $f$ in $d$ variables $\mathbf{x}=(x_{1},\ldots,x_{d})$ is defined as the function:
 $f(\mathbf{x})=\bigvee_{i=1}^{n}\left\{\mathbf{a}_{i}^{T}\mathbf{x}+b_{i}\right\}=\max_{i\in[n]}\left\{\mathbf{a}_{i}^{T}\mathbf{x}+b_{i}\right\},$ (1)
-where $[n]=\{1,...,n\}$.
+where $[n]:=\{1,...,n\}$.
 Here, $n$ represents the rank of the tropical polynomial.
@@
 Figures 1(a), 1(b) illustrate examples of tropical polynomials in one and multiple variables, respectively.
-Tropical rational mappings extend the concept of rational polynomial functions to tropical algebra.
-A tropical rational mapping or function is defined as the tropical multiplication of a tropical polynomial $p$ by the tropical multiplicative inverse $q^{-1}$ of another tropical polynomial $q$.
+Tropical rational functions extend the concept of rational polynomial functions to tropical algebra.
+A tropical rational function is defined as the tropical multiplication of a tropical polynomial $p$ by the tropical multiplicative inverse $q^{-1}$ of another tropical polynomial $q$.
 In conventional arithmetic, this operation corresponds to the difference of two tropical polynomials $p$ and $q$:
 $r(\mathbf{x})=p(\mathbf{x})-q(\mathbf{x})$
-Tropical rational mappings correspond to general piecewise linear functions.
-Specifically, every tropical rational mapping is a continuous piecewise linear function, and every continuous piecewise linear function can be expressed (though not uniquely) as a tropical rational mapping.
-Figures 1(c), 1(d) provide examples of tropical rational mappings.
+Tropical rational functions correspond to general piecewise linear functions.
+Specifically, every tropical rational function is a continuous piecewise linear function, and every continuous piecewise linear function can be expressed (though not uniquely) as a tropical rational function.
+Figures 1(c), 1(d) provide examples of tropical rational functions.
 (a) $f(x)=\max\{x,1\}$
@@
 (d) $f(x,y)=\max\{x,-x\}-\max\{y,-y\}$
-Figure 1: (a) depicts a single-variate tropical polynomial, (b) depicts a multi-variate tropical polynomial, (c) depicts a single-variate tropical rational mapping, (d) depicts a multi-variate tropical rational mapping
+Figure 1: (a) depicts a single-variate tropical polynomial, (b) depicts a multi-variate tropical polynomial, (c) depicts a single-variate tropical rational function, (d) depicts a multi-variate tropical rational function
 3.1.2 Newton Polytopes
@@
 If $(\mathbf{a}^{T},b)$ lies below the upper envelope of $P$, then $\forall\mathbf{x}\in\mathbb{R}^{d},\mathbf{a}^{T}\mathbf{x}+b\leq p(\mathbf{x})$.
-The inequality is strict if $(\mathbf{a}_{i}^{T},b_{i})$ lies strictly below the upper envelope.
+The inequality is strict if $(\mathbf{a}^{T},b)$ lies strictly below the upper envelope.
 An immediate consequence of the above lemma is the following theorem:
@@
 $(f\vee g)(x,y)=\max\{x+1,y+1,-y+1,-x+1\},$
-which correspond to the vertices of the upper envelope of $\mathrm{ENewt}(f\vee g)$.
+which corresponds to the vertices of the upper envelope of $\mathrm{ENewt}(f\vee g)$.
 (a) $\mathrm{ENewt}(f)$
@@
 $\mathbf{u}=\mathbf{s}+\sum_{i\in I}\mathbf{v}_{i},$
-where $I\subset[n]$.
-It can be shown that zonotopes have up to exponential many verticesIn fact, for generators in general positions it has exactly $2\sum_{j=0}^{d-1}nCr(n-1,j)$ vertices (Gritzmann and Sturmfels, 1993).
+where $I\subseteq[n]$.
+It can be shown that zonotopes have up to exponential many verticesIn fact, for generators in general positions they have exactly $2\sum_{j=0}^{d-1}nCr(n-1,j)$ vertices (Gritzmann and Sturmfels, 1993).
 Zonotopes exhibit several interesting properties.
@@
 Thus, $p_{j}$ and $q_{j}$ are tropical polynomials formed by the addition of tropical polynomials.
-Consequently, $v_{j}$ is a tropical rational mapping.
+Consequently, $v_{j}$ is a tropical rational function.
 We call $p_{j}$ the positive polynomial and $q_{j}$ the negative polynomial of $v_{j}$.
@@
 Theorem 6 (Zhang et al., 2018)
-A ReLU-activated deep neural network $\mathbf{F}:\mathbb{R}^{d}\rightarrow\mathbb{R}^{m}$ is a tropical rational mapping.
+A ReLU-activated deep neural network $\mathbf{F}:\mathbb{R}^{d}\rightarrow\mathbb{R}^{m}$ is a tropical rational mapping (vector whose elements are tropical rational functions).
 The extended Newton polytope of $|c_{ji}|f_{i}$ is an edge with one endpoint at the origin $\mathbf{0}$ and the other at $|c_{ji}|(\mathbf{a}_{i}^{T},b_{i})$.
@@
 Misiakos et al.
-(2022) derived a bound for the error in approximating the tropical polynomials that represent a neural network.
+(2022) derived a bound for the error of approximating the tropical polynomials that represent a neural network.
 The motivation for geometrically bounding the error of these polynomials stems from Theorem 5.
@@
 where $B=\{\mathbf{x}\in\mathbb{R}^{d}:\|\mathbf{x}\|\leq r\}$ and $\rho=\sqrt{r^{2}+1}$.
-Since $V_{P}\subset P,V_{\tilde{P}}\subset\tilde{P}$, we have that $\mathrm{dist}(\mathbf{u},\tilde{P})\leq\mathrm{dist}(\mathbf{u},V_{\tilde{P}}),\mathrm{dist}(P,\mathbf{v})\leq\mathrm{dist}(V_{P},\mathbf{v})$.
+Since $V_{P}\subseteq P,V_{\tilde{P}}\subseteq\tilde{P}$, we have that $\mathrm{dist}(\mathbf{u},\tilde{P})\leq\mathrm{dist}(\mathbf{u},V_{\tilde{P}}),\mathrm{dist}(P,\mathbf{v})\leq\mathrm{dist}(V_{P},\mathbf{v})$.
 The equality in the aforementioned inequalities is achieved only in special cases and thus the bound $H(P,\tilde{P})$ we provide is in general tighter than $DH(P,\tilde{P})$ from Misiakos et al.
@@
 On the other hand, according to the Hausdorff distance in its standard form, taking the sum is the optimal strategy and gives distance $1/2$ (example 2), better than $\sqrt{3}/2$ (example 3).
-This implies that an algorithm based on the Hausdorff distance would imply a better approximation of $\tilde{P}$ in the above example.
+This implies that an algorithm based on the Hausdorff distance would imply a better approximation of $P$ in the above example.
 We derive the following result, which differs from (Misiakos et al., 2022) by using the standard form of the Hausdorff distance and thus providing a tighter bound.
-For the proof we apply triangle inequality and then repeatedly Proposition 10 to each output of the network.
+For the proof we apply the triangle inequality and then repeatedly Proposition 10 to each output of the network.
 The full proofs of Proposition 10 and Theorem 11 are provided in Appendix A.3.
@@
 First, we focus on single output networks, as the one depicted in Figure 3, with $m=1$.
-The network has weight matrices $\mathbf{A}\in\mathbb{R}^{n\times d}$, $\mathbf{C}\in\mathbb{R}^{1\times n}$, and a single output $v$ with corresponding positive and negative zonotopes $P$ and $Q$.
+The network has weight matrices $\mathbf{A}\in\mathbb{R}^{n\times(d+1)}$, $\mathbf{C}\in\mathbb{R}^{1\times n}$, and a single output $v$ with corresponding positive and negative zonotopes $P$ and $Q$.
 We aim to approximate this network with a smaller one that has fewer neurons in the hidden layer, say $K<n$, resulting in a new hidden layer $\tilde{\mathbf{f}}=(\tilde{f}_{1},\ldots,\tilde{f}_{K})$ with new weight matrices $\widetilde{\mathbf{A}}$ and $\widetilde{\mathbf{C}}$.
@@
 (c) TropNNC
-Figure 5: Example execution of Zonotope K-means and Improved Zonotope K-means
+Figure 5: Example execution of Zonotope K-means and TropNNC
 The result of Example 4 can be generalized. Indeed, the following bound and its corollary hold (proofs in Appendix A.4 and A.5).
@@
 However, in our experiments, this approach yielded worse results.
-Another variation involves clustering neurons based on cosine similarity rather than the L2 distance of the vectors, recognizing that neurons with parallel input weights can be clustered together regardless of their output weights.
+Another variation involves clustering neurons based on cosine similarity rather than the $L^{2}$ distance of the vectors, recognizing that neurons with parallel input weights can be clustered together regardless of their output weights.
 Additionally, one could normalize the vectors based on their input weights by shifting the norm to the output weights before applying K-means.
@@
 Thus, for every cluster $k$, we have unknowns $(\tilde{\mathbf{a}}_{k}^{T},\tilde{b}_{k},\widetilde{\mathbf{C}}_{:,k}^{T})=(\tilde{a}_{k1},\ldots,\tilde{a}_{kd},\tilde{b}_{k},\widetilde{C}_{1k},\ldots)$, and we wish to find a solution which minimizes the following criterion:
-$\sum_{j=1}^{m}\left[\tilde{C}_{jk}(\tilde{\mathbf{a}}_{k}^{T},\tilde{b}_{k})-\sum_{i\in I_{k}}C_{ji}(\mathbf{a}_{i}^{T},b_{i})\right]^{2},$
+$\sum_{j=1}^{m}\left\|\tilde{C}_{jk}(\tilde{\mathbf{a}}_{k}^{T},\tilde{b}_{k})-\sum_{i\in I_{k}}C_{ji}(\mathbf{a}_{i}^{T},b_{i})\right\|^{2},$
 where $m$ is the number of outputs, and $I_{k}$ is the set of neurons of cluster $k$.
@@
   • $\|\cdot\|_{F}$ is the Frobenius norm.
-6 Discussion and Limitations
+6 Extensions, Limitations, and Applications of the Method
 Before proceeding to our experiments, we would like to discuss the applicability of our algorithm, highlight some limitations, and explore potential strategies to overcome these challenges.
@@
 We can incorporate these computations into the weights and bias of the convolutional or linear layer.
+Using these ideas, we can fuse a batch normalization layer with its preceding convolutional or linear layer.
+This enables two approaches for applying our algorithm to networks with batch normalization.
+The first approach involves fusing all layers of the network beforehand, resulting in a fully fused network without batch normalization.
+The second approach performs fusion during the compression process, applying it layer-by-layer as each layer is compressed.
+For the second approach, we made a notable observation during our experiments, although the reasoning behind it remains unclear: using the pre-fusion weights/kernels of the layers for clustering vectors consistently led to significantly better results.
 As for other limitations, we should note that our algorithm is designed to work strictly on layers as described above.
@@
 Layers that apply computations, like batch normalization layers or the loops of residual blocks of ResNet networks fundamentally change the theory for the analysis of the network.
-For example, similary to ThiNet, our methods can not be applied for the compression of the terminal convolutional layer of residual blocks.
-Non-uniform pruning.
+For example, similarly to ThiNet, our methods can not be applied for the compression of the terminal convolutional layer of residual blocks.
+Non-uniform compression.
 The techniques we have presented apply a uniform pruning ratio to all layers of the network.
 However, they can easily be extended to non-uniform pruning if instead of K-means we use hierarchical clustering with a global threshold parameter, like in the CUP framework (Duggal et al., 2021).
-In fact, when used with hierarchical clustering, TropNNC differs from CUP in step 1, where we choose a different approach to build the cluster vectors (or filter features as Duggal et al.
+In fact, when used with hierarchical clustering, TropNNC differs from CUP in step 1, where we choose a different approach to build the clustering vectors (or filter features as Duggal et al.
 (2021) call them) of a convolutional layer, and more importantly in step 3, where we choose a different cluster representative based on tropical geometry.
+We also propose a modification to step 2 of CUP, specifically the hierarchical clustering step using a global distance threshold.
+Since the clustering vectors of different layers have varying dimensions, and vectors in higher-dimensional spaces tend to be more spread out, we introduce two variants for selecting the distance threshold for each layer:
+  • Variant 1: For each layer, take the distance threshold to be some global constant times the square root of the dimension of the clustering vectors of the layer.
+  • Variant 2: For each layer, take the distance threshold to be some global constant times the mean of the norms of the clustering vectors of the layer.
+As demonstrated in our experiments, the first variant excels when minimizing the number of parameters, while the second variant is more effective when our objective is to minimize floating-point operation demands.
 7 Experiments
@@
 Baselines.
-We compare our uniform pruning framework with state-of-the-art methods that conduct structured pruning without requiring re-training (or fine-tuning).
+We compare our framework with state-of-the-art methods that conduct structured pruning without requiring re-training (or fine-tuning).
 Specifically, we compare it with the algorithms proposed by Misiakos et al.
@@
 For the non-uniform pruning variant of our framework, we compare it with CUP.
-CUP assigns to each neuron/channel a feature, performs hierarchical clustering based on these features, and replaces each neuron in each cluster by a cluster representative according to a maximum norm criterion.
-To ensure a fair comparison, we use identical cluster vectors for both frameworks, as defined in CUP, to emphasize the advantages of our tropical geometry-based approach for selecting cluster representatives.
+CUP assigns to each neuron/channel a feature, performs hierarchical clustering based on these features, and replaces each neuron/channel in each cluster by a cluster representative according to a maximum norm criterion.
+As explained, our proposed algorithm enhances all three steps of CUP.
+Through an ablation study, we found that each improvement is essential for achieving a competitive advantage over CUP, indicating that these enhancements complement each other.
+In the presented experiments, we compare CUP exclusively with the fully enhanced version of TropNNC, incorporating all improvements.
 Furthermore, the trivial baseline of random structured pruning discards neurons or channels based on a uniform probability distribution, while L1 structured pruning targets those with the smallest L1 norm of their weights or kernels.
@@
 We evaluate our framework on the MNIST and CIFAR datasets, testing it across various models including simple multi-layer perceptrons (MLPs), convolutional neural networks (CNNs), LeNet (LeCun et al., 1998), AlexNet (Krizhevsky et al., 2012), and VGG (Simonyan and Zisserman, 2015).
-The non-uniform variant of our algorithm is tested on the CIFAR and ImageNet datasets, with its performance evaluated across models such as VGG, ResNet56-bn, and ResNet18 (He et al., 2016).
+The non-uniform variant of our algorithm is tested on the CIFAR and ImageNet datasets, with its performance evaluated across models such as VGG, ResNet56, and ResNet18 (He et al., 2016).
 7.1 MNIST Dataset, Pairs 3-5 and 4-9
@@
 As shown in the results, for both datasets, our algorithm outperforms Neural Path K-means.
-To evaluate the performance of our algorithm in compressing linear layers of deeper networks, we applied TropNNC to ”deepNN,” a fully connected neural network with layer sizes $28\times 28,512,256,128$, and $10$.
+To evaluate the performance of our algorithm in compressing linear layers of deeper networks, we applied TropNNC to ”deepNN”, a fully connected neural network with layer sizes $28\times 28,512,256,128$, and $10$.
 The performance plots are provided in Figures 7(a) and 7(b).
 As illustrated, TropNNC outperforms both Neural Path K-means from (Misiakos et al., 2022) and ThiNet from (Luo et al., 2017).
-To assess the performance of our algorithm in compressing convolutional layers, we applied TropNNC to ”deepCNN2D,” a LeNet-type convolutional neural network with ReLU activations.
+To assess the performance of our algorithm in compressing convolutional layers, we applied TropNNC to ”deepCNN2D”, a LeNet-type convolutional neural network with ReLU activations.
 The performance plots are provided in Figures 7(c) and 7(d).
@@
 Additionally, Figures 8(c) and 8(d) show the compression of VGG’s convolutional layers for these datasets.
+For VGG, we applied compression to the fused networks without batch normalization, as we found that ThiNet’s performance on the original networks was practically identical to its performance on the fused networks.
 The results indicate that for larger datasets, TropNNC consistently outperforms Neural Path K-means from (Misiakos et al., 2022).
@@
 For the compression of convolutional layers of VGG, TropNNC matches ThiNet.
-For the compression of convolutional layers of VGG, we do not present results for iterative TropNNC due to the excessively long execution time required.
+For the compression of convolutional layers of VGG, we do not present results for iterative TropNNC because it had almost identical performance to the non-iterative algorithm.
 (a) AlexNet, linear, CIFAR10
@@
 (d) VGG, conv., CIFAR100
-Figure 8: Compression of linear and convolutional layers of AlexNet on CIFAR datasets.
+Figure 8: Compression of linear layers of AlexNet and convolutional layers of VGG on CIFAR datasets.
 7.4 Non-ReLU Activation
@@
 The results are presented in Figures 9(a), 9(b), 9(c) and 9(d).
-Interestingly, our algorithm demonstrate robust performance even on networks with non-ReLU activations.
+Interestingly, our algorithm demonstrates robust performance even on networks with non-ReLU activations.
 (a) LeNet, linear, MNIST
@@
 7.5 Non-uniform Pruning
-In this section, we evaluate the effectiveness of the non-uniform pruning variant of TropNNC.
-Our experiments here compress various models, such as VGG on CIFAR-10, ResNet56-bn on CIFAR-10, and ResNet18 on ImageNet.
-The results are summarized in Tables 4, 5, and 6.
+In this section, we evaluate the effectiveness of the non-uniform variant of TropNNC.
+Our experiments here compress various models, such as VGG on CIFAR-10, ResNet56 on CIFAR-10, and ResNet18 on ImageNet.
+The results are summarized in Tables 4, 5, 6, and 7.
 We should emphasize that we did not apply any form of fine-tuning or re-training.
-Our findings indicate that the method we use for selecting cluster representatives demonstrates a clear advantage.
+In Table 4 we present the compression of fused models, i.e.
+models whose batch normalization has been fused into its preceding convolutional or linear layer, as explained in a previous section.
+We noticed that the performance of CUP was better on the original networks with batch normalization.
+In Tables 5, 6 and 7 we compare our performance with the performance of CUP compressing the original networks.
+For our method only, we applied batch normalization fusion following the second approach described earlier—fusing the batch normalization layer with the preceding layer on a layer-by-layer basis, followed by compression of the fused layer before moving on with the next layer.
+We observed that our method performed significantly better when the clustering vectors were generated using the pre-fusion weights/kernels of the layers.
+While the reason for this improvement remains unclear, it consistently led to better performance in our experiments.
+Our findings indicate that our method demonstrates a clear advantage.
 Notably, our approach shows a significant performance improvement for the VGG model, while the benefits are comparatively modest for ResNet models.
@@
 Further investigation may be required to fully understand these discrepancies and to optimize our approach across different model types.
+We also found that the first variant of non-uniform TropNNC excelled at reducing the overall network size but was less effective at minimizing inference operations.
+In contrast, the second variant performed well in both tasks, outperforming CUP.
+Upon further analysis, we concluded that the first variant tends to focus more aggressively on the final layers, where parameter count is high due to the increased number of channels, but the number of operations is lower because of smaller image sizes.
+Meanwhile, the second variant, like CUP, also targets the initial layers, where fewer parameters are present, but a greater number of operations is required.
+Model Method Threshold #params $\downarrow$ FLOPS $\downarrow$ Acc. $\uparrow$
+VGG Original - 14.72M 0.63G 93.64
+CIFAR10 CUP 0.15 4.70M 0.40G 56.22
+TropNNC (v2) 1.12 4.70M 0.33G 90.82
+ResNet56 Original - 0.85M 0.25G 93.67
+CIFAR10 CUP 0.55 0.79M 0.18G 45.89
+TropNNC (v2) 1.2 0.65M 0.18G 76.50
+ResNet18 Original - 11.68M 3.63G 69.10
+ImageNet CUP 0.6 11.62M 3.50G 58.20
+TropNNC (v2) 1 11.60M 3.47G 62.60
+Table 4: Comparison of CUP and TropNNC (variant 2) accuracy across different pruning thresholds for fused models without batch normalization.
 Method Threshold #params $\downarrow$ FLOPS $\downarrow$ Acc. $\uparrow$
 Original - 14.7M 0.63G 93.64
-CUP 0.15 6.24M 0.46G 92.34
-TropNNC 0.17 5.90M 0.45G 93.27
-CUP 0.20 4.62M 0.42G 67.82
-TropNNC 0.22 4.58M 0.41G 92.91
-CUP 0.25 3.61M 0.39G 18.10
-TropNNC 0.30 3.42M 0.38G 79.84
-Table 4: Comparison of CUP and TropNNC accuracy across different pruning thresholds on CIFAR10, VGG.
+CUP 0.15 6.24M 0.46G 92.41
+TropNNC (v1) 0.006 5.70M 0.44G 93.58
+CUP 0.20 4.62M 0.42G 67.91
+TropNNC (v1) 0.0085 4.32M 0.41G 93.46
+CUP 0.25 3.61M 0.39G 16.02
+TropNNC (v1) 0.011 3.53M 0.39G 93.21
+Table 5: Comparison of CUP and TropNNC (variant 1) accuracy across different pruning thresholds on CIFAR10, VGG.
 Method Threshold #params $\downarrow$ FLOPS $\downarrow$ Acc. $\uparrow$
 Original - 0.85M 0.25G 93.67
-CUP 0.50 0.83M 0.22G 87.10
-TropNNC-3it. 0.50 0.83M 0.22G 88.34
-CUP 0.55 0.82M 0.21G 82.18
-TropNNC-3it. 0.55 0.82M 0.21G 83.32
-CUP 0.60 0.79M 0.20G 59.97
-TropNNC-3it. 0.60 0.79M 0.20G 64.16
-Table 5: Comparison of CUP and TropNNC (3 iterations) accuracy across different pruning thresholds on CIFAR10, ResNet56-bn.
+CUP 0.55 0.72M 0.20G 87.32
+TropNNC (v2) 1.17 0.71M 0.19G 90.53
+CUP 0.6 0.63M 0.17G 78.83
+TropNNC (v2) 1.23 0.63M 0.17G 87.04
+CUP 0.65 0.54M 0.15G 71.48
+TropNNC (v2) 1.3 0.54M 0.15G 73.03
+TropNNC (v1) 0.067 0.49M 0.19G 82.70
+Table 6: Comparison of CUP and TropNNC accuracy across different pruning thresholds on CIFAR10, ResNet56.
 Method Threshold #params $\downarrow$ FLOPS $\downarrow$ Acc. $\uparrow$
-Original - 11.68M 3.63G 69.10
-CUP 0.60 11.65M 3.55G 62.30
-TropNNC 0.60 11.65M 3.55G 64.80
-CUP 0.65 11.60M 3.48G 54.10
-TropNNC 0.65 11.60M 3.48G 56.60
-CUP 0.70 11.43M 3.32G 29.20
-TropNNC 0.70 11.43M 3.32G 33.20
+Original - 11.69M 3.64G 70.10
+CUP 0.5 11.66M 3.58G 66.60
+TropNNC (v2) 1.1 11.64M 3.44G 68.40
+CUP 0.6 11.49M 3.38G 53.20
+TropNNC (v2) 1.2 11.27M 3.07G 57.90
+CUP 0.65 11.17M 3.15G 29.20
+TropNNC (v2) 1.25 10.71M 2.75G 41.20
+TropNNC (v1) 0.021 9.87M 3.46G 58.10
 Table 7: Comparison of CUP and TropNNC accuracy across different pruning thresholds on ImageNet, ResNet18.
@@
 Leveraging this enhancement we refined the choice of compressed node’s weights that correspond to the K-means clusters from (Misiakos et al., 2022).
-Our novel algorithm, TropNNC, applied to linear and convolutional layers of deep neural networks outperforms Neural Path K-means Misiakos et al.
+Our novel algorithm, TropNNC, applied to linear and convolutional layers of deep neural networks outperforms Neural Path K-means of Misiakos et al.
 (2022), and manages to match or even surpass the performance of ThiNet, particularly for the compression of linear layers.
@@
 $\mathbf{a}^{T}\mathbf{x}+b\leq\mathbf{a}^{T}\mathbf{x}+b^{\prime}=\langle(\mathbf{a}^{T},b^{\prime}),(\mathbf{x},1)\rangle$
-$=\sum_{i=1}^{k}\langle\mathbf{v}_{i},(\mathbf{x},1)\rangle\leq\max_{i=1,\ldots,k}\langle\mathbf{v}_{i},(\mathbf{x},1)\rangle\leq p(\mathbf{x})$
+$=\sum_{i=1}^{k}\lambda_{i}\langle\mathbf{v}_{i},(\mathbf{x},1)\rangle\leq\max_{i=1,\ldots,k}\langle\mathbf{v}_{i},(\mathbf{x},1)\rangle\leq p(\mathbf{x}).$
 If $(\mathbf{a}_{i}^{T},b_{i})$ lies strictly below the upper envelope of $P$, then $b^{\prime}>b$, and the inequality is strict.
@@
 The polytope $P$, the domain of the supremum, is convex and compact. Thus, it suffices to prove that the function
-$f(\mathbf{u})=\operatorname{dist}(\mathbf{u},\tilde{P})=\min_{\tilde{\mathbf{u}}\in\tilde{P}}\|\mathbf{u}-\tilde{\mathbf{u}}\|$
+$f(\mathbf{u})=\operatorname{dist}(\mathbf{u},\tilde{P})=\inf_{\tilde{\mathbf{u}}\in\tilde{P}}\|\mathbf{u}-\tilde{\mathbf{u}}\|$
 is convex in terms of $\mathbf{u}$.
@@
 $=\sum_{j=1}^{m}\lvert(p_{j}(\mathbf{x})-\tilde{p}_{j}(\mathbf{x}))-(q_{j}(\mathbf{x})-\tilde{q}_{j}(\mathbf{x}))\rvert$
-$\leq\sum_{j=1}^{m}\left(\lvert p_{j}(\mathbf{x})-\tilde{p}_{j}(\mathbf{x})\rvert+\lvert q_{j}(\mathbf{x})-\tilde{q}_{j}(\mathbf{x})\rvert\right)$
+$\leq\sum_{j=1}^{m}\left(\lvert p_{j}(\mathbf{x})-\tilde{p}_{j}(\mathbf{x})\rvert+\lvert q_{j}(\mathbf{x})-\tilde{q}_{j}(\mathbf{x})\rvert\right).$
 By Proposition 10 we derive
-$\frac{1}{\rho}\max_{\mathbf{x}\in B}\|\mathbf{v}(\mathbf{x})-\tilde{\mathbf{v}}(\mathbf{x})\|_{1}\leq\sum_{j=1}^{m}\left(H(P_{j},\tilde{P}_{j})+H(Q_{j},\tilde{Q}_{j})\right),$
+$\frac{1}{\rho}\max_{\mathbf{x}\in B}\|\mathbf{v}(\mathbf{x})-\tilde{\mathbf{v}}(\mathbf{x})\|_{1}\leq\sum_{j=1}^{m}\left(H(P_{j},\tilde{P}_{j})+H(Q_{j},\tilde{Q}_{j})\right).$
 A.4 Proof of Proposition 12
 Proof  For the output function, it holds that
-$v(\mathbf{x})=p(\mathbf{x})-q(\mathbf{x}),\quad\tilde{v}(\mathbf{x})=\tilde{p}(\mathbf{x})-\tilde{q}(\mathbf{x})$
+$v(\mathbf{x})=p(\mathbf{x})-q(\mathbf{x}),\quad\tilde{v}(\mathbf{x})=\tilde{p}(\mathbf{x})-\tilde{q}(\mathbf{x}).$
 From the triangle inequality, we deduce
-$\frac{1}{\rho}\left|v(\mathbf{x})-\tilde{v}(\mathbf{x})\right|\leq H(P,\tilde{P})+H(Q,\tilde{Q})$
+$\frac{1}{\rho}\left|v(\mathbf{x})-\tilde{v}(\mathbf{x})\right|\leq H(P,\tilde{P})+H(Q,\tilde{Q}).$
 Thus, it suffices to get a bound on $H(P,\tilde{P})$ and $H(Q,\tilde{Q})$.
-Let $I_{+}\subset[n]$ and $I_{-}\subset[n]$ be the sets of positive and negative generators, respectively.
+Let $I_{+}\subseteq[n]$ and $I_{-}\subseteq[n]$ be the sets of positive and negative generators, respectively.
 First, we deal with $H(P,\tilde{P})$.
@@
 $\operatorname{dist}(P,\mathbf{v})=0,\quad\forall\mathbf{v}\in V_{\tilde{P}}.$
-Let $I_{k}\subset[n]$ be the set of generators that belong to cluster $k$.
-Let $C_{+}\subset[K]$ be the set of clusters of positive generators, and $C_{-}\subset[K]$ the set of clusters of negative generators.
-Consider any vertex $\mathbf{u}$ of $P$. This vertex can be written as the sum of generators $c_{i}(\mathbf{a}_{i}^{T},b_{i})$, for some subset $I_{+}^{\prime}\subset I_{+}$. Thus,
-$\mathbf{u}=\sum_{i\in I_{+}^{\prime}}c_{i}(\mathbf{a}_{i}^{T},b_{i})$
-For every generator $c_{i}(\mathbf{a}_{i}^{T},b_{i})$ that belongs to cluster $k$, define
-$x_{i}=\arg\min_{x}\left\{c_{i}(\mathbf{a}_{i}^{T},b_{i})-x\tilde{c}_{k}(\tilde{\mathbf{a}}_{k}^{T},\tilde{b}_{k})\right\}$
+Let $I_{k}\subseteq[n]$ be the set of generators that belong to cluster $k$.
+Let $C_{+}\subseteq[K]$ be the set of clusters of positive generators, and $C_{-}\subseteq[K]$ the set of clusters of negative generators.
+Consider any vertex $\mathbf{u}$ of $P$. This vertex can be written as the sum of generators $c_{i}(\mathbf{a}_{i}^{T},b_{i})$, for some subset $I_{+}^{\prime}\subseteq I_{+}$. Thus,
+$\mathbf{u}=\sum_{i\in I_{+}^{\prime}}c_{i}(\mathbf{a}_{i}^{T},b_{i}).$
+For every positive generator $c_{i}(\mathbf{a}_{i}^{T},b_{i})$ that belongs to cluster $k$, define
+$x_{i}=\arg\min_{x}\left\|c_{i}(\mathbf{a}_{i}^{T},b_{i})-x\tilde{c}_{k}(\tilde{\mathbf{a}}_{k}^{T},\tilde{b}_{k})\right\|,$
 i.e., project the generator onto its cluster representative.
 Since every inner-cluster pair of generators forms an acute angle, any generator and its cluster representative (sum of generators of the cluster) will also form an acute angle, and thus $x_{i}\geq 0$.
-Define $I_{k}^{\prime}=I_{k}\cap I_{+}^{\prime}$. For every cluster $k$, define
-$\tilde{x}_{k}=\sum_{i\in I_{k}^{\prime}}x_{i}$
+For every cluster $k\in C_{+}$ define $I_{k}^{\prime}=I_{k}\cap I_{+}^{\prime}$ and
+$\tilde{x}_{k}=\sum_{i\in I_{k}^{\prime}}x_{i}.$
 Since the cluster representative is the sum of the generators of the cluster, we have
 $\sum_{i\in I_{k}}x_{i}=1\Rightarrow\tilde{x}_{k}=\sum_{i\in I_{k}^{\prime}}x_{i}\leq 1$
-Thus, for every cluster $k$, the point $\tilde{x}_{k}\tilde{c}_{k}(\tilde{\mathbf{a}}_{k}^{T},\tilde{b}_{k}^{T})$ lies inside the segment $[\mathbf{0},\tilde{c}_{k}(\tilde{\mathbf{a}}_{k}^{T},\tilde{b}_{k}^{T})]$ and thus belongs to $\tilde{P}$.
-For the vertex $\mathbf{u}$, we choose to compare it with the vertex
-$\sum_{k\in C_{+}}\tilde{x}_{k}\tilde{c}_{k}(\tilde{\mathbf{a}}_{k}^{T},\tilde{b}_{k}^{T})\in\tilde{P}.$
+Thus, for every cluster $k\in C_{+}$, the point $\tilde{x}_{k}\tilde{c}_{k}(\tilde{\mathbf{a}}_{k}^{T},\tilde{b}_{k})$ lies inside the segment $[\mathbf{0},\tilde{c}_{k}(\tilde{\mathbf{a}}_{k}^{T},\tilde{b}_{k})]$ and thus belongs to $\tilde{P}$.
+For the vertex $\mathbf{u}$, we choose to compare it with the point
+$\sum_{k\in C_{+}}\tilde{x}_{k}\tilde{c}_{k}(\tilde{\mathbf{a}}_{k}^{T},\tilde{b}_{k})\in\tilde{P}.$
 Thus, we have that
@@
 $\leq\sum_{k\in C_{+}}\sum_{i\in I_{k}^{\prime}}\left\|c_{i}(\mathbf{a}_{i}^{T},b_{i})-x_{i}\tilde{c}_{k}(\tilde{\mathbf{a}}_{k}^{T},\tilde{b}_{k})\right\|$
-$=\sum_{k\in C_{+}}\sum_{i\in I_{k}^{\prime}}\left\|c_{i}(\mathbf{a}_{i}^{T},b_{i})-x_{i}|I_{k}|(c_{i}(\mathbf{a}_{i}^{T},b_{i})+\epsilon_{i})\right\|$
+$=\sum_{k\in C_{+}}\sum_{i\in I_{k}^{\prime}}\min_{x}\left\|c_{i}(\mathbf{a}_{i}^{T},b_{i})-x\tilde{c}_{k}(\tilde{\mathbf{a}}_{k}^{T},\tilde{b}_{k})\right\|$
 $=\sum_{k\in C_{+}}\sum_{i\in I_{k}^{\prime}}\min_{x}\Bigl{\|}c_{i}(\mathbf{a}_{i}^{T},b_{i})-x|I_{k}|(c_{i}(\mathbf{a}_{i}^{T},b_{i})+\epsilon_{i})\Bigr{\|}$
-$=\sum_{k\in C_{+}}\sum_{i\in I_{k}^{\prime}}\min_{x}\left\{|1-x|I_{k}||\cdot\left\|c_{i}(\mathbf{a}_{i}^{T},b_{i})\right\|+|x|I_{k}||\cdot\left\|\epsilon_{i}\right\|\right\}$
+$\leq\sum_{k\in C_{+}}\sum_{i\in I_{k}^{\prime}}\min_{x}\left\{|1-x|I_{k}||\cdot\left\|c_{i}(\mathbf{a}_{i}^{T},b_{i})\right\|+|x|I_{k}||\cdot\left\|\epsilon_{i}\right\|\right\}$
 $=\sum_{k\in C_{+}}\sum_{i\in I_{k}^{\prime}}\min\left\{\left\|c_{i}(\mathbf{a}_{i}^{T},b_{i})\right\|,\left\|\epsilon_{i}\right\|\right\}$
 $\leq\sum_{k\in C_{+}}\sum_{i\in I_{k}^{\prime}}\min\left\{\left\|c_{i}(\mathbf{a}_{i}^{T},b_{i})\right\|,\delta_{\max}\right\}$
-$=\sum_{i\in I_{+}^{\prime}}\min\left\{\left\|c_{i}(\mathbf{a}_{i}^{T},b_{i})\right\|,\delta_{\max}\right\}$
+$=\sum_{i\in I_{+}^{\prime}}\min\left\{\left\|c_{i}(\mathbf{a}_{i}^{T},b_{i})\right\|,\delta_{\max}\right\},$
 where $\epsilon_{i}$ is the error between generator $i$ and the cluster center/mean of K-means.
 The maximum value of the upper bound occurs when $I_{+}^{\prime}=I_{+}$. Thus,
-$\max_{\mathbf{u}\in V_{P}}\operatorname{dist}(\mathbf{u},\tilde{P})\leq\sum_{i\in I_{+}}\min\left\{\left\|c_{i}(\mathbf{a}_{i}^{T},b_{i})\right\|,\delta_{\max}\right\}$
+$\max_{\mathbf{u}\in V_{P}}\operatorname{dist}(\mathbf{u},\tilde{P})\leq\sum_{i\in I_{+}}\min\left\{\left\|c_{i}(\mathbf{a}_{i}^{T},b_{i})\right\|,\delta_{\max}\right\}.$
 Finally, we have
@@
 $\leq\max\left\{\sum_{i\in I_{+}}\min\left\{\left\|c_{i}(\mathbf{a}_{i}^{T},b_{i})\right\|,\delta_{\max}\right\},0\right\}$
-$=\sum_{i\in I_{+}}\min\left\{\left\|c_{i}(\mathbf{a}_{i}^{T},b_{i})\right\|,\delta_{\max}\right\}$
+$=\sum_{i\in I_{+}}\min\left\{\left\|c_{i}(\mathbf{a}_{i}^{T},b_{i})\right\|,\delta_{\max}\right\}.$
 Similarly, for $H(Q,\tilde{Q})$, we have
-$H(Q,\tilde{Q})\leq\sum_{i\in I_{-}}\min\left\{\left\|c_{i}(\mathbf{a}_{i}^{T},b_{i})\right\|,\delta_{\max}\right\}$
+$H(Q,\tilde{Q})\leq\sum_{i\in I_{-}}\min\left\{\left\|c_{i}(\mathbf{a}_{i}^{T},b_{i})\right\|,\delta_{\max}\right\}.$
 Combining, we get
-$\frac{1}{\rho}\left|v(\mathbf{x})-\tilde{v}(\mathbf{x})\right|\leq\sum_{i\in I}\min\left\{\left\|c_{i}(\mathbf{a}_{i}^{T},b_{i})\right\|,\delta_{\max}\right\}$
+$\frac{1}{\rho}\left|v(\mathbf{x})-\tilde{v}(\mathbf{x})\right|\leq\sum_{i\in I}\min\left\{\left\|c_{i}(\mathbf{a}_{i}^{T},b_{i})\right\|,\delta_{\max}\right\},$
 which concludes the proof of Proposition 12.
@@
 Let $I_{jk}=I_{k}\cap I_{j+}$ if $k$ is a positive cluster, else $I_{jk}=I_{k}\cap I_{j-}$
-Consider any vertex $\mathbf{u}$ of $P_{j}$. This vertex can be written as the sum of generators $c_{ji}(\mathbf{a}_{i}^{T},b_{i})$, for some subset $I_{j+}^{\prime}\subset I_{j+}$. Thus,
-$u=\sum_{i\in I_{j+}^{\prime}}c_{ji}(\mathbf{a}_{i}^{T},b_{i})$
-For every generator $c_{ji}(\mathbf{a}_{i}^{T},b_{i})$ that belongs to cluster k, define
-$x_{ji}=arg\min_{x}\left\{c_{ji}(\mathbf{a}_{i}^{T},b_{i})-x\tilde{c}_{jk}(\tilde{\mathbf{a}}_{k}^{T},\tilde{b}_{k})\right\}$
+Consider any vertex $\mathbf{u}$ of $P_{j}$. This vertex can be written as the sum of generators $c_{ji}(\mathbf{a}_{i}^{T},b_{i})$, for some subset $I_{j+}^{\prime}\subseteq I_{j+}$. Thus,
+$\mathbf{u}=\sum_{i\in I_{j+}^{\prime}}c_{ji}(\mathbf{a}_{i}^{T},b_{i}).$
+For every generator $c_{ji}(\mathbf{a}_{i}^{T},b_{i}),i\in I_{j+}$ that belongs to positive cluster $k\in C_{j+}$, define
+$x_{ji}=\arg\min_{x}\left\|c_{ji}(\mathbf{a}_{i}^{T},b_{i})-x\tilde{c}_{jk}(\tilde{\mathbf{a}}_{k}^{T},\tilde{b}_{k})\right\|,$
 i.e., project the generator onto $\tilde{c}_{jk}(\tilde{\mathbf{a}}_{k}^{T},\tilde{b}_{k})$ of its cluster $k$.
 Assume a variant of the algorithm, where the optimization criterion is the following:
-$\sum_{j=1}^{m}\left[|\tilde{C}_{jk}|(\tilde{\mathbf{a}}_{k}^{T},\tilde{b}_{k})-\sum_{i\in I_{jk}}|C_{ji}|(\mathbf{a}_{i}^{T},b_{i})\right]^{2}.$
+$\sum_{j=1}^{m}\left\||\tilde{C}_{jk}|(\tilde{\mathbf{a}}_{k}^{T},\tilde{b}_{k})-\sum_{i\in I_{jk}}|C_{ji}|(\mathbf{a}_{i}^{T},b_{i})\right\|^{2}.$
 The sign of $\tilde{C}_{jk}$ never changes and gets fixed based on the initial solution.
@@
 It is easy to verify that the representative $|\tilde{c}_{jk}|(\tilde{\mathbf{a}}_{k},\tilde{b}_{k})$ that the iterative algorithm produces lies inside the cone of the set of vectors $\{\sum_{i\in I_{jk}}|c_{ji}|(\mathbf{a}_{i}^{T},b_{i})|j\in[m]\}$, which is a subset of the cone of the set of vectors $\{(\mathbf{a}_{i}^{T},b_{i})|i\in I_{k}\}$.
-Thus, the representative forms an acute angle with every vector of the set $\{(\mathbf{a}_{i}^{T},b_{i})|i\in I_{k}\}$, and thus $x_{ji}\geq 0$
-Define $I_{jk}^{\prime}=I_{k}\cap I_{j+}^{\prime}$. For every cluster k, define
-$\tilde{x}_{jk}=\sum_{i\in I_{jk}^{\prime}}x_{ji}$
+Thus, the representative forms an acute angle with every vector of the set $\{(\mathbf{a}_{i}^{T},b_{i})|i\in I_{k}\}$, and thus $x_{ji}\geq 0$.
+For every cluster $k\in C_{j+}$ define $I_{jk}^{\prime}=I_{k}\cap I_{j+}^{\prime}$ and
+$\tilde{x}_{jk}=\sum_{i\in I_{jk}^{\prime}}x_{ji}.$
 Since $x_{ji}\geq 0$, and by the definition of the update step for the output weights, we have
 $\sum_{i\in I_{jk}}x_{ji}=1\Rightarrow\tilde{x}_{jk}=\sum_{i\in I_{jk}^{\prime}}x_{ji}\leq 1$
-Thus, for every cluster $k$, the point $\tilde{x}_{jk}\tilde{c}_{jk}(\tilde{\mathbf{a}}_{k}^{T},\tilde{b}_{k})$ lies inside the segment $[\text{O},\tilde{c}_{jk}(\tilde{\mathbf{a}}_{k}^{T},\tilde{b}_{k})$], and thus belongs to $\tilde{P_{j}}$.
-For every vertex $u$, we choose to compare it with the vertex
-$\sum_{k\in C_{j+}}\tilde{x}_{jk}\tilde{c}_{jk}(\tilde{\mathbf{a}}_{k}^{T},\tilde{b}_{k})\in\tilde{P_{j}}$
+Thus, for every cluster $k\in C_{j+}$, the point $\tilde{x}_{jk}\tilde{c}_{jk}(\tilde{\mathbf{a}}_{k}^{T},\tilde{b}_{k})$ lies inside the segment $[\mathbf{0},\tilde{c}_{jk}(\tilde{\mathbf{a}}_{k}^{T},\tilde{b}_{k})$], and thus belongs to $\tilde{P_{j}}$.
+For the vertex $\mathbf{u}$, we choose to compare it with the point
+$\sum_{k\in C_{j+}}\tilde{x}_{jk}\tilde{c}_{jk}(\tilde{\mathbf{a}}_{k}^{T},\tilde{b}_{k})\in\tilde{P_{j}}.$
 We have that
-$\tilde{c}_{jk}(\tilde{\mathbf{a}}_{k}^{T},\tilde{b}_{k})=\sum_{i\in I_{jk}}c_{ji}(\mathbf{a}_{i}^{T},b_{i})+l_{jk}=|I_{jk}|\frac{\sum_{i\in I_{jk}}c_{ji}(\mathbf{a}_{t}^{T},b_{i})}{|I_{jk}|}+l_{jk}=|I_{jk}|(c_{ji}(\mathbf{a}_{i}^{T},b_{i})+\epsilon_{ji})+l_{jk}$
-where $\sum_{j=1}^{m}l_{jk}^{2}=l_{k}^{2}$ the optimization criterion loss, and $\epsilon_{ji}$ is the error between $c_{ji}(\mathbf{a}_{i}^{T},b_{i})$ and the mean $\frac{\sum_{i\in I_{jk}}c_{ji}(\mathbf{a}_{t}^{T},b_{i})}{|I_{jk}|}$.
+$\tilde{c}_{jk}(\tilde{\mathbf{a}}_{k}^{T},\tilde{b}_{k})=\sum_{i\in I_{jk}}c_{ji}(\mathbf{a}_{i}^{T},b_{i})+l_{jk}=|I_{jk}|\frac{\sum_{i\in I_{jk}}c_{ji}(\mathbf{a}_{t}^{T},b_{i})}{|I_{jk}|}+l_{jk}=|I_{jk}|(c_{ji}(\mathbf{a}_{i}^{T},b_{i})+\epsilon_{ji})+l_{jk},$
+where $\sum_{j=1}^{m}\|l_{jk}\|^{2}=l_{k}^{2}$ the optimization criterion loss, and $\epsilon_{ji}$ is the error between $c_{ji}(\mathbf{a}_{i}^{T},b_{i})$ and the mean $\frac{\sum_{i\in I_{jk}}c_{ji}(\mathbf{a}_{t}^{T},b_{i})}{|I_{jk}|}$.
 It is easy to verify that $\epsilon_{ji}$ tends to zero as $\delta_{max}$ tends to 0.
@@
 $\leq\sum_{k\in C_{j+}}\sum_{i\in I_{jk}^{\prime}}\min\left\{\|c_{ji}(\mathbf{a}_{i}^{T},b_{i})\|,\left\|\frac{l_{jk}}{|I_{jk}|}+\epsilon_{ji}\right\|\right\}+\sum_{i\in N_{j+}}|c_{ji}|\|(\mathbf{a}_{i}^{T},b_{i})\|$
-$\leq\sum_{k\in C_{j+}}\sum_{i\in I_{jk}^{\prime}}\min\left\{\|c_{ji}(\mathbf{a}_{i}^{T},b_{i})\|,\frac{\|l_{jk}\|}{|I_{jk}|}+\|\epsilon_{ji}\|\right\}+\sum_{i\in N_{j+}}|c_{ji}|\|(\mathbf{a}_{i}^{T},b_{i})\|$
+$\leq\sum_{k\in C_{j+}}\sum_{i\in I_{jk}^{\prime}}\min\left\{\|c_{ji}(\mathbf{a}_{i}^{T},b_{i})\|,\frac{\|l_{jk}\|}{|I_{jk}|}+\|\epsilon_{ji}\|\right\}+\sum_{i\in N_{j+}}|c_{ji}|\|(\mathbf{a}_{i}^{T},b_{i})\|.$
 The maximum value of the upper bound occurs when $I_{j+}^{\prime}=I_{j+}$. Thus, we have
-$\max_{\mathbf{u}\in V_{P_{j}}}\operatorname{dist}(u,\tilde{P}_{j})\leq\sum_{k\in C_{j+}}\sum_{i\in I_{jk}}\min\left\{\|c_{ji}(\mathbf{a}_{i}^{T},b_{i})\|,\frac{\|l_{jk}\|}{|I_{jk}|}+\|\epsilon_{ji}\|\right\}+\sum_{i\in N_{j+}}|c_{ji}|\|(\mathbf{a}_{i}^{T},b_{i})\|$
+$\max_{\mathbf{u}\in V_{P_{j}}}\operatorname{dist}(\mathbf{u},\tilde{P}_{j})\leq\sum_{k\in C_{j+}}\sum_{i\in I_{jk}}\min\left\{\|c_{ji}(\mathbf{a}_{i}^{T},b_{i})\|,\frac{\|l_{jk}\|}{|I_{jk}|}+\|\epsilon_{ji}\|\right\}+\sum_{i\in N_{j+}}|c_{ji}|\|(\mathbf{a}_{i}^{T},b_{i})\|.$
 To obtain a bound for $\max_{\mathbf{v}\in V_{\tilde{P}_{j}}}\operatorname{dist}(P_{j},\mathbf{v})$, we write $\mathbf{v}=\sum_{k\in C_{j+}^{\prime}}\tilde{c}_{jk}(\tilde{\mathbf{a}}_{k}^{T},\tilde{b}_{k})\in\tilde{P}_{j}$ and choose vertex $\mathbf{u}=\sum_{i\in I_{j+}^{\prime}}c_{ji}(\mathbf{a}_{i}^{T},b_{i})$, with $I_{j+}^{\prime}=\{i\in I_{j+}|i\in I_{k},k\in C_{j+}^{\prime}\}$.
@@
 At last, we have
-$H(P_{j},\tilde{P}_{j})\leq\sum_{k\in C_{j+}}\sum_{i\in I_{jk}}\min\left\{\|c_{ji}(\mathbf{a}_{i}^{T},b_{i})\|,\frac{\|l_{jk}\|}{|I_{jk}|}+\|\epsilon_{ji}\|\right\}+\sum_{i\in N_{j+}}|c_{ji}|\|(\mathbf{a}_{i}^{T},b_{i})\|$
+$H(P_{j},\tilde{P}_{j})\leq\sum_{k\in C_{j+}}\sum_{i\in I_{jk}}\min\left\{\|c_{ji}(\mathbf{a}_{i}^{T},b_{i})\|,\frac{\|l_{jk}\|}{|I_{jk}|}+\|\epsilon_{ji}\|\right\}+\sum_{i\in N_{j+}}|c_{ji}|\|(\mathbf{a}_{i}^{T},b_{i})\|.$
 Similarly, for $H(Q_{j},\tilde{Q}_{j})$ we have
-$H(Q_{j},\tilde{Q}_{j})\\ \leq\sum_{k\in C_{j-}}\sum_{i\in I_{jk}}\min\left\{\|c_{ji}(\mathbf{a}_{i}^{T},b_{i})\|,\frac{\|l_{jk}\|}{|I_{jk}|}+\|\epsilon_{ji}\|\right\}\\ +\sum_{i\in N_{j-}}|c_{ji}|\|(\mathbf{a}_{i}^{T},b_{i})\|$
+$H(Q_{j},\tilde{Q}_{j})\\ \leq\sum_{k\in C_{j-}}\sum_{i\in I_{jk}}\min\left\{\|c_{ji}(\mathbf{a}_{i}^{T},b_{i})\|,\frac{\|l_{jk}\|}{|I_{jk}|}+\|\epsilon_{ji}\|\right\}\\ +\sum_{i\in N_{j-}}|c_{ji}|\|(\mathbf{a}_{i}^{T},b_{i})\|.$
 Combining, we obtain
-$\frac{1}{\rho}\max_{\mathbf{x}\in B}|v_{j}(\mathbf{x})-\tilde{v}_{j}(\mathbf{x})|\leq\sum_{k=1}^{m}\sum_{i\in I_{jk}}\min\left\{\|c_{ji}(\mathbf{a}_{i}^{T},b_{i})\|,\frac{\|l_{jk}\|}{|I_{jk}|}+\|\epsilon_{ji}\|\right\}\\ +\sum_{i\in N_{j}}|c_{ji}|\|(\mathbf{a}_{i}^{T},b_{i})\|$
-Using the fact that $I_{jk}\subset I_{k}$ and $N_{min}\leq|I_{jk}|,\forall j,k$ we have
-$\frac{1}{\rho}\max_{\mathbf{x}\in B}|v_{j}(\mathbf{x})-\tilde{v}_{j}(\mathbf{x})|\leq\sum_{k=1}^{m}\sum_{i\in I_{k}}\min\left\{\|c_{ji}(\mathbf{a}_{i}^{T},b_{i})\|,\frac{\|l_{jk}\|}{N_{min}}+\|\epsilon_{ji}\|\right\}+\sum_{i\in N_{j}}|c_{ji}|\|(\mathbf{a}_{i}^{T},b_{i})\|$
+$\frac{1}{\rho}\max_{\mathbf{x}\in B}|v_{j}(\mathbf{x})-\tilde{v}_{j}(\mathbf{x})|\leq\sum_{k=1}^{m}\sum_{i\in I_{jk}}\min\left\{\|c_{ji}(\mathbf{a}_{i}^{T},b_{i})\|,\frac{\|l_{jk}\|}{|I_{jk}|}+\|\epsilon_{ji}\|\right\}\\ +\sum_{i\in N_{j}}|c_{ji}|\|(\mathbf{a}_{i}^{T},b_{i})\|.$
+Using the fact that $I_{jk}\subseteq I_{k}$ and $N_{min}\leq|I_{jk}|,\forall j,k$ we have
+$\frac{1}{\rho}\max_{\mathbf{x}\in B}|v_{j}(\mathbf{x})-\tilde{v}_{j}(\mathbf{x})|\leq\sum_{k=1}^{m}\sum_{i\in I_{k}}\min\left\{\|c_{ji}(\mathbf{a}_{i}^{T},b_{i})\|,\frac{\|l_{jk}\|}{N_{min}}+\|\epsilon_{ji}\|\right\}+\sum_{i\in N_{j}}|c_{ji}|\|(\mathbf{a}_{i}^{T},b_{i})\|.$
 We make use of the following inequality, which is a direct consequence of Cauchy-Schwartz Inequality
-$\sum_{j=1}^{m}|u_{j}|\leq\sqrt{m}\sqrt{\sum_{j=1}^{m}|u_{j}|^{2}}=\sqrt{m}\|(u_{1},\dots,u_{m})\|$
+$\sum_{j=1}^{m}|u_{j}|\leq\sqrt{m}\sqrt{\sum_{j=1}^{m}|u_{j}|^{2}}=\sqrt{m}\|(u_{1},\dots,u_{m})\|.$
 We have
-$\sum_{j=1}^{m}|c_{ji}|\leq\sqrt{m}\|\mathbf{C}_{:,i}\|$
-$\sum_{j=1}^{m}\|l_{jk}\|\leq\sqrt{m}\sqrt{\sum_{j=1}^{m}\|l_{jk}\|^{2}}=\sqrt{m}\cdot l_{k}$
-$\sum_{j=1}^{m}\|\epsilon_{ji}\|\leq\sqrt{m}\sqrt{\sum_{j=1}^{m}\|\epsilon_{ji}\|^{2}}=\sqrt{m}\|\epsilon_{:,i}\|_{F}$
+$\sum_{j=1}^{m}|c_{ji}|\leq\sqrt{m}\|\mathbf{C}_{:,i}\|,$
+$\sum_{j=1}^{m}\|l_{jk}\|\leq\sqrt{m}\sqrt{\sum_{j=1}^{m}\|l_{jk}\|^{2}}=\sqrt{m}\cdot l_{k},$
+$\sum_{j=1}^{m}\|\epsilon_{ji}\|\leq\sqrt{m}\sqrt{\sum_{j=1}^{m}\|\epsilon_{ji}\|^{2}}=\sqrt{m}\|\epsilon_{:,i}\|_{F}.$
 Using the above inequalities, the fact that $\sum\min\leq\min\sum$, and the fact that $\max\sum\leq\sum\max$ we get
@@
 $\leq\sqrt{m}\sum_{i=1}^{n}\min\left\{\|\mathbf{C}_{:,i}\|\|(\mathbf{a}_{i}^{T},b_{i})\|,\frac{l_{k(i)}}{N_{min}}+\|\epsilon_{:,i}\|_{F}\right\}$
-$+\sum_{j=1}^{m}\sum_{i\in N_{j}}|c_{ji}|\|(\mathbf{a}_{i}^{T},b_{i})\|$
-as desired
+$+\sum_{j=1}^{m}\sum_{i\in N_{j}}|c_{ji}|\|(\mathbf{a}_{i}^{T},b_{i})\|,$
+as desired.

### v3

@@
 Abstract
-We present TropNNC, a framework for compressing neural networks with linear and convolutional layers and ReLU activations.
-TropNNC is a structured compression framework based on a geometrical approach to machine/deep learning, using tropical geometry and extending the work of Misiakos et al.
-(2022).
-We use the Hausdorff distance of zonotopes in its standard continuous form to achieve a tighter approximation bound for tropical polynomials compared to previous work.
-This enhancement leads to the development of an effective compression algorithm that achieves superior functional approximations of neural networks.
-Our method is significantly easier to implement compared to other frameworks, and does not depend on the availability of training data samples.
-We validate our framework through extensive empirical evaluations on the MNIST, CIFAR, and ImageNet datasets.
-Our results demonstrate that TropNNC achieves performance on par with state-of-the-art methods like ThiNet (even surpassing it in compressing linear layers) and CUP.
-To the best of our knowledge, it is the first method that achieves this using tropical geometry.
-Keywords: tropical geometry, zonotopes, Hausdorff distance, neural network compression
+We present TropNNC, a framework for compressing neural networks with linear and convolutional layers and ReLU activations using tropical geometry.
+By representing a network’s output as a tropical rational function, TropNNC enables structured compression via reduction of the corresponding tropical polynomials.
+Our method refines the geometric approximation of previous work by adaptively selecting the weights of retained neurons.
+Key contributions include the first application of tropical geometry to convolutional layers and the tightest known theoretical compression bound.
+TropNNC requires only access to network weights – no training data – and achieves competitive performance on MNIST, CIFAR, and ImageNet, matching strong baselines such as ThiNet and CUP.
+11footnotetext: Corresponding author.
 1 Introduction
-In recent years, deep neural networks (DNNs) have significantly advanced the field of computer vision, excelling in tasks such as image classification (Krizhevsky et al., 2012; Simonyan and Zisserman, 2015; He et al., 2016; Rawat and Wang, 2017), object detection (Zou et al., 2023), semantic segmentation (Noh et al., 2015; Guo et al., 2018), and image captioning (Jia et al., 2015; Hossain et al., 2019).
-These networks have outperformed traditional methods reliant on manually crafted visual features.
-However, despite their success, deploying DNNs on resource-constrained devices, such as mobile phones or embedded systems, remains challenging due to their substantial computational and storage demands.
-For example, the VGG-16 model (Simonyan and Zisserman, 2015), with its approximately $138$ million parameters, requires over $500$MB of storage and performs about $31$ billion FLOPs to classify a single image.
-Such resource-intensive models exceed the capabilities of small devices and pose significant challenges, including energy limitations recently highlighted by the deployment of large language models (Samsi et al., 2023).
-This has led to a growing focus on network compression techniques (Blalock et al., 2020) as a critical area of research.
-Initial attempts to reduce network complexity date back to the 1990s, when LeCun et al.
-(1989) demonstrated that removing negligible weights from a network had minimal impact on performance.
-This approach, known as neural network pruning, has since evolved.
-Han et al.
-(2015) proposed a simple yet effective pruning strategy: iteratively removing connections with weights below a certain threshold and fine-tuning the network to recover accuracy, resulting in highly sparse models.
-These methods belong to the category of unstructured pruning, which alters the network’s structure by eliminating individual weights.
-While effective in reducing network size, unstructured pruning presents challenges in practical applications.
-Such pruned networks often require specialized hardware and software for efficient inference due to issues like poor cache locality and jumping memory accesses, which can negate the benefits of sparsity and sometimes even degrade performance (Wen et al., 2016).
-To overcome the limitations of unstructured pruning, structured pruning methods, such as channel-level pruning, have been proposed (He and Xiao, 2024).
+Deploying deep neural networks on resource-constrained devices remains challenging due to their substantial computational and storage demands.
+Initial attempts to reduce network complexity include unstructured pruning.
+which alters the network’s structure by eliminating individual weights.
+While effective in reducing network size, unstructured pruning presents challenges in practical applications (Wen et al., 2016).
+To overcome its limitations, structured pruning methods, such as channel-level pruning, have been proposed (He and Xiao, 2024).
 Notably, the ThiNet framework by Luo et al.
 (2017) prunes entire filters or channels, maintaining the network’s original structure.
-In another work, Wen et al.
-(2016) proposed the structured sparsity learning approach, which also includes filter- or channel-wise pruning.
-This structured approach ensures compatibility with existing deep learning libraries and offers several advantages: it significantly reduces memory footprint, facilitates further compression and acceleration through methods like parameter quantization (Gong et al., 2014; Chen et al., 2015; Wu et al., 2016), and accelerates various vision tasks.
-Parallel to these advancements, tropical geometry (Maclagan and Sturmfels, 2021) has emerged as a promising mathematical framework with applications in neural networks and machine learning (Maragos et al., 2021).
-Tropical geometry, rooted in algebraic geometry, operates over the tropical semiring (Cuninghame-Green, 1994), where conventional addition and multiplication are replaced by min and plus operations, respectively.
-This transformation turns polynomials into piecewise linear functions, making tropical geometry highly relevant to the study of neural networks with piecewise linear activations like ReLU.
-Tropical mathematics encompasses a wide range of applications, which include operations research and scheduling (Cuninghame-Green, 1979), control and optimization (Baccelli et al., 2001; Cohen et al., 2004; Butkovič, 2010; Akian et al., 2012), mathematical physics (Litvinov et al., 2001), speech recognition (Mohri et al., 2002), polyhedral geometry (Gaubert and Katz, 2011), tropical geometry (Maclagan and Sturmfels, 2021), dynamical systems on weighted lattices (Maragos, 2017), sparsity theory on weighted lattices (Tsilivis et al., 2022), matrix factorization (Kordonis et al., 2024), finite state transducers (Theodosis and Maragos, 2018), convex regression (Maragos and Theodosis, 2020), and machine learning (Maragos et al., 2021; Gärtner and Jaggi, 2008).
-Recently, tropical geometry has been applied to the theoretical study of neural networks.
+This structured approach ensures compatibility with existing deep learning libraries and offers several advantages: it significantly reduces memory footprint, and facilitates further compression and acceleration through methods like parameter quantization.
+Parallel to these advancements, tropical geometry (Maclagan and Sturmfels, 2021) has emerged as a promising mathematical framework with applications in machine learning (Maragos et al., 2021; Gärtner and Jaggi, 2008) and beyond.
+Recently, it has been applied to the theoretical study of neural networks.
 For example, Zhang et al.
-(2018) demonstrated the equivalence of ReLU-activated neural networks with tropical rational mappings, and used zonotopes to compute a bound on the number of linear regions of networks, equal to the one in (Montufar et al., 2014).
-Other works, like those of Charisopoulos and Maragos (2018); Alfarra et al.
-(2023); Smyrnis and Maragos (2020), have used tropical geometry to compute bounds on the number of linear regions in neural networks and to develop methods for pruning and compressing these networks.
-In this paper, we explore the application of tropical geometry to neural network compression.
-We build on the work of Misiakos et al.
-(2022), who utilized tropical geometry for structured neural network compression by geometrically approximating the zonotopes corresponding to the network layers.
-Our contributions include:
-  • Obtaining a tighter bound for the functional approximation of tropical polynomials than that presented in (Misiakos et al., 2022) using the standard, continuous Hausdorff distance.
-  • Designing an algorithm that leverages tropical geometry and the Hausdorff distance for the structured compression of neural networks.
-Our algorithm refines the choice of weights in the compressed network from the clustering done in (Misiakos et al., 2022) and we also provide an iterative technique to further optimize them.
-Our method is simple to implement and does not depend on the availability of training data samples like other frameworks such as ThiNet (Luo et al., 2017).
-It is also the first method based on the theory of tropical geometry that compresses convolutional layers.
-  • Evaluating our algorithm through experiments on MNIST, Fashion-MNIST, CIFAR-10, CIFAR-100, and ImageNet datasets.
-Our method outperforms that of Misiakos et al.
-(2022) and achieves competitive or superior performance compared to ThiNet, especially in compressing linear layers, and superior performance compared to CUP (Duggal et al., 2021), particularly in the VGG architecture.
-This work demonstrates the potential of tropical geometry in enhancing neural network compression techniques, providing a theoretical foundation and a practical algorithm that improves efficiency and maintains accuracy.
-2 Related Work
-Seminal works have showcased that neural networks tend to be hyperparameterized.
-Denil et al.
-(2013) demonstrated that neural networks can be reconstructed using a small subset of parameters, which sparked significant interest in network compression techniques.
-Blalock et al.
-(2020) reviewed various methods for reducing model size without compromising performance, highlighting the ongoing advancements in this area.
-Early methods for network compression focused on unstructured pruning.
-LeCun et al.
-(1989) showed that removing unimportant weights leads to minimal performance loss.
-This approach was further refined by Han et al.
-(2015), who proposed pruning weights below a certain threshold followed by fine-tuning the network.
-Despite their effectiveness, these methods require specialized hardware for efficient inference due to their unstructured nature.
-Structured pruning techniques maintain the network’s architecture, facilitating better compatibility and efficiency.
-Wen et al.
-(2016) introduced structured sparsity learning to compress networks by finding compact representations of filters and channels, preserving the architecture.
-Luo et al.
-(2017) proposed ThiNet, a method that greedily prunes filters based on their impact on the next layer, thereby maintaining network structure and efficiency.
-Yu et al.
-(2018) presented the Neuron Importance Score Propagation (NISP) algorithm, which extended the ideas of ThiNet by pruning neurons and channels based on their impact on the final layer output.
-Lin et al.
-(2020) introduced HRank, which uses feature map information to rank and prune filters.
-Duggal et al.
-(2021) proposed Cluster Pruning (CUP), which prunes similar filters by clustering them based on features derived from both incoming and outgoing weight connections.
-CUP addresses the limitations of prior work by efficiently determining the ideal number of filters to prune in each layer and integrating pruning into the initial training phase, leading to significant savings in training time and maintaining performance.
-Wang et al.
-(2021) introduced a novel approach to pruning convolutional neural networks that focuses on reducing structural redundancy rather than merely pruning individual neurons or channels.
-Their method identifies and eliminates redundant structures within the network architecture, thereby achieving more efficient pruning without compromising the network’s performance.
-Smyrnis and Maragos (2020) proposed a tropical division algorithm for structured network compression, further enhancing the capabilities of structured pruning techniques.
-Misiakos et al.
-(2022) developed Neural Path K-means, which uses clustering techniques based on tropical geometry and functional approximation for structured compression.
-Beyond pruning, other state-of-the-art strategies for network compression include parameter quantization and low-rank approximation.
-Parameter quantization, as explored by Zhou et al.
-(2017), Zhou et al.
-(2018) and Han et al.
-(2015), reduces the number of bits per parameter, significantly lowering memory and computational requirements.
-Behdin et al.
-(2023) presented an optimization-based framework for efficient post-training quantization of large language models.
-Polino et al.
-(2018) presented quantized distillation, a method that combines quantization with teacher-student distillation.
-Low-rank approximation (LRA) techniques decompose weight matrices into smaller matrices, reducing parameters and computations.
-Studies by Denton et al.
-(2014), Jaderberg et al.
-(2014), Sindhwani et al.
-(2015), Ioannou et al.
-(2016), Tai et al.
-(2016), and Chen et al.
-(2018) have shown that LRA can achieve significant speedups and compression with minimal accuracy loss, although it typically requires iterative optimization.
-These methods can be combined with structured pruning approaches like ThiNet to achieve further model compression and efficiency.
-Our framework is a structured pruning method that extends the approach of Misiakos et al.
-(2022).
-It prunes networks layer-by-layer, aiming to minimize the error introduced to the input of the next layer.
-In that sense, it is similar to the ThiNet framework.
-However, our algorithm is simpler to implement and does not require a training dataset.
-Our framework leverages clustering and can be naturally extended to non-uniform pruning, drawing parallels to the CUP framework.
-In fact, it improves upon steps 1, 2, and 3 of CUP.
-For step 1, we use similar clustering vectors (referred to as filter features by CUP authors Duggal et al.
-(2021)) for linear layers, and introduce a new method for constructing clustering vectors in convolutional layers.
-For step 2, we introduce two approaches for choosing the distance threshold of the hierarchical clustering.
-Our main contribution lies in step 3, where we employ tropical geometry to select a better representative for each cluster, rather than simply choosing the neuron or channel with the most energy.
-While it is tempting to compare our framework with those of NISP (Yu et al., 2018) and HRank (Lin et al., 2020) given that they are both structured pruning methods, it is important to recognize the distinctions between them and ours.
-NISP and HRank are advanced frameworks specifically designed to address the limitations of layer-by-layer pruning approaches, such as those employed by ThiNet and our own method.
-In contrast, our framework leverages a rich theoretical foundation and does not rely on training data samples, offering a unique advantage in certain contexts.
-We believe our results are promising and present significant advantages.
-Our work highlights the potential of tropical geometrical techniques and paves the way for further research that could enhance their competitiveness with current state-of-the-art methods.
+(2018) demonstrated the equivalence of ReLU-activated neural networks with tropical rational mappings.
+Together with other works, like those of Charisopoulos and Maragos (2018); Alfarra et al.
+(2023), they used tropical geometry to compute bounds on the number of linear regions of neural networks equal to the one in (Montufar et al., 2014).
+Smyrnis and Maragos (2020) use tropical geometry to develop pruning methods.
+Contributions.
+In this paper, we explore the application of tropical geometry in the compression of ReLU neural networks. Our contributions include:
+  • Proposing TropNNC, an algorithm that leverages tropical geometry and the Hausdorff distance for the structured compression of neural networks.
+TropNNC compresses convolutional networks layer-wise by approximating the zonotopes corresponding to each layer using an iterative process.
+Our algorithm is data agnostic: it does not require a training dataset.
+  • Refining the bound for the functional approximation of tropical polynomials presented in (Misiakos et al., 2022), and using it to provide stronger theoretical compression guarantees.
+  • Evaluating our algorithm empirically on MNIST, CIFAR and ImageNet datasets.
+Our method outperforms prior tropical geometrical pruning methods, and achieves competitive performance compared to the data-driven ThiNet, and superior performance compared to CUP (Duggal et al., 2021), particularly in the VGG architecture.
+This work demonstrates the potential of tropical geometry in enhancing neural network compression techniques.
+To the best of our knowledge, it is the first tropical geometric pruning method that compresses convolutional layers and is competitive with strong baselines.
+Even if currently not state-of-the-art, we believe that our contribution is a necessary step towards this direction.
+Related work is provided after the definition of our algorithm to enable detailed comparisons.
+Proofs of theoretical results are provided in the appendix.
 2 Tropical Geometry of Neural Networks
-In this section, we introduce the fundamental background of tropical geometry and its applications to neural network compression.
-3.1 Background on Tropical Geometry
-Tropical algebra is a field of mathematics that studies the tropical semiring.
+Tropical algebra studies matrix-vector operations based on the arithmetic of the tropical semiring Cuninghame-Green (1979); Butkovič (2010).
 Tropical geometry is the counterpart of algebraic geometry in the tropical setting.
-The tropical semiring can refer to either the min-plus semiring or the max-plus semiring (Butkovič, 2010).
-In this work, we adhere to the convention of using the max-plus semiring $(\mathbb{R}_{\mathrm{max}},\vee,+)$, defined as the set $\mathbb{R}_{\mathrm{max}}=\mathbb{R}\cup\{-\infty\}$ equipped with two binary operations: $\vee$ (ordinary max) and $+$ (ordinary sum).
-3.1.1 Tropical Polynomials and Rational Functions
-Within the max-plus semiring, we can define polynomials. A tropical polynomial $f$ in $d$ variables $\mathbf{x}=(x_{1},\ldots,x_{d})$ is defined as the function:
-$f(\mathbf{x})=\bigvee_{i=1}^{n}\left\{\mathbf{a}_{i}^{T}\mathbf{x}+b_{i}\right\}=\max_{i\in[n]}\left\{\mathbf{a}_{i}^{T}\mathbf{x}+b_{i}\right\},$ (1)
-where $[n]:=\{1,...,n\}$.
-Here, $n$ represents the rank of the tropical polynomial.
-Each monomial term $\{\mathbf{a}_{i}^{T}\mathbf{x}+b_{i}\}$ of the polynomial has an exponent or slope $\mathbf{a}_{i}\in\mathbb{R}^{d}$ and a coefficient or bias $b_{i}\in\mathbb{R}$.
-Each monomial term corresponds to a plane in $\mathbb{R}^{d+1}$.
-Consequently, tropical polynomials are piecewise linear convex functions.
-Specifically, every tropical polynomial is a continuous piecewise linear convex function, and every continuous piecewise linear convex function can be expressed (though not uniquely) as a tropical polynomial (Maclagan and Sturmfels, 2021).
-The set of tropical polynomials in $\mathbf{x}$ defines the semiring $\mathbb{R}_{\mathrm{max}}[\mathbf{x}]$.
-Figures 1(a), 1(b) illustrate examples of tropical polynomials in one and multiple variables, respectively.
-Tropical rational functions extend the concept of rational polynomial functions to tropical algebra.
-A tropical rational function is defined as the tropical multiplication of a tropical polynomial $p$ by the tropical multiplicative inverse $q^{-1}$ of another tropical polynomial $q$.
-In conventional arithmetic, this operation corresponds to the difference of two tropical polynomials $p$ and $q$:
-$r(\mathbf{x})=p(\mathbf{x})-q(\mathbf{x})$
-Tropical rational functions correspond to general piecewise linear functions.
-Specifically, every tropical rational function is a continuous piecewise linear function, and every continuous piecewise linear function can be expressed (though not uniquely) as a tropical rational function.
-Figures 1(c), 1(d) provide examples of tropical rational functions.
-(a) $f(x)=\max\{x,1\}$
-(b) $f(x,y)=\max\{x,y,1\}$
-(c) $f(x)=\max\{2x-2,-x\}-\max\{-2x-2,x\}$
-(d) $f(x,y)=\max\{x,-x\}-\max\{y,-y\}$
-Figure 1: (a) depicts a single-variate tropical polynomial, (b) depicts a multi-variate tropical polynomial, (c) depicts a single-variate tropical rational function, (d) depicts a multi-variate tropical rational function
-3.1.2 Newton Polytopes
-As with algebraic geometry (Hartshorne, 2013), in tropical geometry we can define the Newton polytope (Monical et al., 2019) of a tropical polynomial.
-Newton polytopes can be used to analyse the behavior of a polynomial.
-They connect tropical geometry with polytope theory, an extensively studied field of mathematics (Ziegler, 2012).
-For a tropical polynomial $f$ as defined in (1), we define its Newton polytope as the convex hull of the slopes $\mathbf{a}_{i}$ of $f$
-$\mathrm{Newt}(f):=\mathrm{conv}\left\{\mathbf{a}_{i},i\in[n]\right\}.$
-Additionally, we define the extended Newton polytope of a tropical polynomial $f$ as the convex hull of the slopes $\mathbf{a}_{i}$ of $f$ extended in their last dimension by the coefficient $b_{i}$
-$\mathrm{ENewt}(f):=\mathrm{conv}\left\{(\mathbf{a}_{i}^{T},b_{i}),i\in[n]\right\}.$
-The following proposition allows us to calculate the (extended) Newton polytope of expressions of tropical polynomials.
-Proposition 1 (Zhang et al., 2018)
-Let $f,g\in\mathbb{R}_{\mathrm{max}}[\mathbf{x}]$ be two tropical polynomials in $\mathbf{x}$. For the extended Newton polytope, the following holds:
-$\mathrm{ENewt}\left(f\vee g\right)$ $=\mathrm{conv}\left\{\mathrm{ENewt}(f)\cup\mathrm{ENewt}(g)\right\}$
-$\mathrm{ENewt}(f+g)$ $=\mathrm{ENewt}(f)\oplus\mathrm{ENewt}(g).$
-Here, $\vee$ denotes the tropical addition (taking the maximum), and $+$ denotes the tropical multiplication (ordinary addition).
-The Minkowski sum $\oplus$ of two polytopes (or more generally subsets of $\mathbb{R}^{d}$) $P$ and $Q$ is defined as:
-$P\oplus Q:=\{p+q\mid p\in P,q\in Q\}.$
-Using the principle of induction, Proposition 1 can be generalized to any finite tropical expression of tropical polynomials.
-The extended Newton polytope provides a geometrical interpretation for studying tropical polynomials.
-The following propositions establish the relationship between tropical polynomials and their extended Newton polytopes.
-The upper envelope or upper hull $UF(P)$ of an extended Newton polytope is defined as the set of all points $(\mathbf{a}^{T},b)$ of the polytope $P$ that are not ”shadowed” by any other part of the polytope when viewed from above (last dimension).
-This means that there is no $b^{\prime}>b$ such that $(\mathbf{a}^{T},b^{\prime})$ belongs to $P$.
-We have the following useful lemma (proof in Appendix A.1).
-Lemma 2
-Let $p\in\mathbb{R}_{\mathrm{max}}[\mathbf{x}]$ a tropical polynomial in $d$ variables with extended Newton polytope $P=\mathrm{ENewt}(p)$.
-If $(\mathbf{a}^{T},b)$ lies below the upper envelope of $P$, then $\forall\mathbf{x}\in\mathbb{R}^{d},\mathbf{a}^{T}\mathbf{x}+b\leq p(\mathbf{x})$.
-The inequality is strict if $(\mathbf{a}^{T},b)$ lies strictly below the upper envelope.
-An immediate consequence of the above lemma is the following theorem:
-Theorem 3
-Let $f\in\mathbb{R}_{\mathrm{max}}[\mathbf{x}]$ be a tropical polynomial. The values of $f$ are fully determined by the upper envelope $UF(\mathrm{ENewt}(f))$ of its extended Newton polytope.
-Theorem 4 (Charisopoulos and Maragos, 2018)
-The linear regions of a tropical polynomial $f\in\mathbb{R}_{\mathrm{max}}[\mathbf{x}]$ are in one-to-one correspondence with the vertices of $UF(\mathrm{ENewt}(f))$.
-In fact, each vertex $(\mathbf{a}_{i}^{T},b_{i})\in UF(\mathrm{ENewt}(f))$ gives exactly one linear region $D$ where $f(\mathbf{x})=\mathbf{a}_{i}^{T}\mathbf{x}+b_{i}$.
-The above theorem highlights the direct relationship between the geometric structure of the extended Newton polytope and the piecewise linear nature of the tropical polynomial.
-An immideate consequence of Theorems 3 and 4 is the following theorem:
-Theorem 5
-For any two tropical polynomials $f,g\in\mathbb{R}_{\mathrm{max}}[\mathbf{x}]$, the following holds:
-$f=g\Leftrightarrow UF(\mathrm{ENewt}(f))=UF(\mathrm{ENewt}(g))$
-This implies that two tropical polynomials are functionally identical if and only if their extended Newton polytopes have the same upper envelope.
-The above theorems indicate that a tropical polynomial is fully functionally determined by the upper envelope of its extended Newton polytope, as shown by the following example.
-Example 1
-Consider the polynomials corresponding to Figure 2.
-$f(x,y)=\max\{0,-y+1,y+1\}$
-$g(x,y)=\max\{x+1,-x+1\}.$
-We have that
-$(f\vee g)(x,y)=\max\{0,x+1,y+1,-y+1,-x+1\}$
-$(f+g)(x,y)=\max\{x+1,-x+1,x-y+2,-x-y+2,x+y+2,-x+y+2\}.$
-The extended Newton polytopes of $f,g,f\vee g,f+g$ are shown in Figure 2. The polynomial $f\vee g$ can be reduced as follows:
-$(f\vee g)(x,y)=\max\{x+1,y+1,-y+1,-x+1\},$
-which corresponds to the vertices of the upper envelope of $\mathrm{ENewt}(f\vee g)$.
-(a) $\mathrm{ENewt}(f)$
-(b) $\mathrm{ENewt}(g)$
-(c) $\mathrm{ENewt}(f\vee g)$
-(d) $\mathrm{ENewt}(f+g)$
-Figure 2: Operations on tropical polynomials.
-$\mathrm{ENewt}(f\vee g)$ corresponds to the convex hull of the union of the vertices of the polytopes $\mathrm{ENewt}(f),\mathrm{ENewt}(g)$.
-$\mathrm{ENewt}(f+g)$ corresponds to the Minkowski sum of $\mathrm{ENewt}(f),\mathrm{ENewt}(g)$.
-For the polytope $\mathrm{ENewt}(f\vee g)$ we illustrate with blue the upper envelope, which consists of a single face.
-The vertices of the upper envelope are the only non reduntant terms of the polynomial $f\vee g$.
-3.1.3 Zonotopes
-In polytope theory, zonotopes are a special class of convex polytopes that can be defined as the Minkowski sum of a finite set of line segments (or edges).
-Formally, given a set of line segments $g_{1},\ldots,g_{n}$, the zonotope is defined as
-$Z:=\bigoplus_{i\in[n]}g_{i}$
-The line segments $g_{i}$ are referred to as the zonotope’s generators. Each generator contributes to the shape and size of the zonotope, which is formed by the space swept out by these line segments.
-Alternatively, a zonotope can be expressed equivalently by a set of vectors $\mathbf{v}_{1},\ldots,\mathbf{v}_{n}\in\mathbb{R}^{d}$ and a starting point $\mathbf{s}\in\mathbb{R}^{d}$.
-By taking the generators to be the segments $[\mathbf{0},\mathbf{v}_{1}],\ldots,[\mathbf{0},\mathbf{v}_{n}]$ and translating the first segment by $\mathbf{s}$, we obtain the equivalent form:
-$Z=\left\{\mathbf{s}+\sum_{i=1}^{n}\lambda_{i}\mathbf{v}_{i}\mid 0\leq\lambda_{i}\leq 1\right\}$
-In this context, the vectors $\mathbf{v}_{i}$ are sometimes referred to as the zonotope’s generators, meaning the segments $[\mathbf{0},\mathbf{v}_{i}]$.
-When the starting point $\mathbf{s}$ is not mentioned, it is assumed to be the origin $\mathbf{0}$.
-For a zonotope with a starting point $\mathbf{s}\in\mathbb{R}^{d}$ and generators $\mathbf{v}_{1},\ldots,\mathbf{v}_{n}\in\mathbb{R}^{d}$, a vertex $\mathbf{u}\in V_{Z}$ corresponds to points where $\lambda_{i}=0\ \text{or}\ 1$.
-The vertex $\mathbf{u}$ can be expressed as:
-$\mathbf{u}=\mathbf{s}+\sum_{i\in I}\mathbf{v}_{i},$
-where $I\subseteq[n]$.
-It can be shown that zonotopes have up to exponential many verticesIn fact, for generators in general positions they have exactly $2\sum_{j=0}^{d-1}nCr(n-1,j)$ vertices (Gritzmann and Sturmfels, 1993).
-Zonotopes exhibit several interesting properties.
-Notably, they are centrally symmetric.
-This symmetry, along with their convex structure, make zonotopes particularly useful in various fields, including optimization (Bern and Eppstein, 2001) and computational geometry (Ziegler, 2012).
+The tropical semiring can refer to either the min-plus semiring or the max-plus semiring.
+In this work, we adhere to the convention of using the max-plus semiring $(\mathbb{R}_{\mathrm{max}},\vee,+)$, defined as the set $\mathbb{R}_{\mathrm{max}}=\mathbb{R}\cup\{-\infty\}$ equipped with two binary operations: $\vee$ (ordinary max) and ++ (ordinary sum).
+Within the max-plus semiring, we can define tropical polynomials, which correspond to convex piecewise linear functions, and tropical rational functions, which correspond to general continuous piecewise linear functions.
+We can also define Newton polytopes of tropical polynomials, which connect tropical algebra with polytope theory.
+For a detailed introduction to these concepts refer to the appendix.
 2.1 Neural Networks with Piecewise Linear Activations
-Tropical geometry provides a mathematical framework for analyzing neural networks with piecewise linear activation functions.
-Leveraging the properties of tropical polynomials and Newton polytopes, it is used to model the behavior of various activation functions commonly used in neural networks.
-In this work, we focus on ReLU-activated networks.
-In this section, we outline the basic tropical geometrical properties of ReLU-activated networks and establish our notations.
+Tropical geometry provides a mathematical framework for analyzing neural networks with piecewise linear activation functions. In this work, we focus on ReLU-activated networks.
 ReLU Activations.
@@
 Figure 1: Neural network with one hidden ReLU layer.
-The first linear layer has weights $\{\mathbf{a}_{i}^{T}\}$ with bias $\{b_{i}\}$ corresponding to $i$-th node $\forall i\in[n]$ and the second has weights $\{c_{ji}\}$, $\forall j\in[m],i\in[n]$
+The first linear layer has weights $\{\mathbf{a}_{i}^{T}\}$ with bias $\{b_{i}\}$ corresponding to node $i\in[n]$ and the second has weights $\{c_{ji}\}$ between nodes $j\in[m],i\in[n]$
 The output of the ReLU unit $i$ is given by:
 $f_{i}(\mathbf{x})=\mathrm{ReLU}(\mathbf{a}_{i}^{T}\mathbf{x}+b_{i})=\max\{\mathbf{a}_{i}^{T}\mathbf{x}+b_{i},0\}.$
-This expression represents a tropical polynomial of rank 2, with one term being the constant 00.
+This expression represents a tropical polynomial of rank 2, with one term being the constant 0.
 The extended Newton polytope $\mathrm{ENewt}(f_{i})$ of $f_{i}$ is an edge with one endpoint at the origin $\mathbf{0}$ and the other endpoint at $(\mathbf{a}_{i}^{T},b_{i})$.
@@
 We call $p_{j}$ the positive polynomial and $q_{j}$ the negative polynomial of $v_{j}$.
-This result can be extended to deeper networks, as suggested by the following proposition:
-Theorem 6 (Zhang et al., 2018)
-A ReLU-activated deep neural network $\mathbf{F}:\mathbb{R}^{d}\rightarrow\mathbb{R}^{m}$ is a tropical rational mapping (vector whose elements are tropical rational functions).
+Zonotopes.
 The extended Newton polytope of $|c_{ji}|f_{i}$ is an edge with one endpoint at the origin $\mathbf{0}$ and the other at $|c_{ji}|(\mathbf{a}_{i}^{T},b_{i})$.
@@
 3 Approximation based on Hausdorff distance
-In this section, we present our improved theorem, which uses the Hausdorff distance in its standard continuous form to bound the error between the original and approximate neural networks.
-Misiakos et al.
-(2022) derived a bound for the error of approximating the tropical polynomials that represent a neural network.
-The motivation for geometrically bounding the error of these polynomials stems from Theorem 5.
-This theorem indicates that two polynomials with the same extended Newton polytopes are functionally equivalent.
-Consequently, it is expected that two tropical polynomials with approximately equal extended Newton polytopes will attain similar values.
+In this section, we present our refined theorem, which uses the Hausdorff distance in its standard continuous form to bound the error between two tropical polynomials.
+In the following, the distance between two points $\mathbf{u}$ and $\mathbf{v}$ is denoted as $\text{dist}(\mathbf{u},\mathbf{v}):=\|\mathbf{u}-\mathbf{v}\|$, where $\|\cdot\|$ denotes the standard $L^{2}$ Euclidean norm.
+The distance between a point $\mathbf{u}$ and a set $V$ is defined as $\operatorname{dist}(\mathbf{u},V)=\operatorname{dist}(V,\mathbf{u}):=\inf_{\mathbf{v}\in V}\|\mathbf{u}-\mathbf{v}\|$.
+Definition 1 (Hausdorff distance).
+Let $S,\tilde{S}$ be two subsets of $\mathbb{R}^{d}$. The Hausdorff distance H(S,S~)H($\tilde{S}$) of the two sets is defined as
+$H(S,\tilde{S}):=\max\left\{\sup_{\mathbf{u}\in S}\mathrm{dist}(\mathbf{u},\tilde{S}),\sup_{\mathbf{v}\in\tilde{S}}\mathrm{dist}(S,\mathbf{v})\right\}$
+In the case of polytopes $P,\tilde{P}$, due to their convexity and compactness, the suprema in the above expression are attained, and in fact by points in the vertex sets $V_{P},V_{\tilde{P}}$ of the polytopes.
 […]
-Definition 7 (Hausdorff distance)
-Let $S,\tilde{S}$ be two subsets of $\mathbb{R}^{d}$. The Hausdorff distance $H(S,\tilde{S})$ of the two sets is defined as
-$H(S,\tilde{S}):=\max\left\{\sup_{\mathbf{u}\in S}\mathrm{dist}(\mathbf{u},\tilde{S}),\sup_{\mathbf{v}\in\tilde{S}}\mathrm{dist}(S,\mathbf{v})\right\}$
-In the case of polytopes, the following lemma (proof in Appendix A.2) reassures us that the suprema in the above expression are attained, and in fact by points in the vertex sets $V_{P},V_{\tilde{P}}$ of the polytopes.
-Lemma 8
-Due to the convexity and compactness of polytopes, we have that
-$H(P,\tilde{P})=\max\left\{\max_{\mathbf{u}\in V_{P}}\mathrm{dist}(\mathbf{u},\tilde{P}),\max_{\mathbf{v}\in V_{\tilde{P}}}\mathrm{dist}(P,\mathbf{v})\right\}$
-In their work, Misiakos et al.
-(2022) used the discrete form of the Hausdorff distance, defined as the Hausdorff distance of the vertex sets of the two polytopes ($DH(P,\tilde{P}):=H(V_{P},V_{\tilde{P}})$), to obtain the following result.
-Proposition 9 (Misiakos et al., 2022)
-Let $p,\tilde{p}\in\mathbb{R}_{\mathrm{max}}[\mathbf{x}]$ be two tropical polynomials with extended Newton polytopes $P=\mathrm{ENewt}(p)$ and $\tilde{P}=\mathrm{ENewt}(\tilde{p})$. Then,
-$\frac{1}{\rho}\max_{\mathbf{x}\in B}|p(\mathbf{x})-\tilde{p}(\mathbf{x})|\leq DH(P,\tilde{P})$
-where $B=\{\mathbf{x}\in\mathbb{R}^{d}:\|\mathbf{x}\|\leq r\}$ and $\rho=\sqrt{r^{2}+1}$.
-We extend the result of Misiakos et al. (2022). The proof is provided in the appendix.
-Proposition 10
-Let $p,\tilde{p}\in\mathbb{R}_{\mathrm{max}}[\mathbf{x}]$ be two tropical polynomials with extended Newton polytopes $P=\mathrm{ENewt}(p)$ and $\tilde{P}=\mathrm{ENewt}(\tilde{p})$. Then,
-$\frac{1}{\rho}\max_{\mathbf{x}\in B}|p(\mathbf{x})-\tilde{p}(\mathbf{x})|\leq H(P,\tilde{P})$
+Theorem 2.
+[…]
+[…]
 where $B=\{\mathbf{x}\in\mathbb{R}^{d}:\|\mathbf{x}\|\leq r\}$ and $\rho=\sqrt{r^{2}+1}$.
 Since $V_{P}\subseteq P,V_{\tilde{P}}\subseteq\tilde{P}$, we have that $\mathrm{dist}(\mathbf{u},\tilde{P})\leq\mathrm{dist}(\mathbf{u},V_{\tilde{P}}),\mathrm{dist}(P,\mathbf{v})\leq\mathrm{dist}(V_{P},\mathbf{v})$.
-The equality in the aforementioned inequalities is achieved only in special cases and thus the bound $H(P,\tilde{P})$ we provide is in general tighter than $DH(P,\tilde{P})$ from Misiakos et al.
+The equality in the aforementioned inequalities is achieved only in special cases and thus the bound H(P,P~)H($\tilde{P}$) we provide is in general tighter than DH(P,P~)DH($\tilde{P}$) from Misiakos et al.
 (2022).
-We demonstrate one example where the two formulations give equal distances and one where the standard Hausdorff distance is strictly smaller than the discrete one.
-Example 2
-Let $P$ be the zonotope formed by taking the generators $(1,0)$ and $(\cos 60^{\circ}=\frac{1}{2},\sin 60^{\circ}=\frac{\sqrt{3}}{2})$.
-Form zonotope $\tilde{P}$ by taking as a generator the mean of the two generators of $P$, which is the vector $(\frac{3}{4},\frac{2+\sqrt{3}}{4})$.
-Refer to Figure 4(a) for a visual representation.
-The discrete Hausdorff distance between the two polytopes is given by:
-$DH(P,\tilde{P})=\max\left\{\frac{\sqrt{3}}{2},\frac{1}{2}\right\}=\frac{\sqrt{3}}{2}.$
-In this case, the Hausdorff distance is equal to the discrete one:
-$H(P,\tilde{P})=\max\left\{\frac{\sqrt{3}}{2},\sin 30^{\circ}\cdot|\text{generator of }\tilde{P}|\right\}$
-$=\max\left\{\frac{\sqrt{3}}{2},\frac{1}{2}\cdot\cos 30^{\circ}=\frac{\sqrt{3}}{4}\right\}=\frac{\sqrt{3}}{2}.$
-Example 3
-Let zonotope $P$ be formed by taking the generators $(1,0)$ and $(\cos 60^{\circ}=\frac{1}{2},\sin 60^{\circ}=\frac{\sqrt{3}}{2})$.
-Form zonotope $\tilde{P}$ by taking as a generator the sum of the two generators, which results in the vector $(\frac{3}{2},\frac{2+\sqrt{3}}{2})$.
-Refer to Figure 4(b) for a visual representation.
-The discrete Hausdorff distance between the two polytopes is calculated as:
-$DH(P,\tilde{P})=\max\{0,1\}=1.$
-Here, the Hausdorff distance is strictly smaller and equal to:
-$H(P,\tilde{P})=\max\left\{0,\frac{1}{2}\right\}=\frac{1}{2}.$
-(a) Example 2
-(b) Example 3
-Figure 4: Polytopes $P$, $\tilde{P}$ of examples 2, 3
-Notice that if our goal were to approximate the polytope $P$ by the polytope $\tilde{P}$, then according to the discrete Hausdorff distance, taking the mean is the optimal strategy and gives distance $\sqrt{3}/2$ (example 2), better than $1$ (example 3).
-On the other hand, according to the Hausdorff distance in its standard form, taking the sum is the optimal strategy and gives distance $1/2$ (example 2), better than $\sqrt{3}/2$ (example 3).
-This implies that an algorithm based on the Hausdorff distance would imply a better approximation of $P$ in the above example.
-We derive the following result, which differs from (Misiakos et al., 2022) by using the standard form of the Hausdorff distance and thus providing a tighter bound.
-For the proof we apply the triangle inequality and then repeatedly Proposition 10 to each output of the network.
-The full proofs of Proposition 10 and Theorem 11 are provided in Appendix A.3.
-Theorem 11
-Let $v$ and $\tilde{v}$ be the outputs of two neural networks as in Figure 3. Then, the following inequality holds:
-$\frac{1}{\rho}\max_{\mathbf{x}\in B}\|v(\mathbf{x})-\tilde{v}(\mathbf{x})\|_{1}\leq\sum_{j=1}^{m}\left(H(P_{j},\tilde{P}_{j})+H(Q_{j},\tilde{Q}_{j})\right),$
+By applying the triangle inequality and Theorem 2 we obtain the following result for two neural networks of compatible input and output dimension.
+Corollary 3.
+Let $\mathbf{v}$ and $\tilde{\mathbf{v}}$ be the outputs of two neural networks as in Figure 1. Then, the following inequality holds:
+$\frac{1}{\rho}\max_{\mathbf{x}\in B}\|\mathbf{v}(\mathbf{x})-\tilde{\mathbf{v}}(\mathbf{x})\|_{1}\leq\sum_{j=1}^{m}\left(H(P_{j},\tilde{P}_{j})+H(Q_{j},\tilde{Q}_{j})\right),$
 where $\|\cdot\|_{1}$ denotes the $L^{1}$ norm.
 4 Compression Algorithm
-In this section we present our algorithm TropNNC for neural network compression.
-Our method, based on Theorem 11, is a structured neural network compression method that reduces the number of neurons in the hidden layers and produces a compressed network that is a functional approximation of the original.
-We differentiate between single-output and multi-output networks and introduce novel approaches for both scenarios.
-The main achievement of our method is to preserve the neural network as a function, achieving performance comparable to state-of-the-art frameworks, all without relying on training data samples.
+From Cor.
+3 it is evident that to compress a ReLU network consisting of a pair of consecutive linear layers, one has to choose compressed weight matrices $\widetilde{\mathbf{A}}\in\mathbb{R}^{K\times(d+1)}$ and $\widetilde{\mathbf{C}}\in\mathbb{R}^{m\times K}$ with $K<n$ such that $\widetilde{P}_{j}\approx P_{j}$ and $\widetilde{Q}_{j}\approx Q_{j}$ for all $j\in[m]$, where the approximation relation is in terms of the Hausdorff distance between the zonotopes.
 4.1 Single output
-First, we focus on single output networks, as the one depicted in Figure 3, with $m=1$.
-The network has weight matrices $\mathbf{A}\in\mathbb{R}^{n\times(d+1)}$, $\mathbf{C}\in\mathbb{R}^{1\times n}$, and a single output $v$ with corresponding positive and negative zonotopes $P$ and $Q$.
-We aim to approximate this network with a smaller one that has fewer neurons in the hidden layer, say $K<n$, resulting in a new hidden layer $\tilde{\mathbf{f}}=(\tilde{f}_{1},\ldots,\tilde{f}_{K})$ with new weight matrices $\widetilde{\mathbf{A}}$ and $\widetilde{\mathbf{C}}$.
-Our goal is to ensure that $\tilde{v}(\mathbf{x})\approx v(\mathbf{x})$ for all $\mathbf{x}\in B$.
-According to Theorem 11, it is sufficient to choose the new weight matrices $\widetilde{\mathbf{A}}$ and $\widetilde{\mathbf{C}}$ such that $\widetilde{P}\approx P$ and $\widetilde{Q}\approx Q$, where the approximation relation is in terms of the Hausdorff distance between the zonotopes.
-In the original network with $n$ hidden neurons, each neuron contributes a generator $|c_{i}|(\mathbf{a}_{i}^{T},b_{i})$, which can be either positive or negative.
-In the approximating network with $K$ hidden neurons, each neuron contributes a generator $|\tilde{c}_{i}|(\tilde{\mathbf{a}}_{i}^{T},\tilde{b}_{i})$.
-We aim to pick generators $|\tilde{c_{i}}|(\tilde{\mathbf{a}}_{i}^{T},\tilde{b}_{i})$ such that the generated positive and negative zonotopes $\widetilde{P}$ and $\widetilde{Q}$ approximate the positive and negative zonotopes $P$ and $Q$, respectively.
-In essence, we have translated the neural network approximation problem into a zonotope approximation problem—specifically, the task of approximating a zonotope with another zonotope that has fewer generators.
-This problem is known in the literature as zonotope order reduction (Yang and Scott, 2018).
+If we have $m=1$ then the problem can readily be translated into a zonotope approximation problem—the task of approximating a zonotope with another zonotope that has fewer generators.
+This problem is known as zonotope order reduction (Yang and Scott, 2018).
 In our case, the approximation must happen in terms of the Hausdorff distance.
-5.1.1 Algorithm for single output network
-Misiakos et al.
-(2022) proposed the algorithm ”Zonotope K-means” for compressing single-output neural networks.
-Their approach uses K-means clustering on the set of positive and negative generators, and replaces the generators of each formed cluster by a single representative, the center of the cluster (i.e.
-the mean of the generators of the cluster).
-In contrast, our approach also uses K-means to cluster the generators, but instead of taking the representative to be the center of the cluster, we take the representative to be the sum of the generators of the cluster.
-We show later that this compression produces a tighter approximation of the neural network compared to (Misiakos et al., 2022).
-We demonstrate this with Example 4 and prove this formally with Proposition 12 and Corollary 13.
-Our Algorithm 1 is depicted below.
+Algorithm for single output network.
+For single-output networks, our approach uses K-Means to cluster positive and negative generators, and replaces each cluster with a single representative.
+Unlike Misiakos et al.
+(2022), who define the representative as the center/mean of the cluster, we take the representative to be the sum of the cluster.
+Our Algorithm 1 is depicted below. Figure 2 illustrates an example execution of TropNNC compared to the method of Misiakos et al. (2022). For our algorithm, the following bound and its corollary hold.
 Algorithm 1 TropNNC, Single output
 1: Split the generators $|c_{i}|(\mathbf{a}_{i}^{T},b_{i})$ into positive and negative generators:
-[…]
-[…]
+$\{$ $|c_{i}|(\mathbf{a}_{i}^{T},b_{i}):c_{i}>0\},$
+$\{$ $|c_{i}|(\mathbf{a}_{i}^{T},b_{i}):c_{i}<0\}.$
 2: Execute K-means with $K/2$ centers on the positive generators $\{|c_{i}|(\mathbf{a}_{i}^{T},b_{i}):c_{i}>0\}$, and $K/2$ centers on the negative generators $\{|c_{i}|(\mathbf{a}_{i}^{T},b_{i}):c_{i}<0\}$.
 3: Obtain positive and negative cluster representatives:
-[…]
-[…]
+$\{$ $|\tilde{c}_{i}|(\tilde{\mathbf{a}}_{i}^{T},\tilde{b}_{i}):i\in C_{+}\},$
+$\{$ $|\tilde{c}_{i}|(\tilde{\mathbf{a}}_{i}^{T},\tilde{b}_{i}):i\in C_{-}\},$
 where $C_{+}\sqcup C_{-}=[K]$ and $|\tilde{c}_{i}|(\tilde{\mathbf{a}}_{i}^{T},\tilde{b}_{i})$ is the sum of the generators of cluster $i$.
 4: For each $i\in[K]$ construct a (hidden layer) neuron with input weights and bias $|\tilde{c}_{i}|(\tilde{\mathbf{a}}_{i}^{T},\tilde{b}_{i})$.
-5:  For each constructed neuron $i$ set the output weight to $1$ if the neuron corresponds to a representative of a positive cluster ($i\in C_{+}$), otherwise set it to -1.
-Example 4
-Suppose we have a positive zonotope consisting of 3 positive generators $g_{1},g_{2},g_{3}$ like the pink zonotope (hexagon) in Figure 5(a), and suppose we have $K/2=2$.
-The Zonotope K-means algorithm of Misiakos et al.
-(2022) will form 2 clusters of generators, the black cluster which consists of a single generator $g_{1}$, and the pink cluster which consists of the pink generators $g_{2},g_{3}$.
-Zonotope K-means will take as cluster representatives the centers of the clusters.
-In our example, this will be the black generator $\tilde{g}_{1}=g_{1}$ for the black cluster, and the blue generator $\tilde{g}_{2}=\frac{g_{2}+g_{3}}{2}$ for the pink cluster, as in Figure 5(b).
-The blue zonotope is the zonotope generated by the representatives of the clusters (black and blue generators $\tilde{g}_{1},\tilde{g}_{2}$ respectively), and it is the approximation that Zonotope K-means uses to approximate the pink zonotope.
-In contrast, TropNNC will also form the same 2 clusters of generators, but instead of taking the center, it will take the sum of the generators of each cluster as the cluster representative.
-In our example, these will be the black generator $\tilde{g}_{1}=g_{1}$ and the blue generator $\tilde{g}_{2}=g_{2}+g_{3}$ of Figure 5(c), which form a different blue zonotope which is used as the approximation of the pink zonotope.
-In this example, one can visually verify that according to the Hausdorff distance, the approximation of TropNNC is better than the approximation of Zonotope K-Means.
-(a) Original Zonotope
-(b) Zonotope K-means
+5: For each constructed neuron $i$ set the output weight to 11 if the neuron corresponds to a representative of a positive cluster ($i\in C_{+}$), otherwise set it to -1.
+(a) Original
+(b) Misiakos et al.
 (c) TropNNC
-Figure 5: Example execution of Zonotope K-means and TropNNC
-The result of Example 4 can be generalized. Indeed, the following bound and its corollary hold (proofs in Appendix A.4 and A.5).
-Proposition 12
+Figure 2: Example execution of TropNNC compared to the method of Misiakos et al. (2022).
+Proposition 4.
 Suppose the clusters K are enough so that for every cluster, no 2 generators of the cluster form an obtuse angle.
 Then, single-output TropNNC produces a neural network with output $\tilde{v}$ satisfying:
-$\frac{1}{\rho}\max_{\mathbf{x}\in B}|u(\mathbf{x})-\tilde{v}(\mathbf{x})|\leq\sum_{i\in I}\min\{||c_{i}(\mathbf{a}_{i}^{T},b_{i})||,\delta_{max}\},$
+[…]
 where $\delta_{max}$ is the largest distance from a point to its corresponding cluster center.
-Corollary 13
+Corollary 5.
 The above bound is tighter than the bound of Zonotope K-means of Misiakos et al. (2022).
 4.2 Multi-output
-Now we consider the case of multi-output networks as in Figure 3, for general $m\in\mathbb{N}$.
-We consider a network with output $\mathbf{v}=(v_{1},\ldots,v_{m})$ and corresponding positive and negative zonotopes $P_{j}$ and $Q_{j}$ for each output node $v_{j}$.
-An interesting property of these polytopes is that they share the directions $(\mathbf{a}_{i}^{T},b_{i})$ of their generators.
+We now consider the multi-output case with $m\in\mathbb{N}$.
+Notice an interesting property of the zonotopes $P_{j},Q_{j}$: they share the directions $(\mathbf{a}_{i}^{T},b_{i})$ of their generators.
 For instance, output $v_{1}$ might have a positive generator $|c_{1i}|(\mathbf{a}_{i}^{T},b_{i})$ of zonotope $P_{1}$, while output $v_{2}$ might have a negative generator $|c_{2i}|(\mathbf{a}_{i}^{T},b_{i})$ of zonotope $Q_{2}$.
 These generators are parallel to each other, with common direction $(\mathbf{a}_{i}^{T},b_{i})$.
-As before, we aim to approximate this network with a smaller network that has fewer neurons in the hidden layer, say $K<n$, resulting in a new hidden layer $\tilde{\mathbf{f}}=(\tilde{f}_{1},\ldots,\tilde{f}_{K})$ with new weight matrices $\widetilde{\mathbf{A}}$ and $\widetilde{\mathbf{C}}$.
-Our goal is to ensure that $\tilde{\mathbf{v}}(\mathbf{x})\approx\mathbf{v}(\mathbf{x})$ for all $\mathbf{x}\in B$.
-As we have seen, this can be translated to having $\widetilde{P}_{j}\approx P_{j}$ and $\widetilde{Q}_{j}\approx Q_{j}$, where the approximation relation is in terms of the Hausdorff distance between the zonotopes.
-Notice that for each new neuron $\tilde{f}_{i}$, its input weight-bias $(\tilde{\mathbf{a}}_{i}^{T},\tilde{b}_{i})$ defines a generator direction for every output $\tilde{v}_{j}$, where $j\in[m]$.
-For this generator direction, the magnitudes of the generators change across different outputs by adjusting the output weight $\tilde{c}_{ji}$ for each $j\in[m]$, resulting in parallel generators for each output: $|\tilde{c}_{1i}|(\tilde{\mathbf{a}}_{i}^{T},\tilde{b}_{i}),\ldots,|\tilde{c}_{mj}|(\tilde{\mathbf{a}}_{i}^{T},\tilde{b}_{i})$.
-These generators can be either positive or negative according to the sign of $\tilde{c}_{ji}$.
-Thus, we start with original positive and negative zonotopes, with parallel generators, and aim to approximate them with new positive and negative zonotopes, also with parallel generators.
+Hence, our original positive and negative zonotopes have parallel generators, and we aim to approximate them with new positive and negative zonotopes, also with parallel generators.
 We refer to this complex approximation problem as simultaneous zonotope approximation.
-Example 5
-Suppose we have a neural network with a single hidden layer, as in Figure 3, with dimensions $d+1=n=m=2$.
-Consider input weights $(\mathbf{a}_{1}^{T},b_{1})=(1,0),(\mathbf{a}_{2}^{T},b_{2})=(0,1)$ and output weights $c_{11}=3,c_{12}=5,c_{21}=4,c_{22}=2$.
-In this example, for simplicity we took all output weights to be positive so that we only deal with positive zonotopes.
-The zonotopes of the two outputs will be two parallelograms with parallel edges, as illustrated in the Figure 6.
-The zonotope of the first output is generated by $c_{11}(\mathbf{a}_{1}^{T},b_{1})=3(1,0)=(3,0)$ and $c_{12}(\mathbf{a}_{2}^{T},b_{2})=5(0,1)=(0,5)$, and of the second output by $c_{21}(\mathbf{a}_{1}^{T},b_{1})=4(1,0)=(4,0)$ and $c_{22}(\mathbf{a}_{2}^{T},b_{2})=2(0,1)=(0,2)$.
-Say we want to reduce the hidden neurons to $K=1,\tilde{\mathbf{f}}=(\tilde{f_{1}})$.
-If we could approximate each output’s zonotope separately, we could simply apply the single output algorithm and approximate each parallelogram by its diagonal.
-However, these diagonals are not parallel to each other, and thus can not occur by a single hidden neuron $\tilde{f}_{1}$ with input weights $(\tilde{\mathbf{a}}_{1}^{T},\tilde{b}_{1})$.
-Instead, we have to choose a single common direction $(\tilde{\mathbf{a}}_{1}^{T},\tilde{b}_{1})$ for both output zonotopes.
-We can however choose a different magnitude for each output along this common direction.
-As will be presented in the algorithm below, for the common direction we choose the vector $(\tilde{\mathbf{a}}_{i}^{T},\tilde{b}_{i})=\frac{(\mathbf{a}_{1}^{T},b_{1})+(\mathbf{a}_{2}^{T},b_{2})}{2}=(0.5,0.5)$.
-For each output $j$, $\tilde{c}_{j1}$ is chosen so that the edge is as close to the diagonal as possible.
-Specifically, we choose $\tilde{c}_{11}=3+5=8$ and $\tilde{c}_{21}=2+4=6$.
-The approximation procedure can be seen in Figure 6(c).
+Example 1.
+[…]
 (a) Zonotope of $1^{st}$ output
 (b) Zonotope of $2^{nd}$ output
-(c) Simultaneous zonotope approximation
+(c) Sim. zonotope approximation
 Figure 3: Example of simultaneous zonotope approximation for a network with 2 outputs and 2 hidden neurons
-5.2.1 Non-iterative Algorithm for multi-output network
-To tackle this problem of simultaneous zonotope approximation, Misiakos et al.
-(2022) proposed the ”Neural Path K-means” algorithm.
-Their approach to the problem is to use K-means clustering on the set of vectors $\left\{(\mathbf{a}_{i}^{T},b_{i},\mathbf{C}_{:,i}^{T}),i\in[n]\right\}$, and replace the vectors of each formed cluster by a single representative, the center of the cluster (i.e.
-the mean of the vectors of the cluster).
-Our approach also uses K-means to cluster the same vectors, but instead of taking the representative to be the center of the cluster, we form the representative using the following heuristic:
-Suppose you have 2 neurons with input weights $(\mathbf{a}_{1}^{T},b_{1}),(\mathbf{a}_{2}^{T},b_{2})$ and output weights $\mathbf{C}_{:,1}^{T},\mathbf{C}_{:,2}^{T}$, that get clustered together.
-If $(\mathbf{a}_{1}^{T},b_{1})\approx(\mathbf{a}_{2}^{T},b_{2})\approx\frac{(\mathbf{a}_{1}^{T},b_{1})+(\mathbf{a}_{2}^{T},b_{2})}{2}$, then for every output $j$ the following holds:
-$C_{j,1}(\mathbf{a}_{1}^{T},b_{1})+C_{j,2}(\mathbf{a}_{2}^{T},b_{2})\approx(C_{j,1}+C_{j,2})\frac{(\mathbf{a}_{1}^{T},b_{1})+(\mathbf{a}_{2}^{T},b_{2})}{2}.$
-Thus, in order to approximate the 2 neurons, it suffices to take the mean of their input weights and the sum of their output weights.
-This heuristic can also be motivated by viewing at the network: it corresponds to taking the mean in terms of the input activations of the two neurons, and taking the sum as the joint output activations of the two neurons.
-We generalize the above to clusters with more neurons.
-For every cluster $k\in[K]$ with clustered neuron indexes $I_{k}$ and vectors $\left\{(\mathbf{a}_{i}^{T},b_{i},\mathbf{C}_{:,i}^{T}),i\in I_{k}\right\}$ take $(\tilde{\mathbf{a}}_{k}^{T},\tilde{b}_{k})$ to be the mean of $\left\{(\mathbf{a}_{i}^{T},b_{i}),i\in I_{k}\right\}$, take $\widetilde{\mathbf{C}}_{:,k}^{T}$ to be the sum of $\left\{\mathbf{C}_{:,i}^{T},i\in I_{k}\right\}$, and form the representative of the cluster $(\tilde{\mathbf{a}}_{k}^{T},\tilde{b}_{k},\widetilde{\mathbf{C}}_{:,k}^{T})$.
-This leads to Algorithm 2.
+4.2.1 Non-iterative Algorithm for multi-output network.
+For the multi-output case, we perform clustering of similar neurons based on the set of clustering vectors $\left\{(\mathbf{a}_{i}^{T},b_{i},\mathbf{C}_{:,i}^{T}),i\in[n]\right\}$, and replace each cluster with a single representative.
+Unlike Misiakos et al.
+(2022), who take the cluster center/mean as the representative, we form the representative as follows: For every cluster $k\in[K]$ with clustered neuron indexes $I_{k}$ and vectors $\left\{(\mathbf{a}_{i}^{T},b_{i},\mathbf{C}_{:,i}^{T}),i\in I_{k}\right\}$ take $(\tilde{\mathbf{a}}_{k}^{T},\tilde{b}_{k})$ to be the mean of $\left\{(\mathbf{a}_{i}^{T},b_{i}),i\in I_{k}\right\}$, take $\widetilde{\mathbf{C}}_{:,k}^{T}$ to be the sum of $\left\{\mathbf{C}_{:,i}^{T},i\in I_{k}\right\}$, and form the representative of the cluster $(\tilde{\mathbf{a}}_{k}^{T},\tilde{b}_{k},\widetilde{\mathbf{C}}_{:,k}^{T})$.
+The complete procedure is shown in Algorithm 2.
 Algorithm 2 (Non-iterative) TropNNC for Multi-output networks
-1:  Execute KMeans with $K$ centers on the vectors $(\mathbf{a}_{i}^{T},b_{i},\mathbf{C}_{:,i}^{T})$ for $i\in[n]$, forming $K$ clusters $\{(\mathbf{a}_{i}^{T},b_{i},\mathbf{C}_{:,i}^{T})\mid i\in I_{k}\}$ for $k\in[K]$.
+1: Execute clustering on the clustering vectors $(\mathbf{a}_{i}^{T},b_{i},\mathbf{C}_{:,i}^{T})$ for $i\in[n]$, forming $K$ clusters $\{(\mathbf{a}_{i}^{T},b_{i},\mathbf{C}_{:,i}^{T})\mid i\in I_{k}\}$ for $k\in[K]$.
 2:For each $k\in[K]$, form the cluster representative $(\tilde{\mathbf{a}}_{k}^{T},\tilde{b}_{k},\widetilde{\mathbf{C}}_{:,k}^{T})$ as follows:
@@
     For the output weights, set the $k$-th column to $\widetilde{\mathbf{C}}_{:,k}$.
-There are several variations of Algorithm 2.
-For instance, one could use the sum instead of the mean for the input weights.
-However, in our experiments, this approach yielded worse results.
-Another variation involves clustering neurons based on cosine similarity rather than the $L^{2}$ distance of the vectors, recognizing that neurons with parallel input weights can be clustered together regardless of their output weights.
-Additionally, one could normalize the vectors based on their input weights by shifting the norm to the output weights before applying K-means.
-Despite testing these methods, none provided significant advantages.
-Therefore, we simply recommend Algorithm 2.
-5.2.2 Iterative Algorithm for multi-output network
-To improve the approximation of Algorithm 2, we again make use of tropical geometry.
-Specifically, we formulate an optimization problem that takes the output of Algorithm 2 and with an iterative process produces weights that achieve a better simultaneous zonotope approximation.
-Motivated by Algorithm 1, assuming the number of null neurons are few (see Misiakos et al.
-(2022) and Appendix A.6), we wish in terms of every output $j$ the cluster representative $\widetilde{C}_{j,k}(\tilde{\mathbf{a}}_{k}^{T},\tilde{b}_{k})$ to be as close to the cluster sum $\sum_{i\in I_{k}}C_{j,i}(\mathbf{a}_{i}^{T},b_{i})$ as possible.
+4.2.2 Iterative Algorithm for multi-output network.
+To improve the approximation of Algorithm 2, we formulate an optimization problem that takes the output of Algorithm 2 and with an iterative process produces weights that achieve a better simultaneous zonotope approximation.
+Motivated by Algorithm 1, assuming the number of null neurons are few (see appendix), we want in terms of every output $j$ the cluster representative $\widetilde{C}_{j,k}(\tilde{\mathbf{a}}_{k}^{T},\tilde{b}_{k})$ to be a close approximation to the cluster sum $\sum_{i\in I_{k}}C_{j,i}(\mathbf{a}_{i}^{T},b_{i})$.
 Thus, for every cluster $k$, we have unknowns $(\tilde{\mathbf{a}}_{k}^{T},\tilde{b}_{k},\widetilde{\mathbf{C}}_{:,k}^{T})=(\tilde{a}_{k1},\ldots,\tilde{a}_{kd},\tilde{b}_{k},\widetilde{C}_{1k},\ldots)$, and we wish to find a solution which minimizes the following criterion:
-$\sum_{j=1}^{m}\left\|\tilde{C}_{jk}(\tilde{\mathbf{a}}_{k}^{T},\tilde{b}_{k})-\sum_{i\in I_{k}}C_{ji}(\mathbf{a}_{i}^{T},b_{i})\right\|^{2},$
+$\sum_{j=1}^{m}\left\|\tilde{C}_{jk}(\tilde{\mathbf{a}}_{k}^{T},\tilde{b}_{k})-\sum_{i\in I_{k}}C_{ji}(\mathbf{a}_{i}^{T},b_{i})\right\|^{2},$ (1)
 where $m$ is the number of outputs, and $I_{k}$ is the set of neurons of cluster $k$.
@@
 1.
-Holding the input weights $(\tilde{\mathbf{a}}_{k}^{T},\tilde{b}_{k})$ constant and minimizing with respect to $\widetilde{\mathbf{C}}_{:,k}$, we notice that the terms of the sum are independent and thus can be minimized separately.
-For each term of the sum, the minimization occurs if we project the sum $\sum_{i\in I_{k}}C_{ji}(\mathbf{a}_{i}^{T},b_{i})$ onto $(\tilde{\mathbf{a}}_{k}^{T},\tilde{b}_{k})$.
+Fixing the input weights $(\tilde{\mathbf{a}}_{k}^{T},\tilde{b}_{k})$ and minimizing with respect to $\widetilde{\mathbf{C}}_{:,k}$, the terms of the sum are independent and thus can be minimized separately.
+For each term of the sum, the optimal error occurs if we project the sum $\sum_{i\in I_{k}}C_{ji}(\mathbf{a}_{i}^{T},b_{i})$ onto $(\tilde{\mathbf{a}}_{k}^{T},\tilde{b}_{k})$.
 We have:
-$\widetilde{C}_{jk}=\frac{\left\langle\sum_{i\in I_{k}}C_{ji}(\mathbf{a}_{i}^{T},b_{i}),(\tilde{\mathbf{a}}_{k}^{T},\tilde{b}_{k})\right\rangle}{\left\|(\tilde{\mathbf{a}}_{k}^{T},\tilde{b}_{k})\right\|^{2}}$
+$\widetilde{C}_{jk}=\frac{\left\langle\sum_{i\in I_{k}}C_{ji}(\mathbf{a}_{i}^{T},b_{i}),(\tilde{\mathbf{a}}_{k}^{T},\tilde{b}_{k})\right\rangle}{\left\|(\tilde{\mathbf{a}}_{k}^{T},\tilde{b}_{k})\right\|^{2}}$ (2)
 2.
-Holding the output weights $\widetilde{\mathbf{C}}_{:,k}$ constant, and minimizing with respect to the input weights $(\tilde{\mathbf{a}}_{k}^{T},\tilde{b}_{k})$, we take the derivative of the criterion with respect to $(\tilde{\mathbf{a}}_{k}^{T},\tilde{b}_{k})$ and set it to zero.
+Fixing the output weights $\widetilde{\mathbf{C}}_{:,k}$, and minimizing with respect to the input weights $(\tilde{\mathbf{a}}_{k}^{T},\tilde{b}_{k})$, we take the derivative of criterion (1) with respect to $(\tilde{\mathbf{a}}_{k}^{T},\tilde{b}_{k})$ and set it to zero.
 We have:
@@
 $\Leftrightarrow(\tilde{\mathbf{a}}_{k}^{T},\tilde{b}_{k})\sum_{j=1}^{m}\widetilde{C}_{jk}^{2}=\sum_{j=1}^{m}\sum_{i\in I_{k}}C_{ji}(\mathbf{a}_{i}^{T},b_{i})\widetilde{C}_{jk}\Leftrightarrow$
-$(\tilde{\mathbf{a}}_{k}^{T},\tilde{b}_{k})=\frac{\sum_{j=1}^{m}\sum_{i\in I_{k}}C_{ji}(\mathbf{a}_{i}^{T},b_{i})\widetilde{C}_{jk}}{\sum_{j=1}^{m}\widetilde{C}_{jk}^{2}}$
-We iterate over the above alternating steps for num_iter iterations.
-Additionally, we can initialize the iteration with the representative obtained from Algorithm 2, which should provide a good starting point.
-This initialization is expected to reduce the number of required epochs and speed up the compression algorithm.
+$(\tilde{\mathbf{a}}_{k}^{T},\tilde{b}_{k})=\frac{\sum_{j=1}^{m}\sum_{i\in I_{k}}C_{ji}(\mathbf{a}_{i}^{T},b_{i})\widetilde{C}_{jk}}{\sum_{j=1}^{m}\widetilde{C}_{jk}^{2}}$ (3)
+We can initialize the iteration with the representative obtained from Algorithm 2.
 The resulting procedure is detailed in Algorithm 3.
-We should note that, in practise, the number of null generators is not negligible, and thus our optimization criterion also constitutes a heuristic method.
+We note that, in practice, the number of null generators is not negligible, and thus criterion (1) also constitutes a heuristic method.
 For this reason, the number of iterations should not be excessive.
-We provide the following bound for the approximation error of our algorithm (proof in Appendix A.6).
+We provide Proposition 6 for the approximation error of our algorithm.
 Algorithm 3 Iterative TropNNC for Multi-output networks
-1:  Execute K-means with K centers on the vectors $(\mathbf{a}_{i}^{T},b_{i},\mathbf{C}_{:,i}^{T}),i\in[n]$, and form K clusters $\{(\mathbf{a}_{i}^{T},b_{i},\mathbf{C}_{:,i}^{T}),i\in I_{k}\},k\in[K]$
-2:  for $k\in[K]$ do
-3:     Form the initial cluster representative $(\tilde{\mathbf{a}}_{k}^{T},\tilde{b}_{k},\widetilde{\mathbf{C}}_{:,k}^{T})$ as follows:
-4:      2i.
-Take $(\tilde{\mathbf{a}}_{k}^{T},\tilde{b}_{k})$ to be the mean of the input weights + bias of the vectors of the cluster, i.e., the mean of the set of vectors $\{(\mathbf{a}_{i}^{T},b_{i}),i\in I_{k}\}$
-5:      2ii. Take $\widetilde{\mathbf{C}}_{:,k}^{T}$ to be the sum of the output weights of the vectors of the cluster, i.e., the sum of the set of vectors $\{\mathbf{C}_{:,i}^{T},i\in I_{k}\}$
+1: Form the initial cluster representatives $(\tilde{\mathbf{a}}_{k}^{T},\tilde{b}_{k},\widetilde{\mathbf{C}}_{:,k}^{T})$ by executing the non-iterative algorithm.
+2:for $\text{iter}=1\text{ to }num\_iter$ do
+3:  for $k\in[K]$ do
+4:   for $j\in[m]$ do
+5:    $\widetilde{C}_{jk}\leftarrow\frac{\langle\sum_{i\in I_{k}}C_{ji}(\mathbf{a}_{i}^{T},b_{i}),(\tilde{\mathbf{a}}_{k}^{T},\tilde{b}_{k})\rangle}{\|(\tilde{\mathbf{a}}_{k}^{T},\tilde{b}_{k})\|^{2}}$
 6:   end for
-7:  for iter=1 to $\text{iter}=1\text{ to }num\_iter$ do
-8:     for $k\in[K]$ do
-9:        for $j\in[m]$ do
-10:           $\widetilde{C}_{jk}\leftarrow\frac{\langle\sum_{i\in I_{k}}C_{ji}(\mathbf{a}_{i}^{T},b_{i}),(\tilde{\mathbf{a}}_{k}^{T},\tilde{b}_{k})\rangle}{\|(\tilde{\mathbf{a}}_{k}^{T},\tilde{b}_{k})\|^{2}}$
-11:        end for
-12:        $(\tilde{\mathbf{a}}_{k}^{T},\tilde{b}_{k})\leftarrow\frac{\sum_{j=1}^{m}\sum_{i\in I_{k}}C_{ji}(\mathbf{a}_{i}^{T},b_{i})\widetilde{C}_{jk}}{\sum_{j=1}^{m}\widetilde{C}_{jk}^{2}}$
-13:     end for
-14:  end for
-15:  Construct the new linear layer:
-16:   For the input weights, the k-th row of weight-bias matrix is $(\tilde{\mathbf{a}}_{k}^{T},\tilde{b}_{k})$
-17:   For the output weights, the k-th column is $\widetilde{\mathbf{C}}_{:,k}^{T}$
-Proposition 14
+7:   $(\tilde{\mathbf{a}}_{k}^{T},\tilde{b}_{k})\leftarrow\frac{\sum_{j=1}^{m}\sum_{i\in I_{k}}C_{ji}(\mathbf{a}_{i}^{T},b_{i})\widetilde{C}_{jk}}{\sum_{j=1}^{m}\widetilde{C}_{jk}^{2}}$
+8:  end for
+9:end for
+10: Construct the new linear layer as in the non-iterative algorithm.
+Proposition 6.
 Suppose the clusters K are enough so that for every cluster, no two $(\mathbf{a}_{i}^{T},b_{i}),(\mathbf{a}_{i^{\prime}}^{T},b_{i^{\prime}})$ of the cluster form an obtuse angle.
-Then, a variant (see Appendix) of Iterative TropNNC produces a neural network with output $\tilde{v}$ satisfying:
-$\frac{1}{\rho}\max_{\mathbf{x}\in B}\|v(\mathbf{x})-\tilde{v}(\mathbf{x})\|_{1}\leq\sqrt{m}\sum_{i=1}^{n}\min\left\{\|\mathbf{C}_{:,i}\|\|(\mathbf{a}_{i}^{T},b_{i})\|,\frac{l_{k(i)}}{N_{min}}+\|\epsilon_{:,i}\|_{F}\right\}\\ +\sum_{j=1}^{m}\sum_{i\in N_{j}}|c_{ji}|\|(\mathbf{a}_{i}^{T},b_{i})\|$
+Then, a variant of Iterative TropNNC produces a neural network with output $\tilde{v}$ satisfying:
+$\frac{1}{\rho}\max_{\mathbf{x}\in B}\|\mathbf{v}(\mathbf{x})-\tilde{\mathbf{v}}(\mathbf{x})\|_{1}\leq\\ \sqrt{m}\sum_{i=1}^{n}\min\left\{\|\mathbf{C}_{:,i}\|\|(\mathbf{a}_{i}^{T},b_{i})\|,\frac{l_{k(i)}}{N_{min}}+\|\epsilon_{:,i}\|_{F}\right\}\\ +\sum_{j=1}^{m}\sum_{i\in N_{j}}|c_{ji}|\|(\mathbf{a}_{i}^{T},b_{i})\|$
 where:
@@
   • $\|\cdot\|_{F}$ is the Frobenius norm.
-6 Extensions, Limitations, and Applications of the Method
-Before proceeding to our experiments, we would like to discuss the applicability of our algorithm, highlight some limitations, and explore potential strategies to overcome these challenges.
+Heuristic Improvements
+We obtain an advantage by dropping the bias from the clustering vectors – which leads to better upper hull approximations – and normalizing the vectors used for clustering.
+These two heuristics are discussed in detail in the appendix.
 Deep and Convolutional Networks.
-Our methods were analyzed previously for the case of networks with a single hidden linear layer with ReLU activations.
-However, they can be generalized both for deep and convolutional neural networks.
-In the first case, where we have a deep network with multiple layers, we can view each layer as an instance for our algorithm, by having an input from the previous layer and providing an output to the next layer.
-Therefore, we can recursively apply our multi-output algorithm to compress each hidden layer of the network.
-In the case of convolutional networks, our methods can be extended for convolutional layers with ReLU and maxout or average pooling layers in between them.
-Here, instead of neurons, we have channels, and instead of weights that connect the neurons, we have kernels that connect the channels.
-For each hidden layer, the input and output weight matrices $\mathbf{A}$ and $\mathbf{C}$ are 4-dimensional, where each element of the row $(\mathbf{a}_{i}^{T},b_{i})$ and the column $\mathbf{C}_{:,i}$ is a kernel.
-Compressing a convolutional layer in a structured manner involves clustering channels.
-To apply our multi-output algorithm, we first flatten the matrices $\mathbf{A}$ and $\mathbf{C}$ so that they become 2-dimensional, by unravelling the kernels row-wise for $\mathbf{A}$ and column-wise for $\mathbf{C}$.
-We then apply our multi-output algorithm on the flattened matrices.
-Finally, after obtaining the new $\widetilde{\mathbf{A}}$ and $\widetilde{\mathbf{C}}$, we reshape the matrices to restore their original dimensions as convolutional layer weight matrices.
-We emphasize that we are the first to apply a tropical geometrical method to convolutional layers.
-Batch normalization and other limitations.
-Our algorithm can not be applied in the case of networks with batch normalization layers in between the convolutional and linear layers.
-The reason for this is that batch normalization layers have a non-negligible effect on the zonotopes of the network we approximate.
-Here, we explain two methodologies that can be used to adapt networks with batch normalization to be applicable for our algorithm.
-These methodologies were used in the experiments to convert the VGG and ResNet models to equivalent models that our algorithm can be applied to.
-We propose two methodologies that allow most networks with batch normalization layers to be expressed equivalently as networks without batch normalization.
-As a first idea, we can fuse together a pair (convolutional (or linear) layer, batch normalization layer) into a single convolutional layer as follows: We take the pair we wish to fuse and form a network with just these two layers.
-We form a second network with a single, new convolutional layer.
-We then train the second network to give the same output as the first network on random noise input.
-This should produce a new convolutional layer that approximates the pair (convolutional (or linear) layer, batch normalization layer) with very little accuracy loss.
-Finally, we replace each such pair with its approximate convolutional (or linear) layer, which gives us an approximately equivalent network without batch normalization.
-An alternative second method is to shift the computations of the normalization layer directly to the convolutional or linear layer.
-Specifically, we can view the batch normalization layer being applied as a multiplication with corrective terms and addition with an extra bias term.
-We can incorporate these computations into the weights and bias of the convolutional or linear layer.
-Using these ideas, we can fuse a batch normalization layer with its preceding convolutional or linear layer.
-This enables two approaches for applying our algorithm to networks with batch normalization.
-The first approach involves fusing all layers of the network beforehand, resulting in a fully fused network without batch normalization.
-The second approach performs fusion during the compression process, applying it layer-by-layer as each layer is compressed.
-For the second approach, we made a notable observation during our experiments, although the reasoning behind it remains unclear: using the pre-fusion weights/kernels of the layers for clustering vectors consistently led to significantly better results.
-As for other limitations, we should note that our algorithm is designed to work strictly on layers as described above.
-Namely, convolutional or linear layers with ReLU and maxout or average pooling layers in between.
-Layers that apply computations, like batch normalization layers or the loops of residual blocks of ResNet networks fundamentally change the theory for the analysis of the network.
-For example, similarly to ThiNet, our methods can not be applied for the compression of the terminal convolutional layer of residual blocks.
+For deep networks, we can view each hidden layer of the network as an instance for our algorithm, having an input from the previous layer and providing an output to the next layer.
+Hence, we can recursively apply our multi-output algorithm to compress each hidden layer.
+For convolutional networks, we unravel the kernels of the weight tensors row-wise for $\mathbf{A}$ and column-wise for $\mathbf{C}$, apply our algorithm, and reshape them back to 4D tensors.
+If we also have batch normalization, then this can be dealt with by fusing the operations of batch normalization into the preceding convolutional or linear layer.
 Non-uniform compression.
-The techniques we have presented apply a uniform pruning ratio to all layers of the network.
-However, they can easily be extended to non-uniform pruning if instead of K-means we use hierarchical clustering with a global threshold parameter, like in the CUP framework (Duggal et al., 2021).
-In fact, when used with hierarchical clustering, TropNNC differs from CUP in step 1, where we choose a different approach to build the clustering vectors (or filter features as Duggal et al.
+In Algorithm 2 we intentionally did not mention how clustering is performed.
+We can either use K-Means with a constant pruning ratio for all layers, or we can use hierarchical clustering with a global threshold parameter, like in the CUP framework.
+When used with hierarchical clustering, TropNNC differs from CUP in step 1, where we choose a different approach to build the clustering vectors (or filter features as Duggal et al.
 (2021) call them) of a convolutional layer, and more importantly in step 3, where we choose a different cluster representative based on tropical geometry.
-We also propose a modification to step 2 of CUP, specifically the hierarchical clustering step using a global distance threshold.
+We also propose a modification to step 2 of CUP.
 Since the clustering vectors of different layers have varying dimensions, and vectors in higher-dimensional spaces tend to be more spread out, we introduce two variants for selecting the distance threshold for each layer:
@@
   • Variant 2: For each layer, take the distance threshold to be some global constant times the mean of the norms of the clustering vectors of the layer.
-As demonstrated in our experiments, the first variant excels when minimizing the number of parameters, while the second variant is more effective when our objective is to minimize floating-point operation demands.
+Limitations.
+Limitations are discussed in the appendix.
+5 Related Work
+Our work falls under the category of structured pruning (He and Xiao, 2024).
+Here, approaches include data-free norm-based pruning (Li et al., 2017; He et al., 2019), KL-divergence based pruning (Luo and Wu, 2020), data-driven activation-based pruning (Lin et al., 2020; Sui et al., 2021), and data-driven filter pruning based on reconstruction loss minimization.
+Within the latter category, ThiNet (Luo et al., 2017) is a representative method that greedily removes filters contributing least to the next layer’s input, based on a data-driven reconstruction objective.
+Our approach resembles ThiNet in its structured layer-wise compression, but unlike ThiNet, it is data-free.
+Although layer-wise pruning is increasingly being replaced by global strategies – such as NISP (Yu et al., 2018) and SPvR (Hassan et al., 2025), which consider the importance of filters relative to the final layer – ThiNet remains a strong structured data-driven baseline to which we compare our method.
+More recently, data-free methods have been proposed.
+For example, CUP (Duggal et al., 2021) hierarchically clusters similar channels and selects representatives by $L^{1}$ norm.
+Our method builds on the same intuition of redundancy reduction via clustering, but introduces three key extensions: (i) modifying the clustering vectors, (ii) adapting the distance thresholds of hierarchical clustering to layer dimensions, and (iii) using tropical geometry to select representatives.
+We built on Misiakos et al.
+(2022), who used a discrete form of the Hausdorff distance to geometrically bound the error between tropical polynomials, framing network reduction as a zonotope order reduction problem, and proposing a basic clustering-based compression method.
+Their method clusters similar neurons and replaces each cluster with the representative obtained by taking the mean of the input and output weights of the cluster.
+We refined their central theorem using the standard Hausdorff distance and introduce a novel algorithm that formulates compression as a simultaneous zonotope approximation problem.
+Our method makes a better initial estimation of the compressed network’s weights, and refines them by an iterative process.
+In the single-output case, our approach leads to stronger theoretical guarantees.
+Our method also has support for convolutional layers.
 6 Experiments
-We conduct experiments targeting the compression of the linear and convolutional layers of both convolutional and deep neural networks.
-Our method TropNNC performs layer-by-layer compression of both linear and convolutional layers and is suitable for regression and classification tasks.
-The empirical results from these experiments demonstrate that our algorithm achieves the expected performance as claimed.
-Throughout our experiments, we reserve the use of the iterative variant of TropNNC for cases where its slower performance is justified by the expectation of significantly better results compared to the non-iterative variant.
 Baselines.
-We compare our framework with state-of-the-art methods that conduct structured pruning without requiring re-training (or fine-tuning).
-Specifically, we compare it with the algorithms proposed by Misiakos et al.
-(2022), ThiNet (Luo et al., 2017), CUP (Duggal et al., 2021) and the simple baselines random and L1 structured pruning.
-The algorithms presented in (Misiakos et al., 2022) are originally designed for linear layers only.
-To enable a fair comparison in compressing convolutional layers, we extended their approach using our proposed technique, making them applicable to convolutional layers as well.
-ThiNet employs a greedy criterion to remove neurons and channels that contribute least to the subsequent layer’s input.
-For the non-uniform pruning variant of our framework, we compare it with CUP.
-CUP assigns to each neuron/channel a feature, performs hierarchical clustering based on these features, and replaces each neuron/channel in each cluster by a cluster representative according to a maximum norm criterion.
+We empirically evaluate TropNNC against baselines that perform structured pruning that do not require retraining.
+Specifically, as baselines we choose Neural Path K-Means by Misiakos et al.
+(2022), ThiNet, CUP, and the simple random and L1 structured pruning baselines.
+Neural Path K-Means is originally designed for linear layers only.
+To enable a fair comparison, we extended their approach using our proposed technique, making them applicable to convolutional layers as well.
+For the non-uniform variant of our framework, we compare it with CUP.
 As explained, our proposed algorithm enhances all three steps of CUP.
-Through an ablation study, we found that each improvement is essential for achieving a competitive advantage over CUP, indicating that these enhancements complement each other.
-In the presented experiments, we compare CUP exclusively with the fully enhanced version of TropNNC, incorporating all improvements.
-Furthermore, the trivial baseline of random structured pruning discards neurons or channels based on a uniform probability distribution, while L1 structured pruning targets those with the smallest L1 norm of their weights or kernels.
+In the presented experiments, we compare CUP exclusively with the fully enhanced version of TropNNC.
+To eliminate randomness, all random methods were performed 5 times on each model and the best performing compressed model was selected for testing.
+All metrics report the average over 5 repetitions of the same experiments with error bars indicating the standard deviation.
 Datasets and networks.
-We evaluate our framework on the MNIST and CIFAR datasets, testing it across various models including simple multi-layer perceptrons (MLPs), convolutional neural networks (CNNs), LeNet (LeCun et al., 1998), AlexNet (Krizhevsky et al., 2012), and VGG (Simonyan and Zisserman, 2015).
-The non-uniform variant of our algorithm is tested on the CIFAR and ImageNet datasets, with its performance evaluated across models such as VGG, ResNet56, and ResNet18 (He et al., 2016).
-7.1 MNIST Dataset, Pairs 3-5 and 4-9
-The first experiment is performed on the binary classification tasks of pairs 3/5 and 4/9 of the MNIST dataset and so we can utilize both single-output and multi-output versions of our algorithm for the compression of the final hidden linear layer.
-In Tables 1 and 2, we compare single-output TropNNC and TropNNC with the corresponding single-output and multi-output algorithms from Misiakos et al.
-(2022).
-We use the same network, a simple CNN network with two fully connected layers and a final hidden linear layer of size $1000$, and the same pruning ratios as in the corresponding experiment of Misiakos et al.
-(2022).
-According to Tables 1 and 2, our algorithm performs better than Misiakos et al.
-(2022) with particular difference in performance when using $0.3\%$ of the nodes of the hidden layer.
-Percentage of Remaining Neurons Zonotope K-means TropNNC, single output Neural Path K-means TropNNC
-100.0 99.42 $\pm$ 0.05 99.42 $\pm$ 0.05 99.44 $\pm$ 0.06 99.44 $\pm$ 0.06
-10.0 99.38 $\pm$ 0.04 99.43 $\pm$ 0.04 99.41 $\pm$ 0.06 99.34 $\pm$ 0.16
-5.0 99.37 $\pm$ 0.06 99.42 $\pm$ 0.06 99.37 $\pm$ 0.04 99.31 $\pm$ 0.16
-1.0 99.33 $\pm$ 0.05 99.41 $\pm$ 0.04 99.32 $\pm$ 0.06 99.33 $\pm$ 0.05
-0.5 99.12 $\pm$ 0.28 99.43 $\pm$ 0.04 99.40 $\pm$ 0.08 99.37 $\pm$ 0.04
-0.3 95.98 $\pm$ 3.16 99.43 $\pm$ 0.04 98.39 $\pm$ 0.99 99.39 $\pm$ 0.10
-Table 1: Comparison of methods on MNIST-3_5
-Percentage of Remaining Neurons Zonotope K-means TropNNC, single output Neural Path K-means TropNNC
-100.0 99.55 $\pm$ 0.04 99.55 $\pm$ 0.04 99.56 $\pm$ 0.04 99.56 $\pm$ 0.04
-10.0 99.60 $\pm$ 0.07 99.54 $\pm$ 0.04 99.52 $\pm$ 0.04 99.50 $\pm$ 0.05
-5.0 99.56 $\pm$ 0.04 99.54 $\pm$ 0.04 99.52 $\pm$ 0.04 99.50 $\pm$ 0.03
-1.0 99.35 $\pm$ 0.29 99.54 $\pm$ 0.04 99.49 $\pm$ 0.14 99.47 $\pm$ 0.11
-0.5 98.03 $\pm$ 0.71 99.54 $\pm$ 0.04 99.01 $\pm$ 0.50 99.48 $\pm$ 0.12
-0.3 86.38 $\pm$ 7.59 99.54 $\pm$ 0.04 94.25 $\pm$ 4.57 99.47 $\pm$ 0.09
-Table 2: Comparison of methods on MNIST-4_9
+We evaluate our framework on the MNIST (Deng, 2012) and CIFAR (Krizhevsky and Hinton, 2009) datasets, testing it across various models including simple multi-layer perceptrons (MLPs), convolutional neural networks (CNNs), AlexNet (Krizhevsky et al., 2012), and VGG (Simonyan and Zisserman, 2015).
+The non-uniform variant of our algorithm is tested on the CIFAR and ImageNet (Deng et al., 2009) datasets, with its performance evaluated across models such as VGG, and ResNet18 (He et al., 2016).
+The selected datasets and models are widely used benchmarks for neural network compression, and offer diversity in their complexity.
+We compare the methods based on test accuracies and FLOPs reduction, following the literature.
+Experimental setup.
+Only a single hyper-parameter needed tuning: ThiNet’s training samples used.
+For each dataset, we gradually increased this number until no further improvement of ThiNet was observed.
+$5000$ was found to suffice for both MNIST and CIFAR.
+The same number can be found by using the method presented in (Luo et al., 2017) for selecting the parameter.
 6.1 MNIST and Fashion-MNIST Datasets
-The second experiment is performed on the MNIST and FASHION-MNIST datasets.
-This classification task is multiclass, and thus only our multi-output algorithm may be used.
-Table 3 compares TropNNC with Neural Path K-means of Misiakos et al.
-(2022) for the same CNN network as above, and for the same pruning ratios as in (Misiakos et al., 2022).
-We again compress the final linear layer.
+The first experiment is performed on the MNIST and Fashion-MNIST datasets.
+Table 1 compares TropNNC with Neural Path K-means for the same CNN network and the same pruning ratios as in (Misiakos et al., 2022).
+We compress the final linear layer.
 As shown in the results, for both datasets, our algorithm outperforms Neural Path K-means.
-To evaluate the performance of our algorithm in compressing linear layers of deeper networks, we applied TropNNC to ”deepNN”, a fully connected neural network with layer sizes $28\times 28,512,256,128$, and $10$.
-The performance plots are provided in Figures 7(a) and 7(b).
-As illustrated, TropNNC outperforms both Neural Path K-means from (Misiakos et al., 2022) and ThiNet from (Luo et al., 2017).
-To assess the performance of our algorithm in compressing convolutional layers, we applied TropNNC to ”deepCNN2D”, a LeNet-type convolutional neural network with ReLU activations.
-The performance plots are provided in Figures 7(c) and 7(d).
-The results demonstrate that TropNNC outperforms Neural Path K-means from (Misiakos et al., 2022) and matches, or even surpasses, ThiNet from (Luo et al., 2017).
+To evaluate the performance of our algorithm in compressing linear layers of deeper networks, we applied TropNNC to "deepNN", a MLP with layer sizes $28\times 28,512,256,128$, and $10$.
+The performance plots are provided in Figures 4(a) and 4(b).
+As illustrated, Neural Path K-Means achieves results comparable to L1-structured pruning.
+By contrast, TropNNC outperforms both Neural Path K-means and ThiNet.
+To assess the performance of our algorithm in compressing convolutional layers, we applied TropNNC to "deepCNN2D", a LeNet-type convolutional neural network with ReLU activations.
+The performance plots are provided in Figures 4(c) and 4(d).
+The results demonstrate that TropNNC significantly outperforms Neural Path K-means, whose effectiveness appears reduced.
+Moreover, TropNNC matches, or even surpasses, ThiNet.
+Table 1: Comparison of Neural Path K-means and TropNNC on MNIST and Fashion-MNIST
 Percentage of Remaining Neurons MNIST Fashion-MNIST
@@
 5.0 95.17 $\pm$ 2.36 97.06 $\pm$ 0.73 81.48 $\pm$ 3.90 87.42 $\pm$ 0.46
-Table 3: Comparison of Neural Path K-means and TropNNC on MNIST and FASHION-MNIST
 (a) deepNN, MNIST
@@
 Figure 4: Compression of linear and convolutional layers of ReLU neural networks on MNIST datasets.
-7.3 CIFAR Datasets
+6.2 CIFAR-10 and -100 Datasets
 In this experiment, we compress AlexNet and VGG trained on the CIFAR-10 and CIFAR-100 datasets to assess the performance of each compression method.
-Figures 8(a) and 8(b) illustrate the compression of the linear layers of AlexNet on CIFAR-10 and CIFAR-100, respectively.
-Additionally, Figures 8(c) and 8(d) show the compression of VGG’s convolutional layers for these datasets.
-For VGG, we applied compression to the fused networks without batch normalization, as we found that ThiNet’s performance on the original networks was practically identical to its performance on the fused networks.
-The results indicate that for larger datasets, TropNNC consistently outperforms Neural Path K-means from (Misiakos et al., 2022).
-Furthermore, it at least matches or even surpasses the performance of ThiNet from (Luo et al., 2017) for the compression of linear layers.
+Figures 5(a) and 5(b) illustrate the compression of the linear layers of AlexNet on CIFAR-10 and CIFAR-100, respectively.
+Additionally, Figures 5(c) and 5(d) show the compression of VGG’s convolutional layers for these datasets.
+The results indicate that for larger datasets, TropNNC consistently outperforms Neural Path K-means, whose performance approaches that of random pruning.
+Furthermore, it matches or even surpasses the performance of ThiNet for the compression of linear layers.
 These findings highlight the effectiveness of TropNNC in handling more complex and larger-scale data scenarios.
 For the compression of convolutional layers of VGG, TropNNC matches ThiNet.
-For the compression of convolutional layers of VGG, we do not present results for iterative TropNNC because it had almost identical performance to the non-iterative algorithm.
-(a) AlexNet, linear, CIFAR10
-(b) AlexNet, linear, CIFAR100
+(a) AlexNet, lin., CIFAR10
+(b) AlexNet, lin., CIFAR100
 (c) VGG, conv., CIFAR10
@@
 Figure 5: Compression of linear layers of AlexNet and convolutional layers of VGG on CIFAR datasets.
-7.4 Non-ReLU Activation
-In this experiment, we evaluate the performance of our algorithm in compressing networks that do not use ReLU activations, such as LeNet (LeCun et al., 1998).
-We conduct experiments on the MNIST and FASHION-MNIST datasets, focusing on compressing the linear and convolutional layers of LeNet.
-The results are presented in Figures 9(a), 9(b), 9(c) and 9(d).
-Interestingly, our algorithm demonstrates robust performance even on networks with non-ReLU activations.
-(a) LeNet, linear, MNIST
-(b) LetNet, linear, F-MNIST
-(c) LeNet, conv., MNIST
-(d) LetNet, conv., F-MNIST
-Figure 9: Compression of LeNet on MNIST Datasets.
 6.3 Non-uniform Pruning
-In this section, we evaluate the effectiveness of the non-uniform variant of TropNNC.
-Our experiments here compress various models, such as VGG on CIFAR-10, ResNet56 on CIFAR-10, and ResNet18 on ImageNet.
-The results are summarized in Tables 4, 5, 6, and 7.
+We evaluate the effectiveness of the non-uniform variant of TropNNC.
+We compress various models, such as VGG on CIFAR-10, and ResNet18 on ImageNet.
+The results are summarized in Tables 2 and 3.
 We should emphasize that we did not apply any form of fine-tuning or re-training.
-In Table 4 we present the compression of fused models, i.e.
-models whose batch normalization has been fused into its preceding convolutional or linear layer, as explained in a previous section.
-We noticed that the performance of CUP was better on the original networks with batch normalization.
-In Tables 5, 6 and 7 we compare our performance with the performance of CUP compressing the original networks.
-For our method only, we applied batch normalization fusion following the second approach described earlier—fusing the batch normalization layer with the preceding layer on a layer-by-layer basis, followed by compression of the fused layer before moving on with the next layer.
-We observed that our method performed significantly better when the clustering vectors were generated using the pre-fusion weights/kernels of the layers.
-While the reason for this improvement remains unclear, it consistently led to better performance in our experiments.
+Table 4 provides an ablation of the iterative variant of TropNNC.
 Our findings indicate that our method demonstrates a clear advantage.
-Notably, our approach shows a significant performance improvement for the VGG model, while the benefits are comparatively modest for ResNet models.
-The substantial advantage observed with VGG remains somewhat unexplained, and we have yet to determine why this effect is less pronounced for ResNet architectures.
+Our approach shows a significant performance improvement for the VGG model, while the benefits are comparatively modest for ResNet18.
+The substantial advantage observed with VGG remains somewhat unexplained, and we have yet to determine why this effect is less pronounced for the ResNet architecture.
 Further investigation may be required to fully understand these discrepancies and to optimize our approach across different model types.
@@
 Meanwhile, the second variant, like CUP, also targets the initial layers, where fewer parameters are present, but a greater number of operations is required.
-Model Method Threshold #params $\downarrow$ FLOPS $\downarrow$ Acc. $\uparrow$
-VGG Original - 14.72M 0.63G 93.64
-CIFAR10 CUP 0.15 4.70M 0.40G 56.22
-TropNNC (v2) 1.12 4.70M 0.33G 90.82
-ResNet56 Original - 0.85M 0.25G 93.67
-CIFAR10 CUP 0.55 0.79M 0.18G 45.89
-TropNNC (v2) 1.2 0.65M 0.18G 76.50
-ResNet18 Original - 11.68M 3.63G 69.10
-ImageNet CUP 0.6 11.62M 3.50G 58.20
-TropNNC (v2) 1 11.60M 3.47G 62.60
-Table 4: Comparison of CUP and TropNNC (variant 2) accuracy across different pruning thresholds for fused models without batch normalization.
+Table 2: Comparison of CUP and TropNNC (variant 1) accuracy across different pruning thresholds on CIFAR10, VGG.
 Method Threshold #params $\downarrow$ FLOPS $\downarrow$ Acc. $\uparrow$
@@
 CUP 0.15 6.24M 0.46G 92.41
-TropNNC (v1) 0.006 5.70M 0.44G 93.58
+TropNNC (3 iters) 0.014 5.22M 0.46G 93.58
 CUP 0.20 4.62M 0.42G 67.91
-TropNNC (v1) 0.0085 4.32M 0.41G 93.46
+TropNNC (3 iters) 0.017 3.61M 0.42G 93.02
 CUP 0.25 3.61M 0.39G 16.02
-TropNNC (v1) 0.011 3.53M 0.39G 93.21
-Table 5: Comparison of CUP and TropNNC (variant 1) accuracy across different pruning thresholds on CIFAR10, VGG.
+TropNNC (3 iters) 0.02 2.75M 0.39G 91.71
+Table 3: Comparison of CUP and TropNNC accuracy across different pruning thresholds on ImageNet, ResNet18.
 Method Threshold #params $\downarrow$ FLOPS $\downarrow$ Acc. $\uparrow$
-Original - 0.85M 0.25G 93.67
-CUP 0.55 0.72M 0.20G 87.32
-TropNNC (v2) 1.17 0.71M 0.19G 90.53
-CUP 0.6 0.63M 0.17G 78.83
-TropNNC (v2) 1.23 0.63M 0.17G 87.04
-CUP 0.65 0.54M 0.15G 71.48
-TropNNC (v2) 1.3 0.54M 0.15G 73.03
-TropNNC (v1) 0.067 0.49M 0.19G 82.70
-Table 6: Comparison of CUP and TropNNC accuracy across different pruning thresholds on CIFAR10, ResNet56.
+Original - 11.69M 3.64G 69.75
+CUP 0.5 11.66M 3.58G 66.80
+TropNNC (v2) 1.1 11.66M 3.48G 68.95
+CUP 0.6 11.49M 3.38G 53.40
+TropNNC (v2) 1.2 11.46M 3.25G 61.30
+CUP 0.65 11.17M 3.15G 28.45
+TropNNC (v2) 1.25 10.98M 2.92G 41.65
+TropNNC (v1) 0.0187 9.84M 3.46G 59.15
+Table 4: Comparison of non-uniform TropNNC and iterative TropNNC (variant 1) accuracy on CIFAR10, VGG.
 Method Threshold #params $\downarrow$ FLOPS $\downarrow$ Acc. $\uparrow$
-Original - 11.69M 3.64G 70.10
-CUP 0.5 11.66M 3.58G 66.60
-TropNNC (v2) 1.1 11.64M 3.44G 68.40
-CUP 0.6 11.49M 3.38G 53.20
-TropNNC (v2) 1.2 11.27M 3.07G 57.90
-CUP 0.65 11.17M 3.15G 29.20
-TropNNC (v2) 1.25 10.71M 2.75G 41.20
-TropNNC (v1) 0.021 9.87M 3.46G 58.10
-Table 7: Comparison of CUP and TropNNC accuracy across different pruning thresholds on ImageNet, ResNet18.
-8 Conclusions and Future work
-We improved upon the theoretical results of Misiakos et al.
-(2022) by using the (usual) Hausdorff distance instead of its discrete counterpart, to obtain a tighter bound for the approximation of tropical polynomials.
-Leveraging this enhancement we refined the choice of compressed node’s weights that correspond to the K-means clusters from (Misiakos et al., 2022).
-Our novel algorithm, TropNNC, applied to linear and convolutional layers of deep neural networks outperforms Neural Path K-means of Misiakos et al.
-(2022), and manages to match or even surpass the performance of ThiNet, particularly for the compression of linear layers.
-In non-uniform pruning it performs better than CUP, with significant improvement in the case of the VGG network architecture.
-These achievements underscore the potential of tropical geometry in the realm of neural network compression.
-One promising direction for further investigation could involve optimizing the approximation techniques by considering the upper hulls of the network’s tropical polynomials.
-Appendix A Proofs of Propositions and Lemmas
-A.1 Proof of Lemma 2
-Proof  Since $(\mathbf{a}^{T},b)$ lies below the upper envelope of $P$, there exists a point
+Original - 14.7M 0.63G 93.64
+TropNNC 0.02 2.75M 0.39G 86.2
+TropNNC (3 iters) 0.021 2.52M 0.38G 91.2
+TropNNC 0.025 1.80M 0.33G 34.89
+TropNNC (3 iters) 0.0255 1.76M 0.33G 69.95
+7 Conclusion
+We proposed TropNNC, a tropical geometrical method for structured, data-free pruning of linear and convolutional layers of ReLU activated neural networks.
+Based on clustering similar neurons/channels, it selects better cluster representatives than previous work by using tropical geometry and optimization, with provable compression guarantees.
+TropNNC significantly outperforms prior work on tropical geometric pruning and manages to match or even surpass the performance of the data-driven ThiNet.
+In non-uniform pruning it outperforms CUP, with significant improvement in the case of the VGG architecture.
+Our findings highlight the potential of tropical geometry in the realm of neural network compression.
+Acknowledgments
+The research project was supported by the Hellenic Foundation for Research and Innovation (H.F.R.I.) under the “2nd Call for H.F.R.I.
+Research Projects to support Faculty Members & Researchers” (Project Number:2656, Acronym: TROGEMAL).
+Appendix
+In this appendix, we first go into more detail regarding the heuristic improvements of our algorithm.
+Afterwards, we introduce the background of tropical algebra which was omitted from the main text.
+Finally, we provide the proofs for the theorems and propositions stated in the main text.
+Before we proceed with the technical content of this appendix, we first provide an ethical statement for this work and a disclosure of the limitations of our method.
+Ethical statement
+This work does not introduce any new ethical concerns beyond those already present in prior research on network pruning and model compression.
+As with existing pruning methods, our approach aims to reduce model size and computational cost, and does not involve the use of sensitive data, human subjects, or deployment in high-risk applications.
+Limitations
+A key limitation of our method is that it relies on the presence of consecutive convolutional or linear layers – that is, the existence of intermediate hidden layers.
+In addition, it only holds for networks with ReLU activations.
+Heuristic Improvements
+In the experiments comparing TropNNC and CUP, TropNNC was enhanced with 2 heuristic improvements.
+These improvements are not rigorously backed by theory; however, we try to provide intuitive explanations as to why they work.
+Heuristic 1.
+First, let us focus on the single-output case, where the problem of network compression reduces to the problem of zonotope approximation.
+In this setting, Algorithm 1 suggests clustering similar generators based on their $L^{2}$ distances and replacing each cluster with the sum of its elements.
+The intuition is as follows: if the generators within a cluster are exactly equal, then through their Minkowski sum, these generators become equivalent to a single generator given by their sum.
+We can take this one step further: if we have generators that are parallel to each other, their Minkowksi sum reduces to a single equivalent generator without introducing any error.
+This suggest that generator similarity is determined by direction (i.e.
+cosine similarity) and not $L^{2}$ distance.
+Hence, an improvement to the single-output algorithm is to cluster the generators based on cosine similarity.
+In practice, this can be achieved by applying K-Means clustering to normalized generators, while still using the original generators in the summation step.
+A similar heuristic can also be applied to the multi-output algorithm.
+By normalizing the input weights of the clustering vectors (i.e.
+the vectors used for clustering similar neurons/channels, not the weights used for calculating the representative of each cluster) we obtain an improvement to the algorithm.
+Heuristic 2.
+The second heuristic is motivated by Theorem 9, which states that two tropical polynomials are functionally identical if the upper hulls of their extended Newton polytopes coincide (i.e.
+only the vertices of the upper hull contribute to non-redundant terms).
+Indeed, this theorem also serves as motivation for Theorem 2, which shows that approximate equality of the extended Newton polytopes implies approximate functional equivalence of the corresponding tropical polynomials.
+Although we do not formally prove it, this theorem can be further refined by bounding the error of two tropical polynomials by the Hausdorff distance of the upper hulls of their extended Newton polytopes:
+$)\frac{1}{\rho}\max_{\mathbf{x}\in B}|p(\mathbf{x})-\tilde{p}(\mathbf{x})|\leq H(UF(P),UF(\tilde{P}$))
+Based on this refined result, it becomes clear that to compress networks, it suffices to approximate the upper hulls of the zonotopes of the network, rather than the zonotopes themselves.
+We can take this idea one step further.
+Consider the upper hull of an extended Newton polytope, and project it by eliminating the bias term.
+Then, the points on the convex hull of the resulting projection (which coincides with the convex hull of the (non-extended) Newton polytope) correspond to the slopes of the outer linear regions of the polynomials, i.e.
+the regions which extend to infinity.
+We can view these points as being the "outer" points of the upper hull.
+When approximating a tropical polynomial, these unbounded regions are particularly important, because they have the potential to make the sup-norm of the approximation error infinite (something which is not captured by Theorem 2 because we divide by $\rho$ and take the $L^{1}$-norm error on a bounded set).
+Based on this insight, our goal is to design an algorithm that prioritizes the approximation of the upper hulls of the zonotopes – particularly the vertices corresponding to the unbounded regions of the tropical polynomials.
+(i.e.
+the "outer" vertices of the upper hulls).
+Let us first see how vanilla TropNNC behaves on a simple example.
+Consider the tropical polynomial $v(x)=\max(0,-x+5)+\max(0,x+5)+\max(0,x)$.
+Such a polynomial and its corresponding zonotope are depicted in Figure 6(a).
+Notice that the upper hull has 4 points, each corresponding to a linear region of the polynomial.
+It also has 2 "outer" points, which correspond to the 2 outer regions of the polynomial extending to infinity.
+We reduce the terms from 3 to $K=2$.
+First, TropNNC performs clustering on the generators of the zonotope, forming the clusters $\{(-1,5),(1,5)\}$ and $\{(1,0)\}$.
+It then replaces the first cluster with the representative $(0,10)$ and the second cluster with the representative $(1,0)$.
+The reduced tropical polynomial is $\tilde{v}(x)=\max(0,10)+\max(0,x)=10+\max(0,x)$.
+The compression procedure and the corresponding reduced function is depicted in Figure 6(b).
+It is evident that this approximation is not ideal: while the reduced polynomial accurately captures the two central regions, it fails to preserve the unbounded regions.
+This causes the approximation error becomes unbounded.
+Since we have $K=2$, the best possible approximation using only 2 terms is $v^{\prime}(x)=\max(0,-x+5)+\max(0,2x+5)$.
+This approximation captures both the important outer regions with zero error, and also one of the central regions.
+The question however remains: how could we have arrived at such an approximation?
+To answer this, notice that $v^{\prime}(x)=\max(0,-x+5)+\max(0,(x+5)+(x))$.
+Thus, if TropNNC had initially formed the clusters $\{(-1,5)\}$ and $\{(1,5),(1,0)\}$, then it would have been able to find the better approximation.
+Now, the modification we have to make becomes clear: instead of forming the clusters based on the generators themselves, we can form the clusters based on the reduced generators with the bias suppressed (and, of course, for the summation step the bias remains).
+The modified compression procedure is depicted in Figure 6(c).
+The new approximation clearly provides a closer fit to the upper hull and captures the “outer” points exactly, resulting in zero error in the outer unbounded regions.
+To conclude, another heuristic improvement comes from dropping the bias from the clustering vectors, seeing as this helps the algorithm achieve better upper hull approximations and prioritize the important "outer" vertices of the upper hull.
+(a) Example tropical polynomial $v(x)=\max(0,-x+5)+\max(0,x+5)+\max(0,x)$ and corresponding zonotope.
+(b) Approximation using vanilla TropNNC.
+(c) Approximation using Heuristic 2.
+Figure 6: Example of reduction of tropical polynomial.
+Background on Tropical Algebra
+We begin with the definitions of tropical polynomials, tropical rational functions, Newton polytopes, and zonotopes.
+Tropical Polynomials and Rational Functions.
+Within the max-plus semiring, we can define polynomials. A tropical polynomial $f$ in $d$ variables $\mathbf{x}=(x_{1},\ldots,x_{d})$ is defined as the function:
+$f(\mathbf{x})=\bigvee_{i=1}^{n}\left\{\mathbf{a}_{i}^{T}\mathbf{x}+b_{i}\right\}=\max_{i\in[n]}\left\{\mathbf{a}_{i}^{T}\mathbf{x}+b_{i}\right\},$ (4)
+where $n]:=\{1,...,n\}$.
+Here, $n$ represents the rank of the tropical polynomial.
+Each monomial term $\{\mathbf{a}_{i}^{T}\mathbf{x}+b_{i}\}$ of the polynomial has an exponent or slope $\mathbf{a}_{i}\in\mathbb{R}^{d}$ and a coefficient or bias $b_{i}\in\mathbb{R}$.
+Each monomial term corresponds to a plane in $\mathbb{R}^{d+1}$.
+Consequently, tropical polynomials are piecewise linear convex functions.
+Specifically, every tropical polynomial is a continuous piecewise linear convex function, and every continuous piecewise linear convex function can be expressed (though not uniquely) as a tropical polynomial (Maclagan and Sturmfels, 2021).
+The set of tropical polynomials in $\mathbf{x}$ defines the semiring $\mathbb{R}_{\mathrm{max}}[\mathbf{x}]$.
+Figures 7(a), 7(b) illustrate examples of tropical polynomials in one and multiple variables, respectively.
+Tropical rational functions are defined as the difference of two tropical polynomials $p$ and $q$:
+$r(\mathbf{x})=p(\mathbf{x})-q(\mathbf{x})$
+Tropical rational functions correspond to general piecewise linear functions.
+Specifically, every tropical rational function is a continuous piecewise linear function, and every continuous piecewise linear function can be expressed (though not uniquely) as a tropical rational function.
+Figures 7(c), 7(d) provide examples of tropical rational functions.
+Figure 7: (a) depicts a single-variate tropical polynomial, (b) depicts a multi-variate tropical polynomial, (c) depicts a single-variate tropical rational function, (d) depicts a multi-variate tropical rational function
+Newton Polytopes.
+For a tropical polynomial $f$ as defined in (4), we define its Newton polytope as the convex hull of the slopes $\mathbf{a}_{i}$ of $f$
+$\mathrm{Newt}(f):=\mathrm{conv}\left\{\mathbf{a}_{i},i\in[n]\right\}.$
+Additionally, we define the extended Newton polytope of a tropical polynomial $f$ as the convex hull of the slopes $\mathbf{a}_{i}$ of $f$ extended in their last dimension by the coefficient $b_{i}$
+$\mathrm{ENewt}(f):=\mathrm{conv}\left\{(\mathbf{a}_{i}^{T},b_{i}),i\in[n]\right\}.$
+The following proposition allows us to calculate the (extended) Newton polytope of expressions of tropical polynomials, and it was used in the main text to deduce that the extended Newton polytopes of the outputs of the network are zonotopes.
+Proposition 7 (Zhang et al., 2018).
+Let $f,g\in\mathbb{R}_{\mathrm{max}}[\mathbf{x}]$ be two tropical polynomials in $\mathbf{x}$. For the extended Newton polytope, the following holds:
+$\mathrm{ENewt}\left(f\vee g\right)$ $=\mathrm{conv}\left\{\mathrm{ENewt}(f)\cup\mathrm{ENewt}(g)\right\}$
+$\mathrm{ENewt}(f+g)$ $=\mathrm{ENewt}(f)\oplus\mathrm{ENewt}(g).$
+The Minkowski sum $\oplus$ of two polytopes (or more generally subsets of $\mathbb{R}^{d}$) $P$ and $Q$ is defined as:
+$P\oplus Q:=\{p+q\mid p\in P,q\in Q\}.$
+Using the principle of induction, Proposition 7 can be generalized to any finite tropical expression of tropical polynomials.
+The upper envelope or upper hull $UF(P)$ of an extended Newton polytope is defined as the set of all points $(\mathbf{a}^{T},b)$ of the polytope $P$ that are not "shadowed" by any other part of the polytope when viewed from above (last dimension).
+This means that there is no $b^{\prime}>b$ such that $(\mathbf{a}^{T},b^{\prime})$ belongs to $P$.
+We have the following useful lemma.
+Lemma 8.
+Let $p\in\mathbb{R}_{\mathrm{max}}[\mathbf{x}]$ a tropical polynomial in $d$ variables with extended Newton polytope $P=\mathrm{ENewt}(p)$.
+If $(\mathbf{a}^{T},b)$ lies below the upper envelope of $P$, then $\forall\mathbf{x}\in\mathbb{R}^{d},\mathbf{a}^{T}\mathbf{x}+b\leq p(\mathbf{x})$.
+The inequality is strict if $(\mathbf{a}^{T},b)$ lies strictly below the upper envelope.
+Proof.
+Since $(\mathbf{a}^{T},b)$ lies below the upper envelope of $P$, there exists a point
 $(\mathbf{a}^{T},b^{\prime})=\sum_{i=1}^{k}\lambda_{i}\mathbf{v}_{i},\quad\sum_{i=1}^{k}\lambda_{i}=1$
@@
 $\mathbf{a}^{T}\mathbf{x}+b\leq\mathbf{a}^{T}\mathbf{x}+b^{\prime}=\langle(\mathbf{a}^{T},b^{\prime}),(\mathbf{x},1)\rangle$
-$=\sum_{i=1}^{k}\lambda_{i}\langle\mathbf{v}_{i},(\mathbf{x},1)\rangle\leq\max_{i=1,\ldots,k}\langle\mathbf{v}_{i},(\mathbf{x},1)\rangle\leq p(\mathbf{x}).$
-If $(\mathbf{a}_{i}^{T},b_{i})$ lies strictly below the upper envelope of $P$, then $b^{\prime}>b$, and the inequality is strict.
-A.2 Proof of Lemma 8
-Proof  We will prove that
-$\sup_{\mathbf{u}\in P}\operatorname{dist}(\mathbf{u},\tilde{P})=\max_{\mathbf{u}\in V_{P}}\operatorname{dist}(\mathbf{u},\tilde{P}),$
+$\sum_{i=1}^{k}\lambda_{i}\langle\mathbf{v}_{i},(\mathbf{x},1)\rangle\leq\max_{i=1,\ldots,k}\langle\mathbf{v}_{i},(\mathbf{x},1)\rangle\leq p(\mathbf{x}).$
+If $(\mathbf{a}_{i}^{T},b_{i})$ lies strictly below the upper envelope of $P$, then $b^{\prime}>b$, and the inequality is strict. ∎
+The extended Newton polytope provides a geometrical interpretation for studying tropical polynomials. For instance, the following theorem holds.
+Theorem 9.
+For any two tropical polynomials $f,g\in\mathbb{R}_{\mathrm{max}}[\mathbf{x}]$, the following holds:
+$f=g\Leftrightarrow UF(\mathrm{ENewt}(f))=UF(\mathrm{ENewt}(g))$
+This implies that two tropical polynomials are functionally identical if and only if their extended Newton polytopes have the same upper envelope.
+The above theorems indicate that a tropical polynomial is fully functionally determined by the upper envelope of its extended Newton polytope, as shown by the following example.
+Example 2.
+Consider the polynomials corresponding to Figure 8.
+$f(x,y)=\max\{0,-y+1,y+1\}$
+$g(x,y)=\max\{x+1,-x+1\}.$
+We have that
+$(f\vee g)(x,y)=\max\{0,x+1,y+1,-y+1,-x+1\}$
+$(f+g)(x,y)=\max\{x+1,-x+1,x-y+2,$
+xy+2,x+y+2,x+y+2}.-x-y+2,x+y+2,-x+y+2\}.
+The extended Newton polytopes of $f,g,f\vee g,f+g$ are shown in Figure 8. The polynomial $f\vee g$ can be reduced as follows:
+$(f\vee g)(x,y)=\max\{x+1,y+1,-y+1,-x+1\},$
+which corresponds to the vertices of the upper envelope of $\mathrm{ENewt}(f\vee g)$.
+(a) $f$
+(b) $g$
+(c) $f\vee g$
+(d) $f+g$
+Figure 8: Operations on tropical polynomials.
+$\mathrm{ENewt}(f\vee g)$ corresponds to the convex hull of the union of the vertices of the polytopes $\mathrm{ENewt}(f),\mathrm{ENewt}(g)$.
+$\mathrm{ENewt}(f+g)$ corresponds to the Minkowski sum of $\mathrm{ENewt}(f),\mathrm{ENewt}(g)$.
+For the polytope $\mathrm{ENewt}(f\vee g)$ we illustrate with blue the upper envelope, which consists of a single face.
+The vertices of the upper envelope are the only non reduntant terms of the polynomial $f\vee g$.
+Zonotopes.
+In polytope theory, zonotopes are a special class of convex polytopes that can be defined as the Minkowski sum of a finite set of line segments (or edges).
+Formally, given a set of line segments $g_{1},\ldots,g_{n}$, the zonotope is defined as
+$Z:=\bigoplus_{i\in[n]}g_{i}$
+The line segments $g_{i}$ are referred to as the zonotope’s generators.
+Alternatively, a zonotope can be expressed equivalently by a set of vectors $\mathbf{v}_{1},\ldots,\mathbf{v}_{n}\in\mathbb{R}^{d}$ and a starting point $\mathbf{s}\in\mathbb{R}^{d}$.
+By taking the generators to be the segments $\mathbf{0},\mathbf{v}_{1}],\ldots,[\mathbf{0},\mathbf{v}_{n}]$ and translating the first segment by $\mathbf{s}$, we obtain the equivalent form:
+$Z=\left\{\mathbf{s}+\sum_{i=1}^{n}\lambda_{i}\mathbf{v}_{i}\mid 0\leq\lambda_{i}\leq 1\right\}$
+In this context, the vectors $\mathbf{v}_{i}$ are sometimes referred to as the zonotope’s generators, meaning the segments $\mathbf{0},\mathbf{v}_{i}]$.
+When the starting point $\mathbf{s}$ is not mentioned, it is assumed to be the origin $\mathbf{0}$.
+For a zonotope with a starting point $\mathbf{s}\in\mathbb{R}^{d}$ and generators $\mathbf{v}_{1},\ldots,\mathbf{v}_{n}\in\mathbb{R}^{d}$, a vertex $\mathbf{u}\in V_{Z}$ corresponds to points where $\lambda_{i}=0\ \text{or}\ 1$.
+The vertex $\mathbf{u}$ can be expressed as:
+$\mathbf{u}=\mathbf{s}+\sum_{i\in I}\mathbf{v}_{i},$
+where $I\subseteq[n]$.
+Tropical polynomial approximation based on Hausdorff distance
+Before we proceed with the proof of our refined Theorem 2, we first provide an auxiliary lemma, stated implicitly in the main text, which is necessary to see how our bound is indeed tighter than the bound of Misiakos et al.
+(2022).
+This lemma is also used for the proofs of Theorem 2, and Propositions 4 and 6.
+Lemma 10.
+Due to the convexity and compactness of polytopes, we have that
+$H(P,\tilde{P})=\max\left\{\max_{\mathbf{u}\in V_{P}}\mathrm{dist}(\mathbf{u},\tilde{P}),\max_{\mathbf{v}\in V_{\tilde{P}}}\mathrm{dist}(P,\mathbf{v})\right\}$
+Proof.
+We will prove that
+[…]
 i.e., the supremum is attained at some vertex of $P$.
@@
 $\lambda f(\mathbf{u}_{1})+(1-\lambda)f(\mathbf{u}_{2})\geq\|\lambda\mathbf{u}_{1}+(1-\lambda)\mathbf{u}_{2}-\tilde{\mathbf{u}}\|\geq f(\lambda\mathbf{u}_{1}+(1-\lambda)\mathbf{u}_{2}),$
-which concludes the proof.
-A.3 Proof of Proposition 10 and Theorem 11
-Proof  Consider a point $\mathbf{x}\in B$ and assume that $p(\mathbf{x})=\mathbf{a}^{T}\mathbf{x}+b$ and $\tilde{p}(\mathbf{x})=\mathbf{c}^{T}\mathbf{x}+d$.
+which concludes the proof. ∎
+We continue with the proof of Theorem 2.
+Proof (Theorem 2).
+Consider a point $\mathbf{x}\in B$ and assume that $p(\mathbf{x})=\mathbf{a}^{T}\mathbf{x}+b$ and $\tilde{p}(\mathbf{x})=\mathbf{c}^{T}\mathbf{x}+d$.
 Take an arbitrary $(\mathbf{u}^{T},v)\in\tilde{P}$.
@@
 Finally, we obtain that
-$-\max_{(\mathbf{c}^{T},d)\in V_{\tilde{P}}}\operatorname{dist}((\mathbf{c}^{T},d),P)\cdot\rho\leq p(\mathbf{x})-\tilde{p}(\mathbf{x})\leq\max_{(\mathbf{a}^{T},b)\in V_{P}}\operatorname{dist}((\mathbf{a}^{T},b),\tilde{P})\cdot\rho,$
+$\max_{(\mathbf{c}^{T},d)\in V_{\tilde{P}}}\operatorname{dist}((\mathbf{c}^{T},d),P)\cdot\rho\leq p(\mathbf{x})-\tilde{p}(\mathbf{x})\leq\max_{(\mathbf{a}^{T},b)\in V_{P}}\operatorname{dist}((\mathbf{a}^{T},b),\tilde{P})\cdot\rho,$
 which implies
@@
 Therefore, by Lemma 10, we have,
-$\frac{1}{\rho}\max_{\mathbf{x}\in B}\lvert p(\mathbf{x})-\tilde{p}(\mathbf{x})\rvert\leq H(P,\tilde{P}),$
-which concludes the proof of Proposition 10.
-Finally, notice that we may write
-$\|\mathbf{v}(\mathbf{x})-\tilde{\mathbf{v}}(\mathbf{x})\|_{1}$ $=\sum_{j=1}^{m}\lvert\mathbf{v}_{j}(\mathbf{x})-\tilde{\mathbf{v}}_{j}(\mathbf{x})\rvert$
-$=\sum_{j=1}^{m}\lvert(p_{j}(\mathbf{x})-q_{j}(\mathbf{x}))-(\tilde{p}_{j}(\mathbf{x})-\tilde{q}_{j}(\mathbf{x}))\rvert$
-$=\sum_{j=1}^{m}\lvert(p_{j}(\mathbf{x})-\tilde{p}_{j}(\mathbf{x}))-(q_{j}(\mathbf{x})-\tilde{q}_{j}(\mathbf{x}))\rvert$
-$\leq\sum_{j=1}^{m}\left(\lvert p_{j}(\mathbf{x})-\tilde{p}_{j}(\mathbf{x})\rvert+\lvert q_{j}(\mathbf{x})-\tilde{q}_{j}(\mathbf{x})\rvert\right).$
-By Proposition 10 we derive
-$\frac{1}{\rho}\max_{\mathbf{x}\in B}\|\mathbf{v}(\mathbf{x})-\tilde{\mathbf{v}}(\mathbf{x})\|_{1}\leq\sum_{j=1}^{m}\left(H(P_{j},\tilde{P}_{j})+H(Q_{j},\tilde{Q}_{j})\right).$
-A.4 Proof of Proposition 12
-Proof  For the output function, it holds that
+[…]
+Proofs of Proposition 4 and Corollary 5
+We continue with the proofs of Proposition 4 and Corollary 5.
+Proof (Prop. 4).
+For the output function, it holds that
 $v(\mathbf{x})=p(\mathbf{x})-q(\mathbf{x}),\quad\tilde{v}(\mathbf{x})=\tilde{p}(\mathbf{x})-\tilde{q}(\mathbf{x}).$
 From the triangle inequality, we deduce
-$\frac{1}{\rho}\left|v(\mathbf{x})-\tilde{v}(\mathbf{x})\right|\leq H(P,\tilde{P})+H(Q,\tilde{Q}).$
-Thus, it suffices to get a bound on $H(P,\tilde{P})$ and $H(Q,\tilde{Q})$.
+[…]
+Thus, it suffices to get a bound on H(P,P~)H($\tilde{P}$) and H(Q,Q~)H($\tilde{Q}$).
 Let $I_{+}\subseteq[n]$ and $I_{-}\subseteq[n]$ be the sets of positive and negative generators, respectively.
-First, we deal with $H(P,\tilde{P})$.
+First, we deal with H(P,P~)H($\tilde{P}$).
 Notice that $\forall\mathbf{v}\in V_{\tilde{P}},\mathbf{v}$ is the sum of generators of some clusters of $P$.
@@
 $\sum_{i\in I_{k}}x_{i}=1\Rightarrow\tilde{x}_{k}=\sum_{i\in I_{k}^{\prime}}x_{i}\leq 1$
-Thus, for every cluster $k\in C_{+}$, the point $\tilde{x}_{k}\tilde{c}_{k}(\tilde{\mathbf{a}}_{k}^{T},\tilde{b}_{k})$ lies inside the segment $[\mathbf{0},\tilde{c}_{k}(\tilde{\mathbf{a}}_{k}^{T},\tilde{b}_{k})]$ and thus belongs to $\tilde{P}$.
+Thus, for every cluster $k\in C_{+}$, the point $\tilde{x}_{k}\tilde{c}_{k}(\tilde{\mathbf{a}}_{k}^{T},\tilde{b}_{k})$ lies inside the segment $\mathbf{0},\tilde{c}_{k}(\tilde{\mathbf{a}}_{k}^{T},\tilde{b}_{k})]$ and thus belongs to $\tilde{P}$.
 For the vertex $\mathbf{u}$, we choose to compare it with the point
@@
 Thus, we have that
-$\operatorname{dist}(\mathbf{u},\tilde{P})$ $\leq\left\|\sum_{i\in I_{+}^{\prime}}c_{i}(\mathbf{a}_{i}^{T},b_{i})-\sum_{k\in C_{+}}\tilde{x}_{k}\tilde{c}_{k}(\tilde{\mathbf{a}}_{k}^{T},\tilde{b}_{k})\right\|$
+[…]
 $\leq\sum_{k\in C_{+}}\left\|\sum_{i\in I_{k}^{\prime}}c_{i}(\mathbf{a}_{i}^{T},b_{i})-\tilde{x}_{k}\tilde{c}_{k}(\tilde{\mathbf{a}}_{k}^{T},\tilde{b}_{k})\right\|$
@@
 Finally, we have
-$H(P,\tilde{P})$ $=\max\left\{\max_{\mathbf{u}\in V_{P}}\operatorname{dist}(\mathbf{u},\tilde{P}),\max_{\mathbf{v}\in V_{\tilde{P}}}\operatorname{dist}(P,\mathbf{v})\right\}$
+[…]
 $\leq\max\left\{\sum_{i\in I_{+}}\min\left\{\left\|c_{i}(\mathbf{a}_{i}^{T},b_{i})\right\|,\delta_{\max}\right\},0\right\}$
 $=\sum_{i\in I_{+}}\min\left\{\left\|c_{i}(\mathbf{a}_{i}^{T},b_{i})\right\|,\delta_{\max}\right\}.$
-Similarly, for $H(Q,\tilde{Q})$, we have
+Similarly, for H(Q,Q~)H($\tilde{Q}$), we have
 $H(Q,\tilde{Q})\leq\sum_{i\in I_{-}}\min\left\{\left\|c_{i}(\mathbf{a}_{i}^{T},b_{i})\right\|,\delta_{\max}\right\}.$
@@
 $\frac{1}{\rho}\left|v(\mathbf{x})-\tilde{v}(\mathbf{x})\right|\leq\sum_{i\in I}\min\left\{\left\|c_{i}(\mathbf{a}_{i}^{T},b_{i})\right\|,\delta_{\max}\right\},$
-which concludes the proof of Proposition 12.
-A.5 Proof of Corollary 13
-Proof  The bound of Misiakos et al. (2022) is the following:
+which concludes the proof of Proposition 4. ∎
+Proof (Corollary 5).
+The bound of Misiakos et al. (2022) is the following:
 $\frac{1}{\rho}\max_{\mathbf{x}\in B}|v(\mathbf{x})-\tilde{v}(\mathbf{x})|\leq K\delta_{\max}+\left(1-\frac{1}{N_{\max}}\right)\sum_{i=1}^{n}|c_{i}|\|(\mathbf{a}_{i}^{T},b_{i})\|.$
@@
 $\sum_{i\in I_{k}}\min\left\{0,\delta_{\max}-|c_{i}|\|(\mathbf{a}_{i}^{T},b_{i})\|\right\}+\frac{1}{|I_{k}|}\sum_{i\in I_{k}}|c_{i}|\|(\mathbf{a}_{i}^{T},b_{i})\|\\ \leq\delta_{\max}-\max_{i\in I_{k}}|c_{i}|\|(\mathbf{a}_{i}^{T},b_{i})\|+\frac{1}{|I_{k}|}\sum_{i\in I_{k}}|c_{i}|\|(\mathbf{a}_{i}^{T},b_{i})\|\leq\delta_{\max},$
-which concludes the proof.
-A.6 Proof of Proposition 14
+which concludes the proof. ∎
+Proof of Proposition 6
 Before we proceed with the proof of Proposition 6, we first give the definition of null neurons and generators.
-Definition 15 (Null neuron/generator)
+Definition 11 (Null neuron/generator).
 A neuron/generator $i\in[n]$ that belongs to cluster $k\in[K]$ is a null neuron/generator with respect to output $j\in[m]$ if $\tilde{c}_{jk}c_{ji}\leq 0$.
 $N_{j}$ is the set of all null neurons with respect to output $j$.
-Proof  Assume the algorithm iterative scheme has reached a stationary point (otherwise, assume the last step is an output weight update and the proof works fine).
+Proof (Prop. 6).
+Assume the algorithm iterative scheme has reached a stationary point (otherwise, assume the last step is an output weight update and the proof works fine).
 First, we focus on a single output, say $j$-th output.
@@
 $\sum_{i\in I_{jk}}x_{ji}=1\Rightarrow\tilde{x}_{jk}=\sum_{i\in I_{jk}^{\prime}}x_{ji}\leq 1$
-Thus, for every cluster $k\in C_{j+}$, the point $\tilde{x}_{jk}\tilde{c}_{jk}(\tilde{\mathbf{a}}_{k}^{T},\tilde{b}_{k})$ lies inside the segment $[\mathbf{0},\tilde{c}_{jk}(\tilde{\mathbf{a}}_{k}^{T},\tilde{b}_{k})$], and thus belongs to $\tilde{P_{j}}$.
+Thus, for every cluster $k\in C_{j+}$, the point $\tilde{x}_{jk}\tilde{c}_{jk}(\tilde{\mathbf{a}}_{k}^{T},\tilde{b}_{k})$ lies inside the segment $\mathbf{0},\tilde{c}_{jk}(\tilde{\mathbf{a}}_{k}^{T},\tilde{b}_{k})$], and thus belongs to $\tilde{P_{j}}$.
 For the vertex $\mathbf{u}$, we choose to compare it with the point
@@
 $+\sum_{j=1}^{m}\sum_{i\in N_{j}}|c_{ji}|\|(\mathbf{a}_{i}^{T},b_{i})\|,$
-as desired.
+as desired. ∎